### v1

# Chordality and Shellability of Higher Independence Complexes

Abstract.
This paper investigates the $r$-independence complex $\mathcal{I}_{r}(G)$ of a graph $G$ and its associated hypergraph $\mathcal{C}_{r}(G)$ from combinatorial and algebraic perspectives. For an integer $r\geq 1$, the $r$-independence complex $\mathcal{I}_{r}(G)$ consists of all vertex subsets whose induced subgraphs have components of size at most $r$. We study the hypergraph $\mathcal{C}_{r}(G)$, whose independence complex matches $\mathcal{I}_{r}(G)$, focusing on its chordality in the sense of Woodroofe. Hypergraph chordality ensures shellability of $\mathcal{I}_{r}(G)$—implying that the edge ideal of the hypergraph $\mathcal{C}_{r}(G)$ is sequentially Cohen-Macaulay—and guarantees linear resolutions for the edge ideal of the complement hypergraph $\overline{\mathcal{C}_{r}(G)}$, generalizing Fröberg’s theorem.
Contrary to a recent claim [1], we demonstrate that $\mathcal{C}_{r}(G)$ need not be chordal even when $G$ is a tree, and we provide explicit counterexamples. Our results reveal deep connections between graph structure and hypergraph chordality. For general graphs with $n\geq 3$ vertices, $\mathcal{C}_{r}(G)$ is a chordal hypergraph when $n-2\leq r\leq n-1$. For cycles $C_{n}$, this holds precisely in the same range, while forests satisfy chordality for all $r\geq n-5$.
For block graphs, we prove $\mathcal{C}_{2}(G)$ is always chordal, and if the diameter is at most 4, $\mathcal{C}_{r}(G)$ is a chordal hypergraph for all $r\geq 2$. Forests of diameter at most 5 also satisfy chordality for all $r\geq 2$. Furthermore, we characterize when $\mathcal{C}_{3}(G)$ is a chordal hypergraph for block graphs in terms of forbidden substructures, and for trees, this always holds.
We extend these results to specialized block graph classes: for $\mathcal{T}_{2}$-graphs (generalizing lobster trees), $\mathcal{C}_{4}(G)$ is a chordal hypergraph, while for $\mathcal{T}_{1}$-graphs (generalizing caterpillars), $\mathcal{C}_{r}(G)$ is a chordal hypergraph for all $r\geq 5$.
Additionally, we establish shellability conditions for $\mathcal{I}_{r}(G)$ via three key constructions: star-clique attachments to vertex covers (with tight thresholds for chordal graphs), generalized whiskering operations preserving shellability under size constraints, and cyclic clique arrangements with star-clique attachments, where shellability depends critically on attachment sizes. These results generalize earlier findings on the existence of shellable independence complexes to the setting of higher independence complexes.
Key words and phrases:
chordal graph, hypergraph, shellable, linear resolution, $r$-independence complex
AMS Classification 2020: 05E45, 13F55, 05E40
1. Introduction
Let $G=(V(G),E(G))$ be a simple (no loops, no multiple edges), undirected graph with vertex set $V(G)=\{x_{1},x_{2},\dots,x_{n}\}$ and edge set $E(G)$. A subset $A\subseteq V(G)$ is called an independent set if no two vertices in $A$ are adjacent. Equivalently, the induced subgraph $G[A]$—defined as the subgraph of $G$ formed by the vertex set $A$ and all edges in $E(G)$ whose endpoints both lie in $A$—has no edges. The independence complex of $G$, denoted $\mathcal{I}(G)$, is the simplicial complex whose faces are precisely the independent sets of $G$. Formally, $\mathcal{I}(G)=\{A\subseteq V(G)\mid A\text{ is an independent set in }G\}.$ The Stanley-Reisner ideal of $\mathcal{I}(G)$ coincides with the edge ideal of $G$, denoted $I(G)$. This ideal is generated by the squarefree monomials corresponding to the edges of $G$: $I(G)=(x_{i}x_{j}\mid\{x_{i},x_{j}\}\in E(G))\subseteq\mathbb{K}[x_{1},\dots,x_{n}],$ where $\mathbb{K}$ is a field, [32]. The study of independence complexes has been a vibrant area of research, with significant attention devoted to their topological, combinatorial, and algebraic properties (see, e.g., [2, 23, 24, 29, 32, 34]). Central questions in this field include:
  1. (1)
    Classification and Characterization: A major focus is the classification of graph families whose independence complexes exhibit desirable properties, such as being vertex decomposable, shellable, or sequentially Cohen-Macaulay (see, e.g., [9, 17, 31, 32, 34, 35]). For simplicial complexes, the following hierarchy of implications holds:
[…]
    Note that these implications are strict.
  2. (2)
    Graph Modifications and Constructions: For an arbitrary graph $G$, a natural question arises: how can $G$ be modified—for instance, by attaching whiskers, adding cliques, or performing other graph operations—to produce a new graph whose independence complex satisfies the above properties? (see, e.g., [5, 8, 14, 16, 22, 31, 32, 34].)
This line of research lies at the intersection of combinatorics, commutative algebra, and topological combinatorics, offering deep insights into the connections between these fields.
In [26], Paolini and Salvetti introduced the $r$-independence complex of a graph $G$, generalizing the notion of the independence complex for any integer $r\geq 1$. A subset $A\subseteq V(G)$ is called $r$-independent if every connected component of the induced subgraph $G[A]$ has at most $r$ vertices. The $r$-independence complex of $G$, denoted $\mathcal{I}_{r}(G)$, is the simplicial complex with vertex set $V(G)$ and simplices given by all $r$-independent subsets of $G$. When $r=1$, this recovers the standard independence complex, i.e., $\mathcal{I}_{1}(G)=\mathcal{I}(G)$. These complexes have been extensively studied in various contexts (see [1, 3, 10, 11, 12, 14, 30]). In the study of $r$-independence complexes, two primary research directions—classification and characterization, and graph modifications and constructions—remain central. This paper focuses on these two directions, approaching them through the framework of chordal hypergraph techniques introduced by Woodroofe [35]. To proceed, we first recall the definitions of hypergraphs.
A hypergraph $\mathcal{H}$ is a pair $(V(\mathcal{H}),E(\mathcal{H}))$, where $V(\mathcal{H})$ is the vertex set, and $E(\mathcal{H})\subseteq 2^{V(\mathcal{H})}$ is the edge set, consisting of subsets of $V(\mathcal{H})$ called edges. A hypergraph is said to be simple if, for every $e\in E(\mathcal{H})$, there does not exist another edge $e^{\prime}\in E(\mathcal{H})$ such that $e^{\prime}\subsetneq e$. Throughout this paper, we assume all hypergraphs to be simple. Given a hypergraph $\mathcal{H}$ with vertex set $V(\mathcal{H})=\{x_{1},\dots,x_{n}\}$, we identify each vertex $x_{i}\in V(\mathcal{H})$ with a corresponding variable $x_{i}$ in the polynomial ring $\mathbb{K}[x_{1},\dots,x_{n}]$. The edge ideal of $\mathcal{H}$, denoted $I(\mathcal{H})$, is the squarefree monomial ideal in $\mathbb{K}[x_{1},\dots,x_{n}]$ generated by monomials associated with the edges of $\mathcal{H}$: $I(\mathcal{H})=\big{(}\prod\limits_{x_{i}\in e}x_{i}\mid e\in E(\mathcal{H})\big{)}.$ This definition generalizes the edge ideal of a graph and plays a fundamental role in the study of higher-degree squarefree monomial ideals, [19].
The study of chordal graphs—those containing no induced cycles of length greater than 3—has inspired multiple hypergraph generalizations, collectively termed chordal hypergraphs. These generalizations vary depending on which graph-theoretic properties are extended, leading to diverse definitions in the literature (see [6, 15, 19, 25, 31, 35]). Among these, Woodroofe’s [35] definition of $\operatorname{W}$-chordal hypergraphs [^1] The notation honors Russ Woodroofe’s foundational work. stands out as particularly influential. This class of hypergraphs exhibits deep algebraic properties, forming a proper higher-dimensional analog of classical chordal graph theory. For any $\operatorname{W}$-chordal hypergraph $\mathcal{H}$, two key structural results hold. First, the independence complex $\mathcal{I}(\mathcal{H})$ is shellable [35, Corollary 5.4]. Second, the Alexander dual of the independence complex of the complement hypergraph, $\mathcal{I}(\overline{\mathcal{H}})^{\vee}$, is vertex decomposable. This vertex decomposability implies that the edge ideal of the complement hypergraph (which is uniform) has linear quotients and thus admits a linear resolution [35, Corollary 6.10]. Together, these results provide a genuine partial extension of Fröberg’s theorem [18], which characterizes chordal graphs as those for which the complement’s edge ideal has a linear resolution. Note that these resolutions are field-independent.
In this paper, we investigate hypergraphs as natural generalizations of graphs. The construction of such hypergraphs has been extensively studied in the literature [1, 3, 10, 11, 14, 30]. Given a graph $G$ and an integer $r\geq 1$, we define the hypergraph $\mathcal{C}_{r}(G)$ associated with $G$ as follows: the vertex set of $\mathcal{C}_{r}(G)$ is $V(\mathcal{C}_{r}(G))=V(G)$, and the edge set consists of all subsets $S\subseteq V(G)$ of size $r+1$ such that the induced subgraph $G[S]$ is connected. Note that when $r=1$, the hypergraph $\mathcal{C}_{1}(G)$ coincides with the original graph $G$. The independence complex of this class of hypergraphs is particularly intriguing, as it aligns with the $r$-independence complex of $G$, i.e., $\mathcal{I}(\mathcal{C}_{r}(G))=\mathcal{I}_{r}(G)$ for all $r\geq 1$ [1, Proposition 2.6]. It is worth noting that even if $G$ is a chordal graph, $\mathcal{C}_{r}(G)$ may not be a $\operatorname{W}$-chordal hypergraph for $r\geq 2$ (see Proposition 4.1). However, identifying $\operatorname{W}$-chordal hypergraphs remains a non-trivial challenge, requiring sophisticated combinatorial techniques. Recently, [1, Theorem 2.7] claimed that if $G$ is a tree (a connected graph with no cycles), then $\mathcal{C}_{r}(G)$ is $\operatorname{W}$-chordal for all $r\geq 2$. However, we construct a class of counterexamples to this result using Proposition 4.1. Consequently, characterizing the trees $G$ and integers $r\geq 1$ for which $\mathcal{C}_{r}(G)$ is $\operatorname{W}$-chordal remains an open problem. To the best of our knowledge, no prior work has explored this direction.
In this paper, we study block graphs, a well-known family of graphs introduced by Harary [20]. A graph $G$ is a block graph if every block — a maximal connected subgraph that remains connected upon the removal of any single vertex — is a complete graph. Structurally, these graphs generalize trees by allowing cliques to be interconnected through shared vertices, forming a tree-like hierarchy of complete subgraphs. Block graphs exhibit two fundamental properties. First, they are closed under induced subgraphs: if $H$ is an induced subgraph of a block graph $G$, then $H$ itself is a block graph. Second, as established in [4], any two distinct vertices $v$ and $w$ in a connected block graph are connected by a unique shortest path, and this path is necessarily induced. This second property directly extends the analogous behavior of paths in trees, demonstrating how block graphs naturally generalize forest structures while introducing additional complexity through their clique-based construction.
Our main results establish deep connections between block graphs and the $\operatorname{W}$-chordality of their associated hypergraphs. Specifically, we prove the following theorems:
  1. (1)
    Theorem 4.3: Let $G=(V(G),E(G))$ be a graph with $|V(G)|=n\geq 3$. If $n-2\leq r\leq n-1$, then $\mathcal{C}_{r}(G)$ is a $\operatorname{W}$-chordal hypergraph.
  2. (2)
    Proposition 4.5: Let $C_{n}$ be a cycle of length $n\geq 4$. Then, $\mathcal{C}_{r}(C_{n})$ is $\operatorname{W}$-chordal if and only if $n-2\leq r\leq n-1$.
  3. (3)
    Theorem 4.7: Let $G$ be a forest. Then, for all $r\geq n-5$, the hypergraph $\mathcal{C}_{r}(G)$ is $\operatorname{W}$-chordal.
  4. (4)
    Theorem 5.1: If $G$ is a block graph, then the hypergraph $\mathcal{C}_{2}(G)$ is $\operatorname{W}$-chordal.
  5. (5)
    Theorem 5.2: If $G$ is a block graph with diameter $\operatorname{diam}(G)\leq 4$, then the hypergraph $\mathcal{C}_{r}(G)$ is $\operatorname{W}$-chordal for all $r\geq 2$.
  6. (6)
    Theorem 5.4: If $G$ is a forest with $\operatorname{diam}(G)\leq 5$, then the hypergraph $\mathcal{C}_{r}(G)$ is $\operatorname{W}$-chordal for all $r\geq 2$.
  7. (7)
    Theorem 5.6: For a block graph $G$, the hypergraph $\mathcal{C}_{3}(G)$ is $\operatorname{W}$-chordal if and only if $G$ is $\operatorname{T}_{3}$-free, i.e., $G$ does not contain specific substructures. As a consequence, if $G$ is a forest, then $\mathcal{C}_{3}(G)$ is $\operatorname{W}$-chordal (Corollary 5.7).
  8. (8)
    Theorem 5.12: If $G$ is a $\mathcal{T}_{2}$-graph (a generalization of lobster tree graphs; see Definition 5.10 for details), then the hypergraph $\mathcal{C}_{4}(G)$ is $\operatorname{W}$-chordal.
  9. (9)
    Corollary 5.13: For a tree $G$, $G$ is a lobster tree graph if and only if the hypergraph $\mathcal{C}_{4}(G)$ is $\operatorname{W}$-chordal.
  10. (10)
    Theorem 5.15: If $G$ is a $\mathcal{T}_{1}$-graph (a generalization of caterpillar tree graphs; see Definition 5.9 for details), then the hypergraph $\mathcal{C}_{r}(G)$ is $\operatorname{W}$-chordal for all $r\geq 5$.
As a consequence of these results, we establish one of the main contributions of this paper: the $r$-independence complex $\mathcal{I}_{r}(G)$ is shellable in each of the cases mentioned above. Furthermore, the Alexander dual of the independence complex of the complement hypergraph is vertex decomposable. This vertex decomposability implies that the edge ideal of the complement hypergraph $\overline{\mathcal{C}_{r}(G)}$ has linear quotients and thus admits a linear resolution in each of the above cases (Theorem 5.16).
The next focus of this paper is to investigate how various graph modifications—such as attaching whiskers, adding cliques, or performing other structural operations—affect the shellability of the higher independence complex. A foundational result by Villarreal [32] shows that for any graph $G$, the independence complex of its whiskered version $W(G)$—formed by attaching a pendant vertex (whisker) to each vertex of $G$—is Cohen–Macaulay. This was refined by Dochtermann and Engström [14] and independently by Woodroofe [34], who showed that $\mathcal{I}(W(G))$ is vertex decomposable. Building on this, Cook and Nagel [8] introduced the vertex clique-whiskered graph $G^{\pi}$ and proved that $\mathcal{I}(G^{\pi})$ is both unmixed and vertex decomposable. Subsequently, Biermann et al. [5] provided sufficient conditions on a subset $S\subseteq V(G)$ such that $\mathcal{I}(G\cup W(S))$ is vertex decomposable, where only the vertices in $S$ are whiskered. In a related direction, Hibi et al. [22] showed that attaching a complete graph to each vertex of $G$ results in a graph whose independence complex remains unmixed and vertex decomposable.
Motivated by these developments, we study the shellability of $r$-independence complexes $\mathcal{I}_{r}(G)$ for a family of graphs constructed via star-clique attachments. Given a graph $H$ and a vertex cover $S\subseteq V(H)$, we define, for a fixed integer $r\geq 1$, the graph $\operatorname{CCG}(H,S,r)$ as the graph obtained by attaching a star-clique graph $\operatorname{SC}(x)$ with $|\operatorname{SC}(x)|\geq r+1$ to each vertex $x\in S$. Our main result establishes the following:
Theorem 6.1. Let $G=\operatorname{CCG}(H,S,r)$, where $r\geq 1$. Then $\mathcal{I}_{r}(G)$ is shellable.
This result is further extended to a broader class of chordal graphs. In particular, we show that if $H$ is a chordal graph and $G=\mathrm{CCG}(H,V(H),t)$ for some integer $t\geq 1$, then $\mathcal{I}_{r}(G)$ is shellable whenever $r\leq 2t+1$ (Theorem 6.2). Moreover, we present a counterexample demonstrating that this bound is sharp: the conclusion fails when $r>2t+1$ (Example 6.3).
We also generalize the clique whiskering construction introduced by Cook and Nagel [8]. Given a graph $G$, a clique vertex-partition $\pi=\{W_{1},\ldots,W_{k}\}$, and an integer $r\geq 1$, the $r$-clique whiskering of $G$ with respect to $\pi$, denoted $G^{\pi}_{r}$, is formed by adding $t_{i}\geq r$ new vertices to each clique $W_{i}$ and making the union $W_{i}\cup\{x_{i,1},\ldots,x_{i,t_{i}}\}$ a clique. We prove the following:
Theorem 6.4. Let $G$ be a graph, and let $\pi=\{W_{1},\ldots,W_{k}\}$ be a clique vertex-partition of $G$. Then $\mathcal{I}_{r}(G^{\pi}_{r})$ is shellable.
Finally, we introduce the class of clique cycle graphs, which generalize ordinary cycles by replacing edges with cliques. An $n$-clique cycle $\mathrm{CC}(B_{1},\ldots,B_{n})$ is defined as a cyclic arrangement of $n\geq 3$ cliques where each consecutive pair shares exactly one vertex. We consider graphs $G(r)$ constructed from an $n$-clique cycle by attaching a star-clique graph $\operatorname{SC}(x)$ to each vertex in a subset $S$ of the connecting vertices, requiring that $|V(\operatorname{SC}(x))|\geq r+1$ for at least one $x\in S$. Our final main result is:
Theorem 6.5. Let $G=G(r)$ for some $r\geq 1$. Then $\mathcal{I}_{r}(G)$ is shellable.
To conclude, we provide an example showing that the conclusions of Theorems 6.1, 6.4, and 6.5 may fail for $\operatorname{CCG}(H,S,t)$, $G^{\pi}_{t}$, or $G(t)$ when $t<r$ (Example 6.6).
Our paper is structured as follows. In Section 2, we introduce the necessary terminology and foundational results that underpin our work. Section 3 presents key technical lemmas and propositions essential for establishing our main results. In Sections 4 and 5, we investigate the $\operatorname{W}$-chordality of hypergraphs $\mathcal{C}_{r}(G)$, providing sufficient conditions and characterizations for various classes of graphs, including cycles and block graphs. Finally, in Section 6, we examine the shellability of $r$-independence complexes in graph constructions.
2. Notation and Preliminaries
In this section, we establish the basic definitions and notation required for the main results. Let $R=\mathbb{K}[x_{1},\dots,x_{n}]$ be a polynomial ring over a field $\mathbb{K}$, and let $I\subseteq R$ be a homogeneous ideal. The minimal graded free resolution of $I$ is an exact sequence
$0\to\bigoplus_{j\in\mathbb{N}}R(-j)^{\beta_{p,j}(I)}\to\cdots\to\bigoplus_{j\in\mathbb{N}}R(-j)^{\beta_{0,j}(I)}\to I\to 0,$
where $p\leq n$ is the projective dimension of $I$, and $R(-j)$ denotes the twisted graded $R$-module with $R(-j)_{d}=R_{d-j}$. The integers $\beta_{i,j}(I)$, called the graded Betti numbers of $I$, encode the multiplicity of degree-$j$ generators in the $i$-th syzygy module of the resolution. The (Castelnuovo-Mumford) regularity of $I$ is defined as $\operatorname{reg}(I)=\max\{j-i\mid\beta_{i,j}(I)\neq 0\}.$ Intuitively, $\operatorname{reg}(I)$ captures the highest degree of a generator appearing in any syzygy module of $I$, serving as a key measure of the ideal’s algebraic complexity.
For a homogeneous ideal $I=(f_{1},\dots,f_{t})$, we highlight two fundamental properties. A homogeneous ideal $I$ has a linear resolution if all its minimal generators share the same degree $d$, i.e., $\deg f_{1}=\cdots=\deg f_{t}=d$, and the Betti numbers satisfy $\beta_{i,i+j}(I)=0$ for all $i\geq 1$ and $j\neq d$. Equivalently, the minimal free resolution takes the form
$0\to R(-d-p)^{\beta_{p,p+d}(I)}\to\cdots\to R(-d)^{\beta_{0,d}(I)}\to I\to 0$
The ideal $I$ has linear quotients if there exists an ordering $f_{\sigma(1)},\dots,f_{\sigma(t)}$ of its minimal generators such that for each $i=2,\dots,t$, the colon ideal $(f_{\sigma(1)},\dots,f_{\sigma(i-1)}):(f_{\sigma(i)})$ is generated by linear forms. These properties are closely related: if a homogeneous ideal has linear quotients and all its minimal generators share the same degree, then it admits a linear resolution [21, Proposition 8.2.1]. Moreover, if a homogeneous ideal with minimal generators of degree $d$ has a linear resolution, then $\operatorname{reg}(I)=d$.
A simplicial complex $\Delta$ on a vertex set $V=\{x_{1},\ldots,x_{n}\}$ is a collection of subsets of $V$ satisfying two conditions: (i) every singleton $\{x_{i}\}$ belongs to $\Delta$ for each $x_{i}\in V$, and (ii) $\Delta$ is closed under taking subsets; that is, if $F\in\Delta$, then every subset $F^{\prime}\subseteq F$ also belongs to $\Delta$. The elements of $\Delta$ are called faces, and the maximal faces under inclusion are called facets. Several important constructions are associated with a face $F\in\Delta$: the link of $F$ is defined as $\operatorname{link}_{\Delta}(F)=\{F^{\prime}\subseteq V\mid F^{\prime}\cap F=\emptyset\text{ and }F^{\prime}\cup F\in\Delta\}$; the deletion of $F$ is $\Delta\setminus F=\{H\in\Delta\mid H\cap F=\emptyset\}$; and the Alexander dual of $\Delta$ is the complex $\Delta^{\vee}=\{V\setminus F\mid F\notin\Delta\}$.
The concept of vertex decomposability for simplicial complexes was first introduced by Provan and Billera [27] in the pure case (where all maximal faces have the same cardinality) and later extended to arbitrary complexes by Björner and Wachs [7]. A simplicial complex $\Delta$ is vertex decomposable if it is either a simplex (i.e., it contains all subsets of its vertex set) or admits a shedding vertex $v\in V$ such that both the deletion $\Delta\setminus v$ and the link $\operatorname{link}_{\Delta}(v)$ are vertex decomposable, with the additional condition that no facet of $\operatorname{link}_{\Delta}(v)$ is a facet of $\Delta\setminus v$.
A related notion, shellability, requires that the facets of $\Delta$ can be ordered $F_{1},\ldots,F_{s}$ (a shelling order) such that for every $i<j$, there exists a vertex $v\in F_{j}\setminus F_{i}$ and an index $\ell<j$ satisfying $F_{j}\setminus F_{\ell}=\{v\}$.
A graded $R$-module $M$ is called sequentially Cohen–Macaulay over a field $\mathbb{K}$ if it admits a finite filtration $0=M_{0}\subset M_{1}\subset\cdots\subset M_{r}=M$ of graded submodules where each quotient $M_{i}/M_{i-1}$ is Cohen–Macaulay and the dimensions $\dim(M_{i}/M_{i-1})$ strictly increase. Stanley [29] showed that shellability of $\Delta$ implies the sequential Cohen–Macaulay property for its Stanley–Reisner ring. Furthermore, Herzog and Hibi [21, Proposition 8.2.5] established that $\Delta$ is shellable if and only if the Stanley–Reisner ideal of its Alexander dual has a linear quotient.
Let $G$ be a finite simple graph without isolated vertices. For a graph $G$, let $V(G)$ and $E(G)$ denote the vertex set and edge set of $G$, respectively. The degree of a vertex $x\in V(G)$, denoted by $\deg_{G}(x)$ (or simply $\deg(x)$), is the number of edges incident to $x$. A subgraph $H\subseteq G$ is called induced if for any two vertices $u,v\in V(H)$, the edge $\{u,v\}$ belongs to $E(H)$ if and only if $\{u,v\}\in E(G)$. For a subset $A\subseteq V(G)$, the induced subgraph of $G$ on $A$ is denoted by $G[A]$; that is, $G[A]$ is the subgraph of $G$ whose vertex set is $A$ and whose edge set consists of all edges of $G$ with both endpoints in $A$. For a subset $\{u_{1},\ldots,u_{r}\}\subseteq V(G)$, the neighborhood of $u_{1},\ldots,u_{r}$, denoted by $N_{G}(u_{1},\ldots,u_{r})$, is defined as $N_{G}(u_{1},\ldots,u_{r})=\{v\in V(G)\mid\{u_{i},v\}\in E(G)\text{ for some }1\leq i\leq r\}.$ The closed neighborhood of $u_{1},\ldots,u_{r}$, denoted by $N_{G}[u_{1},\ldots,u_{r}]$, is defined as $N_{G}[u_{1},\ldots,u_{r}]=N_{G}(u_{1},\ldots,u_{r})\cup\{u_{1},\ldots,u_{r}\}.$ For a subset $U\subseteq V(G)$, we denote by $G\setminus U$ the induced subgraph of $G$ on the vertex set $V(G)\setminus U$.
A path in $G$ is a sequence of distinct vertices $v_{0}v_{1}\cdots v_{k}$ such that $\{v_{i},v_{i+1}\}\in E(G)$ for all $10\leq i\leq k-$ 1. The length of a path is the number of edges it contains. A graph $G$ is called connected if for any two vertices $u,v\in V(G)$, there exists a path from $u$ to $v$. A connected component of $G$ is a maximal connected subgraph of $G$. The distance between two vertices $u$ and $v$ in $G$, denoted by $\operatorname{d}(u,v)$, is the length of the shortest path connecting $u$ and $v$. The diameter of $G$, denoted by $\operatorname{diam}(G)$, is the maximum distance between any pair of vertices in $G$. Formally, $\operatorname{diam}(G)=\max_{u,v\in V(G)}\operatorname{d}(u,v).$ A vertex $v\in V(G)$ is called a cut-vertex if the removal of $v$ from $G$ increases the number of connected components of $G$. A vertex $v\in V(G)$ is called a simplicial vertex if the subgraph induced by its neighbors is a complete graph (or clique); that is, for every pair of neighbors $u$ and $w$ of $v$, the edge $\{u,w\}\in E(G)$.
A hypergraph $\mathcal{H}$ is a pair $(V(\mathcal{H}),E(\mathcal{H}))$, where $V(\mathcal{H})$ is the vertex set and $E(\mathcal{H})\subseteq 2^{V(\mathcal{H})}$ is the edge set, consisting of subsets of $V(\mathcal{H})$ called the edges. A subset $S\subseteq V(\mathcal{H})$ is called independent if there is no $e\in E(\mathcal{H})$ such that $e\subseteq S$. The independence complex of a hypergraph $\mathcal{H}$, denoted $\mathcal{I}(\mathcal{H})$, is the simplicial complex whose simplices correspond to the independent subsets of $V(\mathcal{H})$. The deletion of the vertex $v$, denoted by $\mathcal{H}\setminus v$, is the hypergraph defined as:
$V(\mathcal{H}\setminus v)=V(\mathcal{H})\setminus\{v\},\quad E(\mathcal{H}\setminus v)=\{e\in E(\mathcal{H})\mid v\notin e\}.$
The contraction of the vertex $v$, denoted by $\mathcal{H}/v$, is the hypergraph defined as:
$V(\mathcal{H}/v)=V(\mathcal{H})\setminus\{v\},\quad E(\mathcal{H}/v)=\min\bigl{\{}e\setminus\{v\}\mid e\in E(\mathcal{H})\bigr{\}},$
where $\min(S)$ denotes the inclusion-wise minimal subsets of a collection of sets $S$. Contractions and deletions of distinct vertices can be performed sequentially, and it is well known that the resulting hypergraph does not depend on the order of these operations [28, Lemma 3].
A hypergraph $\mathcal{H}^{\prime}$ obtained from $\mathcal{H}$ by deleting a subset of vertices $V_{d}\subseteq V(\mathcal{H})$ and contracting another subset $V_{c}\subseteq V(\mathcal{H})$, with $V_{d}\cap V_{c}=\emptyset$, is called a minor of $\mathcal{H}$, denoted by $\mathcal{H}\setminus V_{d}/V_{c}$. If $V_{d}=\emptyset$, the hypergraph $\mathcal{H}^{\prime}$ is called a $c$-minor of $\mathcal{H}$. Furthermore, if a $c$-minor $\mathcal{H}^{\prime}$ contains no singleton edges, it is called a $c^{\prime}$-minor of $\mathcal{H}$.
The concept of $\operatorname{W}$-chordality, introduced by Woodroofe, generalizes the notion of chordality from graphs to hypergraphs and provides a framework for studying shellability and linear resolutions.
Definition 2.1 ([35, Definition 4.2 and Definition 4.3]).
  1. (1)
    A vertex $v\in V(\mathcal{H})$ in a hypergraph $\mathcal{H}$ is called a simplicial vertex if for any two edges $e_{1},e_{2}\in E(\mathcal{H})$ containing $v$, there exists an edge $e_{3}\in E(\mathcal{H})$ such that $e_{3}\subseteq(e_{1}\cup e_{2})\setminus\{v\}.$
  2. (2)
    A hypergraph $\mathcal{H}$ is said to be $\operatorname{W}$-chordal if every minor of $\mathcal{H}$ contains a simplicial vertex.
These definitions lay the foundation for analyzing the shellability and sequentially Cohen–Macaulay properties of independence complexes.
Theorem 2.2 ([35, Corollary 5.3 and Corollary 5.4]).
  1. (1)
    Let $\mathcal{H}$ be a hypergraph such that every contraction of $\mathcal{H}$ has a simplicial vertex. Then $\mathcal{I}(\mathcal{H})$ is shellable, and therefore the Stanley–Reisner ideal of $\mathcal{I}(\mathcal{H})$ is sequentially Cohen–Macaulay.
  2. (2)
    If $\mathcal{H}$ is a $\operatorname{W}$-chordal hypergraph, then $\mathcal{I}(\mathcal{H})$ is shellable, and therefore the Stanley–Reisner ideal of $\mathcal{I}(\mathcal{H})$ is sequentially Cohen–Macaulay.
For any undefined terminology and further basic definitions and properties, we refer the reader to [21] and [33].
3. Technical Preliminaries and Key Results
In this section, we establish the technical groundwork necessary for the proof of our main theorem. We present a series of key lemmas and propositions that serve as the foundation for our arguments
A collection $\mathcal{F}$ of graphs is called a hierarchy if, for every nonempty graph $G\in\mathcal{F}$, the graph $G\setminus u$ also belongs to $\mathcal{F}$ for every vertex $u\in V(G)$. The following proposition establishes a connection between the structural properties of a graph hierarchy and the $\operatorname{W}$-chordality of their associated hypergraphs. In particular, it provides a sufficient condition for $\mathcal{C}_{r}(G)$ to be $\operatorname{W}$-chordal, based on the existence of simplicial vertices in its minors.
Proposition 3.1.
Let $\mathcal{F}$ be a hierarchy of graphs and $r$ a positive integer. If every $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$ contains a simplicial vertex for any $G\in\mathcal{F}$, then $\mathcal{C}_{r}(G)$ is $\operatorname{W}$-chordal.
Proof.
We first show that every $c$-minor of $\mathcal{C}_{r}(G)$ contains a simplicial vertex for any $G\in\mathcal{F}$. Let $\mathcal{H}$ be a $c$-minor of $\mathcal{C}_{r}(G)$, i.e., $\mathcal{H}=\mathcal{C}_{r}(G)/V_{c}$ for some $V_{c}\subseteq V(G)$. We consider two cases based on the presence of singleton edges in $\mathcal{H}$: Case 1: If $\mathcal{H}$ has no singleton edges, then $\mathcal{H}$ is a $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$. By assumption, $\mathcal{H}$ contains a simplicial vertex. Case 2: Suppose $\mathcal{H}$ contains singleton edges. Let $S$ denote the set of all singleton edges in $\mathcal{H}$. Consider the hypergraph $\mathcal{H}\setminus S=(\mathcal{C}_{r}(G)/V_{c})\setminus S.$ By [28, Lemma 3], we have
$\mathcal{H}\setminus S=(\mathcal{C}_{r}(G)\setminus S)/V_{c}.$
Additionally, for any vertex $x\in V(G)$ and $r\geq 1$, we observe that $\mathcal{C}_{r}(G)\setminus x=\mathcal{C}_{r}(G\setminus x).$ Extending this identity to the set $S$, we obtain
$\mathcal{H}\setminus S=\mathcal{C}_{r}(G\setminus S)/V_{c}.$
Since $G\setminus S\in\mathcal{F}$ (as $\mathcal{F}$ is a hierarchy), and $\mathcal{H}\setminus S$ is a $c^{\prime}$-minor of $\mathcal{C}_{r}(G\setminus S)$, the hypothesis implies that $\mathcal{H}\setminus S$ contains a simplicial vertex. Thus, $\mathcal{H}$ must also contain a simplicial vertex.
In both cases, we conclude that every $c$-minor of $\mathcal{C}_{r}(G)$ contains a simplicial vertex. Since any minor can be obtained through deletions and contractions, and deletions correspond to graphs in the hierarchy by definition, it follows that all minors of $\mathcal{C}_{r}(G)$ possess a simplicial vertex. Hence, $\mathcal{C}_{r}(G)$ is $\operatorname{W}$-chordal for every $G\in\mathcal{F}$. ∎
In [10, Theorem 4.3], Dao and Schweig proved that a vertex $v$ of a graph $G$ is simplicial if and only if it remains simplicial in $\mathcal{C}_{r}(G)$ for all $r\geq 1$. However, this equivalence does not hold for general graphs. For example, consider the cycle graph $C_{4}$. While no vertex in $C_{4}$ is simplicial, every vertex in $\mathcal{C}_{r}(C_{4})$ becomes simplicial for $r=2$ (or $r=3$ under their construction). This demonstrates that the claimed equivalence fails in general.
Motivated by this observation, we examine the persistence of simplicial vertices in minors of the hypergraph $\mathcal{C}_{r}(G)$.
Lemma 3.2.
Let $G$ be a graph, and let $\mathcal{H}$ be a minor of the hypergraph $\mathcal{C}_{r}(G)$ for some $r\geq 1$. If $x$ is a simplicial vertex of $G$ and $x\in V(\mathcal{H})$, then $x$ remains a simplicial vertex in $\mathcal{H}$.
Proof.
Suppose first that $x$ belongs to exactly one edge of $\mathcal{H}$; then $x$ is simplicial by definition. Now consider the case where $x$ belongs to two distinct edges $e_{1},e_{2}\in E(\mathcal{H})$, with $e_{1}\neq e_{2}$. By construction of the minor, there exist edges $f_{1},f_{2}\in E(\mathcal{C}_{r}(G))$ such that $e_{i}=f_{i}\cap V(\mathcal{H})\quad\text{for }i=1,2.$ Since $x$ is simplicial in $G$, the induced subgraph of $G$ on $(f_{1}\cup f_{2})\setminus\{x\}$ is connected. Moreover, the fact that $f_{1}\neq f_{2}$ implies that $|f_{1}\cup f_{2}|\geq r+2,$ and hence, $|(f_{1}\cup f_{2})\setminus\{x\}|\geq r+1.$ Select a subset $f\subseteq(f_{1}\cup f_{2})\setminus\{x\}$ of size $r+1$ such that the induced subgraph on $f$ is connected (which is possible since the entire set is connected and of size at least $r+1$). Define $e_{3}:=f\cap V(\mathcal{H}).$ Clearly, $e_{3}\subseteq e_{1}\cup e_{2}\setminus\{v\}$. If $e_{3}\in E(\mathcal{H})$, we are done. Otherwise, there exists $e_{3}^{\prime}\in E(\mathcal{H})$ such that $e_{3}^{\prime}\subseteq e_{3}\subseteq e_{1}\cup e_{2}\setminus\{v\}$. This proves that $v$ is a simplicial vertex in $\mathcal{H}$. ∎
A graph is called a star-clique graph, denoted $\operatorname{SC}(x)$, if it is formed by attaching a collection of cliques $B_{1},\ldots,B_{t}$ to a central vertex $x$, where each $B_{i}$ is a complete graph (clique) that may vary in size.
Corollary 3.3.
If $G$ is a star-clique graph, then $\mathcal{C}_{r}(G)$ is a $\operatorname{W}$-chordal hypergraph for all $r\geq 1$.
Proof.
Let $V_{1}$ be the set of all simplicial vertices of $G$. Consider a minor $\mathcal{H}$ of the hypergraph $\mathcal{C}_{r}(G)$. If $V_{1}\cap V(\mathcal{H})\neq\emptyset$, then by Lemma 3.2, $\mathcal{H}$ contains a simplicial vertex. If $V_{1}\cap V(\mathcal{H})=\emptyset$, then all simplicial vertices of $G$ are excluded from $\mathcal{H}$, which can only happen if $\mathcal{H}$ has exactly one vertex. In that case, $\mathcal{H}$ trivially contains a simplicial vertex. Therefore, every minor $\mathcal{H}$ of $\mathcal{C}_{r}(G)$ has a simplicial vertex, which implies that $\mathcal{C}_{r}(G)$ is $\operatorname{W}$-chordal. ∎
The following observation highlights a key property of chordal graphs, leveraging the structure of simplicial vertices and their iterative removal.
Observation 3.4.
Let $G$ be a chordal graph with $|V(G)|\geq 2$. Define $V_{1}$ as the set of all simplicial vertices of $G$. In [13], Dirac provided a fundamental characterization of chordal graphs in terms of simplicial vertices: every chordal graph contains at least one simplicial vertex, and any chordal graph that is not complete contains at least two non-adjacent simplicial vertices. Therefore, we have $|V_{1}|\geq 2$.
For $i\geq 2$, let $V_{i}$ denote the set of simplicial vertices in the subgraph obtained by successively removing the sets $V_{1},V_{2},\dots,V_{i-1}$ from $G$, i.e., $G\setminus\bigcup_{j=1}^{i-1}V_{j}$. This process continues, and we define $V_{\kappa}$ as the set of simplicial vertices in the subgraph $G\setminus\bigcup_{j=1}^{\kappa-1}V_{j}$, where the graph $G\setminus\bigcup_{j=1}^{\kappa}V_{j}$ becomes empty for some $\kappa\geq 1$.
Since $G$ is chordal, and the subgraph $G\setminus\bigcup_{j=1}^{t}V_{j}$ remains chordal for all $1\leq t\leq\kappa$, it follows that each set $V_{i}$ is non-empty. Furthermore, we have $|V_{i}|\geq 2$ for all $1\leq i<\kappa$, and $|V_{\kappa}|\geq 1$. Additionally, the sets $V_{i}$ are pairwise disjoint, meaning $V_{i}\cap V_{j}=\emptyset$ whenever $i\neq j$.
The following observation explores the structure of block graphs, a subclass of chordal graphs. Using the framework established in Observation 3.4, we analyze the cliques associated with vertices in $V_{2}$ and derive a key property regarding their structure and size.
Observation 3.5.
Let $G$ be a block graph. Since every block graph is chordal, we apply the framework introduced in Observation 3.4. For each vertex $v\in V_{2}$, there exists a maximum-sized clique in $G$ that contains $v$, such that all other vertices in this clique belong to $V_{1}$. Denote these cliques by $C_{1},C_{2},\ldots,C_{m}$, where $m\geq 1$. Define
$S_{v}=\bigcup_{i=1}^{m}\left(C_{i}\setminus\{v\}\right).$
As a consequence, we observe that $|S_{v}|\geq 1$.
We now define an $n$-clique path graph, denoted $\operatorname{CP}(B_{1},\ldots,B_{n})$, as a graph consisting of a sequence of $n$ cliques $B_{1},B_{2},\ldots,B_{n}$, arranged linearly, with $n\geq 1$. Each consecutive pair of cliques $B_{i}$ and $B_{i+1}$ shares exactly one vertex, denoted $x_{i}$. The vertices $x_{1},\ldots,x_{n-1}$ are referred to as the connecting vertices of the $n$-clique path graph.
[…]
$\operatorname{CP}(B_{1},B_{2},B_{3},B_{4},B_{5})$
For example, the graph shown left is a 5-clique path graph $\operatorname{CP}(B_{1},B_{2},B_{3},B_{4},B_{5})$, where $B_{1}$ is the clique on 4 vertices, $B_{2}$ is the clique on 3 vertices, $B_{3}$ is the clique on 2 vertices, $B_{4}$ is the clique on 6 vertices, and $B_{5}$ is the clique on 3 vertices. The vertices $x_{1},x_{2},x_{3},x_{4}$ are the connecting vertices, which link each consecutive pair of cliques in this path.
The following setup introduces key notation and assumptions that will be used in subsequent results. Using Observation 3.4, we define a parameter $\ell$ that plays a crucial role in the analysis.
Set-up 3.6.
Let $G$ be a connected block graph. Then there exists a maximum-sized $n$-clique path, denoted $\operatorname{CP}(B_{1},\ldots,B_{n})$, as an induced subgraph with connecting vertices $x_{1},x_{2},\ldots,x_{n-1}$, where $n\geq 1$. Let $\mathcal{H}$ be a $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$ for some $r\geq 1$, and assume that $|E(\mathcal{H})|\geq 3$. Using Observation 3.4 and its associated notation, we define the parameter $\ell$ as:
$\ell=\min\left\{i\mid V_{i}\cap V(\mathcal{H})\neq\emptyset\right\}.$
Lemma 3.7.
Assume the notation from Set-Up 3.6, and suppose $\ell\neq 1$ and $r<n$. If $x_{0},\ldots,x_{r-2}\notin V(\mathcal{H})$, where $x_{0}\in V(B_{1})\cap V_{1}$, then $x_{r-1}\in V(\mathcal{H})$. Furthermore, if $y\in N_{G}(x_{r-1})\setminus\{x_{r-2}\}$, then $y\in V(\mathcal{H})$.
Proof.
Suppose, for contradiction, that $x_{r-1}\notin V(\mathcal{H})$. Choose a vertex $c\in V(\mathcal{H})$ such that
$\mathrm{d}(c,x_{r-1})=\min\{\mathrm{d}(z,x_{r-1})\mid z\in V(\mathcal{H})\}.$
Set $\mathrm{d}(c,x_{r-1})=t$. Then, there exists a shortest path $(x_{r-1}=y_{0})y_{1}\cdots y_{t-1}(y_{t}=c)$ from $x_{r-1}$ to $c$. Since $c$ is the closest vertex in $V(\mathcal{H})$, all intermediate vertices $y_{1},\ldots,y_{t-1}$ must lie outside $V(\mathcal{H})$. Define $S=\{x_{0},\ldots,x_{r-1},y_{1},\ldots,y_{t}\}.$ Clearly, $|S|\geq r+1$ and $S$ is connected. Moreover, $S\cap V(\mathcal{H})=\{c\}$. Since $c\in V(\mathcal{H})$, any $f\in E(\mathcal{C}_{r}(G))$ such that $f\subseteq S$ and $c\in f$ would imply $f\cap V(\mathcal{H})=\{c\}\in E(\mathcal{H})$, contradicting the assumption that $\mathcal{H}$ is a $c^{\prime}$-minor (i.e., no edge of $\mathcal{C}_{r}(G)$ intersects $V(\mathcal{H})$ in exactly one vertex). Hence, we must have $x_{r-1}\in V(\mathcal{H})$.
Now suppose $y\in N_{G}(x_{r-1})\setminus\{x_{r-2}\}$, and assume, for contradiction, that $y\notin V(\mathcal{H})$. Then the set $T=\{x_{0},\ldots,x_{r-1},y\}$ is connected. Since $x_{r-1}\in V(\mathcal{H})$ and all other elements of $T$ lie outside $V(\mathcal{H})$, we have $T\cap V(\mathcal{H})=\{x_{r-1}\}.$ Thus, $T$ defines an edge in $\mathcal{C}_{r}(G)$ whose intersection with $V(\mathcal{H})$ is exactly $\{x_{r-1}\}\in E(\mathcal{H})$, again contradicting the $c^{\prime}$-minor condition. Therefore, $y\in V(\mathcal{H})$. ∎
Lemma 3.8.
Assume the setup in 3.6 and the notation introduced in Observation 3.5. Let $\ell=2$, and suppose that $v_{k}\in e_{1}\cap e_{2}$, where $e_{1},e_{2}\in E(\mathcal{H})$ and $e_{1}\neq e_{2}$. Then, there exist edges $f_{1},f_{2}\in E(\mathcal{C}_{r}(G))$ such that the following conditions hold:
$S_{v_{k}}\subseteq f_{i},\quad f_{i}\cap V(\mathcal{H})=e_{i}\quad\text{for }i=1,2,$
and
$f_{1}\cup f_{2}\setminus\left(S_{v_{k}}\cup\{v_{k}\}\right)$
is connected.
Proof.
Let $S_{v_{k}}=S$. We first claim that there exists $f_{1}\in E(\mathcal{C}_{r}(G))$ such that:
$S\subseteq f_{1}\quad\text{and}\quad f_{1}\cap V(\mathcal{H})=e_{1}.$
Since $\mathcal{H}$ is a $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$, there exists an edge $f^{\prime}\in E(\mathcal{C}_{r}(G))$ such that $f^{\prime}\cap V(\mathcal{H})=e_{1}$. If $S\subseteq f^{\prime}$, the claim is true. Otherwise, since every vertex of $S$ is connected to $v_{k}$, and $v_{k}\in f^{\prime}$, the set $S\cup f^{\prime}$ is connected, and we have $|S\cup f^{\prime}|\geq r+1$. We can then choose a connected subset $f_{1}$ of $S\cup f^{\prime}$ such that $S\subseteq f_{1}$ and $|f_{1}|=r+1$. Consequently, we have: $f_{1}\cap V(\mathcal{H})\subseteq f^{\prime}\cap V(\mathcal{H})=e_{1}.$ If $f_{1}\cap V(\mathcal{H})\subset e_{1}$, this would contradict the assumption that $e_{1}\in E(\mathcal{H})$, since $f_{1}\cap V(\mathcal{H})$ would then be a proper subset of $e_{1}$. Therefore, it must be that $f_{1}\cap V(\mathcal{H})=e_{1}$.
Similarly, there exists $f_{2}\in E(\mathcal{C}_{r}(G))$ such that $S\subseteq f_{2}\quad\text{and}\quad f_{2}\cap V(\mathcal{H})=e_{2}.$
Next, we claim that there exists $z\in f_{1}\setminus(S\cup\{v_{k}\})$ such that $z\notin V_{1}$ and $z\in N_{G}(v_{k})$. There must exist a vertex $t\in N_{G}(v_{k})$ such that $t\notin V_{1}$, because if every vertex in $N_{G}(v_{k})$ were in $V_{1}$, then $v_{k}$ would be in $V_{1}$, contradicting the assumption that $v_{k}\in V_{2}$. Since $e_{1}\neq e_{2}$, there exists $x_{1}^{\prime}\in e_{1}\setminus e_{2}$, and hence $x_{1}^{\prime}\in f_{1}$. Consider the shortest path $(v_{k}=a_{0})a_{1}\cdots a_{n-1}(a_{n}=x_{1}^{\prime})$ in $G$, where $a_{0},\dots,a_{n}\in f_{1}$. Then, $a_{1}\in N_{G}(v_{k})$, and if $a_{1}\in V_{1}$, $a_{1}$ would be a simplicial vertex, implying the path $v_{k}a_{2}\cdots a_{n-1}(a_{n}=x_{1}^{\prime})$ would be a shorter path, leading to a contradiction. Therefore, $a_{1}\notin V_{1}$, and hence $a_{1}\notin S\cup\{v_{k}\}$. This proves the claim.
Similarly, there exists $z\in f_{2}\setminus(S\cup\{v_{k}\})$ such that $z\notin V_{1}$ and $z\in N_{G}(v_{k})$.
Finally, we claim that $f_{1}\cup f_{2}\setminus(S\cup\{v_{k}\})$ is connected. By the previous argument, there exist vertices $z_{1}\in f_{1}\setminus(S\cup\{v_{k}\})$ and $z_{2}\in f_{2}\setminus(S\cup\{v_{k}\})$, both of which are not in $V_{1}$ and are in $N_{G}(v_{k})$. If $z_{1}=z_{2}$, the claim follows immediately. Otherwise, if $z_{1}\neq z_{2}$, there must be an edge between $z_{1}$ and $z_{2}$ because $v_{k}\in V_{2}$, which ensures that $z_{1}$ and $z_{2}$ are connected. Therefore, $f_{1}\cup f_{2}\setminus(S\cup\{v_{k}\})$ is connected, and the proof is complete. ∎
Notation 3.9.
Assume the setup in 3.6 and the notation introduced in Observation 3.5. Let $\ell=2$, and suppose that $v_{k}\in e_{1}\cap e_{2}$, where $e_{1},e_{2}\in E(\mathcal{H})$ and $e_{1}\neq e_{2}$. By Lemma 3.8, there exist edges $f_{1},f_{2}\in E(\mathcal{C}_{r}(G))$ such that:
$S_{v_{k}}\subseteq f_{i},\quad f_{i}\cap V(\mathcal{H})=e_{i},\quad\text{for }i=1,2,$
and the set
$f_{1}\cup f_{2}\setminus(S_{v_{k}}\cup\{v_{k}\})$
induces a connected subgraph. Next, fix an element $x_{1}^{\prime}\in e_{1}\setminus e_{2}$. For each $1\leq i\leq\gamma$, where $\gamma\geq 1$, define a set
$C_{i}=\{c_{i,1},\ldots,c_{i,q_{i}}\}\subseteq V(G),$
satisfying the following conditions:
$\left.\begin{array}[]{ll}\text{(i)}&C_{i}\text{ is connected},\\ \text{(ii)}&c_{i,j}\in N_{G}(x_{1}^{\prime})\text{ for some }j\in\{1,\ldots,q_{i}\},\\ \text{(iii)}&C_{i}\cap V(\mathcal{H})=\emptyset,\\ \text{(iv)}&\operatorname{d}(c_{i,j},v_{k})>\operatorname{d}(v_{k},x_{1}^{\prime})\text{ for all }j\in\{1,\ldots,q_{i}\}.\end{array}\right\}$ (3.1)
Finally, define
$C=\bigcup_{i=1}^{\gamma}C_{i}.$
Lemma 3.10.
Using the notation from Notation 3.9, we have:
  1. (1)
    If $|f_{1}\cup f_{2}\setminus(S_{v_{k}}\cup\{v_{k}\})|\geq r+1$, then $v_{k}$ is a simplicial vertex of $\mathcal{H}$.
  2. (2)
    If $|f_{1}\cup f_{2}\setminus(S_{v_{k}}\cup\{v_{k}\})|\leq r$ and $|C|\geq|S_{v_{k}}|$, then $v_{k}$ is a simplicial vertex of $\mathcal{H}$.
Proof.
(1) Suppose that $\left|(f_{1}\cup f_{2})\setminus(S_{v_{k}}\cup\{v_{k}\})\right|\geq r+1$. Then, there exists a connected subset $f\subseteq(f_{1}\cup f_{2})\setminus(S_{v_{k}}\cup\{v_{k}\})$ with $|f|=r+1$. Let $e_{3}=f\cap V(\mathcal{H})$; since $f$ avoids $\{v_{k}\}$, we have $e_{3}\subseteq(e_{1}\cup e_{2})\setminus\{v_{k}\}$. If $e_{3}$ is itself an edge in $E(\mathcal{H})$, then $v_{k}$ must be a simplicial vertex. Otherwise, by the properties of the hypergraph $\mathcal{H}$, there exists a subset $e_{3}^{\prime}\in E(\mathcal{H})$ such that $e_{3}^{\prime}\subseteq e_{3}$, which again implies that $v_{k}$ is simplicial. In either case, the conclusion follows.
(2) Now assume $|A|\leq r$, where $A=f_{1}\cup f_{2}\setminus(S_{v_{k}}\cup\{v_{k}\})$. Let $S=S_{v_{k}}$ and define $D^{\prime}=f_{2}\setminus(S\cup\{v_{k}\})$. We claim that $D^{\prime}$ is a connected set. Let $\alpha,\beta\in D^{\prime}$. If $\alpha$ and $\beta$ are not adjacent, since $\alpha,\beta\in f_{2}$, there exists a shortest path $\alpha\ y_{1}\cdots y_{p}\ \beta$ within $f_{2}$, where each $y_{i}\in f_{2}$ for $1\leq i\leq p$. Since $y_{i}$ is not a simplicial vertex of $G$ for all $1\leq i\leq p$, we have $y_{i}\notin S\cup\{v_{k}\}$, ensuring that $D^{\prime}$ remains connected.
Next, we claim that $D^{\prime}\cap C=\emptyset$. Suppose, for contradiction, that there exists $c_{i,j}\in D^{\prime}\cap C$, where $c_{i,j}\in C_{i}$ for some $1\leq i\leq\gamma$, $1\leq j\leq q_{i}$. Since $v_{k},x_{1}^{\prime}\in f_{1}$, there exists a shortest path $(v_{k}=z_{0})z_{1}\cdots z_{t-1}(z_{t}=x_{1}^{\prime})$ in $G$, where $z_{i}\in f_{1}$ for all $0\leq i\leq t$. Let $B$ be the maximal clique containing $z_{t-1}$ and $x_{1}^{\prime}$. Similarly, since $v_{k},c_{i,j}\in f_{2}$, there exists a shortest path $P(v_{k},c_{i,j})$ within $f_{2}$. Since $c_{i,j}\in C_{i}$ and conditions (i) and (ii) from (3.1) hold, there exists a shortest path $x_{1}^{\prime}a_{1}\cdots a_{m}c_{i,j}$, where $a_{p}\in C_{i}$ for all $1\leq p\leq m$. We observe that $a_{1}\notin V(B)$, since otherwise, $\operatorname{d}(v_{k},a_{1})\leq\operatorname{d}(v_{k},x_{1}^{\prime})$, contradicting condition (iv) from (3.1). Since $x_{1}^{\prime}$ is a cut vertex, $a_{1}$ and $z_{t-1}$ belong to different components of $G\setminus\{x_{1}^{\prime}\}$. However, the path $z_{t-1}\cdots v_{k}\ P(v_{k},c_{i,j})\ a_{m}\cdots a_{1}$ remains in $G\setminus\{x_{1}^{\prime}\}$, which is a contradiction. Therefore, $D^{\prime}\cap C=\emptyset$.
Since $D^{\prime}\cap\{x_{1}^{\prime}\}=\emptyset$ (because $x_{1}^{\prime}\in e_{2}$), we have $|f_{2}|=r+1=|D^{\prime}|+|S|+1.$ Now, consider $D^{\prime}\cup R\cup C$, where $R=\{z_{1},\ldots,z_{t-1},z_{t}\}$. We can show that $D^{\prime}\cup R\cup C$ is connected. By computing its size, we obtain $|D^{\prime}\cup R\cup C|\geq r+1.$ Select a connected subset $f\subseteq D^{\prime}\cup R\cup C$ with $|f|=r+1$, and define $e=f\cap V(\mathcal{H}).$ This ensures that $e\subseteq e_{1}\cup e_{2}\setminus\{v_{k}\}$. If $e\in E(\mathcal{H})$, then $v_{k}$ is a simplicial vertex. Otherwise, there exists $e_{s}\in E(\mathcal{H})$ such that $e_{s}\subseteq e$, proving that $v_{k}$ is a simplicial vertex. ∎
4. Conditions on $r$ for $\operatorname{W}$-Chordality in $\mathcal{C}_{r}(G)$
This section investigates the $\operatorname{W}$-chordality of the hypergraphs $\mathcal{C}_{r}(G)$. For each $r\geq 4$, we begin by constructing a family of graphs $G_{t}$ by attaching three connected graphs $H_{1}$, $H_{2}$, and $H_{3}$, each of size at least $t$, to a base graph $G$. Proposition 4.1 shows that if $t\geq r-3$, then the hypergraph $\mathcal{C}_{r}(G_{t})$ is not $\operatorname{W}$-chordal. We then establish a sufficient condition for $\operatorname{W}$-chordality in Theorem 4.3: for any graph $G$ with $n\geq 3$ vertices, the hypergraph $\mathcal{C}_{r}(G)$ is $\operatorname{W}$-chordal whenever $n-2\leq r\leq n-1$. Proposition 4.5 demonstrates that this condition is sharp for cycle graphs $C_{n}$, where $\mathcal{C}_{r}(C_{n})$ is $\operatorname{W}$-chordal if and only if $n-2\leq r\leq n-1$. Finally, we extend the lower bound on $r$ to the class of forests in Theorem 4.7.
For each $r\geq 4$, we construct a family of graphs $G_{t}$ such that the hypergraph $\mathcal{C}_{r}(G_{t})$ is not $\operatorname{W}$-chordal. The construction proceeds as follows.
Construction 1.
Let $H_{1}$, $H_{2}$, and $H_{3}$ be connected graphs with pairwise disjoint vertex sets, each satisfying $|V(H_{i})|\geq t$ for $1\leq i\leq 3$. Define a base graph $G$ with vertex set $V(G)=\{v_{1},v_{2},v_{3},v_{4},v_{5},v_{6},v_{7}\}$ and edge set
$E(G)=\{\{v_{1},v_{2}\},\{v_{2},v_{3}\},\{v_{1},v_{4}\},\{v_{4},v_{5}\},\{v_{1},v_{6}\},\{v_{6},v_{7}\}\}.$
[…]
The graph $G_{t}$
The graph $G_{t}$ is formed by taking the disjoint union of $G$, $H_{1}$, $H_{2}$, and $H_{3}$, and then attaching each $H_{i}$ to $G$ via a single edge. Specifically, let $u_{1}\in V(H_{1})$, $u_{2}\in V(H_{2})$, and $u_{3}\in V(H_{3})$ be arbitrary vertices. Then the vertex and edge sets of $G_{t}$ are given by $V(G_{t})=V(G)\cup V(H_{1})\cup V(H_{2})\cup V(H_{3}),$ $E(G_{t})=E(G)\cup E(H_{1})\cup E(H_{2})\cup E(H_{3})\cup\{\{v_{3},u_{1}\},\{v_{5},u_{2}\},\{v_{7},u_{3}\}\}.$
Proposition 4.1.
If $t\geq r-3$ for some $r\geq 4$, then the hypergraph $\mathcal{C}_{r}(G_{t})$ is not $\operatorname{W}$-chordal.
Proof.
For each $1\leq i\leq 3$, choose a subset $S_{i}\subseteq V(H_{i})$ with $|S_{i}|=r-3$ such that the following subsets of $V(G_{t})$ belong to $E(\mathcal{C}_{r}(G_{t}))$:
$S_{1}\cup\{v_{1},v_{2},v_{3},v_{4}\},\quad S_{1}\cup\{v_{1},v_{2},v_{3},v_{6}\},$
$S_{2}\cup\{v_{5},v_{4},v_{1},v_{6}\},\quad S_{2}\cup\{v_{5},v_{4},v_{1},v_{2}\},$
$S_{3}\cup\{v_{7},v_{6},v_{1},v_{4}\},\quad S_{3}\cup\{v_{7},v_{6},v_{1},v_{2}\}.$
Let $S^{\prime}_{i}=V(H_{i})\setminus S_{i}$, and define $S=S^{\prime}_{1}\cup S^{\prime}_{2}\cup S^{\prime}_{3}$. Let $\mathcal{H}$ be the hypergraph obtained from $\mathcal{C}_{r}(G_{t})$ by deleting all vertices in $S$ and contracting the vertices in $S_{1}\cup S_{2}\cup S_{3}$. Since vertex deletion and contraction behave compatibly with $\mathcal{C}_{r}(G)$ (i.e., $\mathcal{C}_{r}(G)\setminus x=\mathcal{C}_{r}(G\setminus x)$ for any vertex $x\in V(G)$), the hypergraph $\mathcal{H}$ has: $V(\mathcal{H})=\{v_{1},v_{2},v_{3},v_{4},v_{5},v_{6},v_{7}\},$ and its edge set is: $E(\mathcal{H})=$$\bigl{\{}\{v_{1},v_{2},v_{3},v_{4}\},\,$$\{v_{1},v_{2},v_{3},v_{6}\},\,$$\{v_{5},v_{4},v_{1},v_{6}\},\,$$\{v_{5},v_{4},v_{1},v_{2}\},\,$$\{v_{7},v_{6},v_{1},v_{4}\},\,$$\{v_{7},v_{6},v_{1},v_{2}\}\bigr{\}}.$ One can verify that $\mathcal{H}$ contains no simplicial vertices. Therefore, $\mathcal{H}$ is not $\operatorname{W}$-chordal, and consequently, $\mathcal{C}_{r}(G_{t})$ is not $\operatorname{W}$-chordal. ∎
Remark 4.2.
In [1], the authors claimed that for any tree $G$, the hypergraph $\mathcal{C}_{r}(G)$ is $\operatorname{W}$-chordal for all $r\geq 2$ (see [1, Theorem 2.7]), and consequently, that the $r$-independence complex $\mathcal{I}_{r}(G)$ is shellable. However, their proof contains a flaw. Proposition 4.1 shows that for every $r\geq 4$, there exists a tree $G$ (constructed by taking each $H_{i}$ to be a tree for $i=1,2,3$) such that $\mathcal{C}_{r}(G)$ is not $\operatorname{W}$-chordal. This provides a counterexample to their claim and, in turn, invalidates the shellability conclusion.
The following theorem provides a sufficient condition for $\mathcal{C}_{r}(G)$ to be $\operatorname{W}$-chordal when $r$ is close to the number of vertices in $G$.
Theorem 4.3.
Let $G=(V(G),E(G))$ be a graph with $|V(G)|=n\geq 3$. If $n-2\leq r\leq n-1$, then $\mathcal{C}_{r}(G)$ is a $\operatorname{W}$-chordal hypergraph.
Proof.
If $r=n-1$, then each edge of $\mathcal{C}_{r}(G)$ has size $n$, which implies that there can be at most one such edge. Hence, $\mathcal{C}_{r}(G)$ is trivially $\operatorname{W}$-chordal.
Now suppose $r=n-2$. Then each edge in $\mathcal{C}_{r}(G)$ has size $n-1$. If $|E(\mathcal{C}_{r}(G))|\leq 2$, then again the hypergraph is trivially $\operatorname{W}$-chordal. Assume $|E(\mathcal{C}_{r}(G))|\geq 3$. We first show that for any two distinct edges $f_{1},f_{2}\in\mathcal{C}_{r}(G)$, we have $|f_{1}\cap f_{2}|=n-2$. Since $|f_{1}|=|f_{2}|=n-1$, we have $|f_{1}\cup f_{2}|=2(n-1)-|f_{1}\cap f_{2}|.$ But $f_{1}\cup f_{2}\subseteq V(G)$, so
$22(n-1)-|f_{1}\cap f_{2}|\leq n\Rightarrow|f_{1}\cap f_{2}|\geq n-$ 2 .
Also, since $f_{1}\neq f_{2}$, we cannot have $|f_{1}\cap f_{2}|=n-1$. Hence, $|f_{1}\cap f_{2}|=n-2$, and so $f_{1}\cup f_{2}=V(G)$.
Now let $\mathcal{H}$ be any minor of $\mathcal{C}_{r}(G)$. We show that $\mathcal{H}$ contains a simplicial vertex. If $|E(\mathcal{H})|\leq 2$, the result is immediate. Assume $|E(\mathcal{H})|\geq 3$. Let $e_{1},e_{2}\in E(\mathcal{H})$ with $e_{1}\neq e_{2}$, and choose vertices $x_{1}^{\prime}\in e_{1}\setminus e_{2}$ and $x_{2}^{\prime}\in e_{2}\setminus e_{1}$.
We claim that $x_{1}^{\prime}$ is a simplicial vertex in $\mathcal{H}$. Let $e_{1}^{\prime}$ and $e_{2}^{\prime}$ be two distinct edges in $E(\mathcal{H})$ such that their intersection $e_{1}^{\prime}\cap e_{2}^{\prime}$ contains the vertex $x_{1}^{\prime}$. Since $x_{1}^{\prime}\in e_{1}\setminus e_{2}$, it follows that $e_{1}^{\prime},e_{2}^{\prime}\neq e_{2}$. Let $f_{1}^{\prime},f_{2}^{\prime},g_{2}\in E(\mathcal{C}_{r}(G))$ be preimages of $e_{1}^{\prime},e_{2}^{\prime},e_{2}$ under the minor operation, so that $f_{i}^{\prime}\cap V(\mathcal{H})=e_{i}^{\prime}$ for $i=1,2$ and $g_{2}\cap V(\mathcal{H})=e_{2}$. From earlier, we know $f_{1}^{\prime}\cup f_{2}^{\prime}=V(G)$, and since $x_{1}^{\prime}\notin g_{2}$, we have
$e_{2}=g_{2}\cap V(\mathcal{H})\subseteq(f_{1}^{\prime}\cup f_{2}^{\prime})\cap V(\mathcal{H})=e_{1}^{\prime}\cup e_{2}^{\prime}\subseteq V(\mathcal{H}).$
Since $x_{1}^{\prime}\notin e_{2}$, this implies $e_{2}\subseteq(e_{1}^{\prime}\cup e_{2}^{\prime})\setminus\{x_{1}^{\prime}\}$. Hence, $x_{1}^{\prime}$ is a simplicial vertex of $\mathcal{H}$. Therefore, every minor of $\mathcal{C}_{r}(G)$ contains a simplicial vertex, and so $\mathcal{C}_{r}(G)$ is $\operatorname{W}$-chordal. ∎
The following example illustrates that the lower bound provided in Theorem 4.3 is sharp, even within the class of chordal graphs.
Example 4.4.
Let $G$ be the graph with $V(G)=\{x_{1},\ldots,x_{15}\}$, as shown below. Consider the hypergraph $\mathcal{C}_{12}(G)$. Define $\mathcal{H}=\mathcal{C}_{12}(G)/x_{1}/x_{2}/x_{3}/x_{4}/x_{12}/x_{13}/x_{14}/x_{15}$.
$x_{1}$$x_{2}$$x_{3}$$x_{4}$$x_{5}$$x_{6}$$x_{7}$$x_{8}$$x_{9}$$x_{10}$$x_{11}$$x_{12}$$x_{13}$$x_{14}$$x_{15}$
Observe that $V(\mathcal{H})=$$\{x_{5}$, $x_{6}$, $x_{7}$, $x_{8}$, $x_{9}$, $x_{10}$, $x_{11}\}$ and the edge set $E(\mathcal{H})$ is given by: $E(\mathcal{H})=\{\{x_{5},x_{7},x_{8},x_{9},x_{11}\}$, $\{x_{5},x_{6},x_{8},x_{10},x_{11}\}$, $\{x_{5},x_{7},x_{8},x_{10},x_{11}\}$, $\{x_{5},x_{6},x_{8},x_{9},x_{11}\}$.
It can be observed that $\mathcal{H}$ contains no simplicial vertices. Therefore, we conclude that $\mathcal{C}_{12}(G)$ is not a $\operatorname{W}$-chordal hypergraph.
It is worth noting that the above result is sharp when $G$ is a cycle graph. Specifically, we establish the following:
Proposition 4.5.
Let $C_{n}$ be a cycle of length $n\geq 4$. Then, $\mathcal{C}_{r}(C_{n})$ is $\operatorname{W}$-chordal if and only if $n-2\leq r\leq n-1$.
Proof.
Let $V(C_{n})=\{x_{1},\ldots,x_{n}\}$ be the vertex set of the cycle $C_{n}$. Case 1: Suppose $r<n-2$. We show that $\mathcal{C}_{r}(C_{n})$ is not $\operatorname{W}$-chordal. Consider two edges $f_{1},f_{2}\in E(\mathcal{C}_{r}(C_{n}))$, where $f_{1}=\{x_{1},x_{2},\ldots,x_{r+1}\}\quad\text{and}\quad f_{2}=\{x_{n},x_{1},x_{2},\ldots,x_{r}\}.$ Note that $f_{1}\cup f_{2}\setminus\{x_{1}\}$ forms a disconnected set, as it consists of two disjoint components. Therefore, there does not exist an edge $f_{3}\in E(\mathcal{C}_{r}(C_{n}))$ such that $f_{3}\subseteq(f_{1}\cup f_{2})\setminus\{x_{1}\}$. Furthermore, by a similar argument, we can show that $\mathcal{C}_{r}(C_{n})$ does not contain any simplicial vertices. This implies that $\mathcal{C}_{r}(C_{n})$ is not $\operatorname{W}$-chordal when $r<n-2$. Case 2: Now, suppose $n-2\leq r\leq n-1$. By Theorem 4.3, it follows that $\mathcal{C}_{r}(C_{n})$ is $\operatorname{W}$-chordal in this case, since the condition $n-2\leq r\leq n-1$ satisfies the requirements for $\mathcal{C}_{r}(G)$ to be $\operatorname{W}$-chordal.
Thus, we conclude that $\mathcal{C}_{r}(C_{n})$ is $\operatorname{W}$-chordal if and only if $n-2\leq r\leq n-1$. ∎
In the following lemma, we establish a relationship between a cut vertex in the tree $G$ and a $c^{\prime}$-minor of the graph $\mathcal{C}_{r}(G)$, where $r\geq 1$.
Lemma 4.6.
Let $G$ be a tree, and let $\mathcal{H}$ be a $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$ for some $r\geq 1$. Following the notation in Set-Up 3.6, assume $\ell\neq 1$. Then, there exists a cut vertex $v$ of $G$ such that:
  1. (1)
    $v\in V(\mathcal{H})$, and
  2. (2)
    $V(\mathcal{H})\setminus\{v\}\subseteq V(G_{1})$,
where $G_{1}$ is a connected component of $G\setminus\{v\}$.
Proof.
Let $v\in V_{\ell}\cap V(\mathcal{H})$. Since $\mathcal{H}$ is a $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$, the vertex $v$ cannot form a singleton edge in $\mathcal{H}$, i.e., $\{v\}\notin E(\mathcal{H})$. Thus, $v$ must be adjacent to at least one vertex in $G\setminus\bigcup_{i=1}^{\ell-1}V_{i}$. As $v$ is a simplicial vertex in the subgraph $G\setminus\bigcup_{i=1}^{\ell-1}V_{i}$, its neighborhood in this subgraph consists of a single vertex: $N_{G\setminus\bigcup_{i=1}^{\ell-1}V_{i}}(v)=\{a\},$ for some $a\in V(G)$. Since $G$ is a tree and $v$ is a cut vertex, the graph $G\setminus\{v\}$ has at least two connected components. Let $G_{1}$ be the component containing $a$.
Now consider any vertex $x\in V(\mathcal{H})\setminus\{v\}$. Since $G$ is a tree, there exists a unique path from $v$ to $x$, say $vz_{1}z_{2}\cdots z_{t}x.$ Because $v\in V_{\ell}$, there is a path $a_{1}a_{2}\cdots a_{\ell-1}v$ with $a_{i}\in V_{i}$ for $11\leq i\leq\ell-$ 1. Also, since $x\in V(\mathcal{H})\setminus\{v\}$, and $x\in V_{j}$ for some $j\geq\ell$, there exists a path $xb_{1}b_{2}\cdots b_{r}$ with $r\geq\ell-1$. Let $z_{i}$ be any internal vertex on the path from $v$ to $x$. Then both
$a_{1}a_{2}\cdots a_{\ell-1}vz_{1}\cdots z_{i-1}z_{i}\quad\text{and}\quad z_{i}z_{i+1}\cdots z_{t}xb_{1}\cdots b_{r}$
are paths of length at least $\ell$. Therefore, $z_{i}\notin V_{s}$ for $1\leq s\leq\ell$, implying all internal vertices lie outside $\bigcup_{s=1}^{\ell}V_{s}$.
Since $a$ is the only neighbor of $v$ in $G\setminus\bigcup_{s=1}^{\ell-1}V_{s}$, the first edge of the path from $v$ to $x$ must be $vz_{1}=va$, so $z_{1}=a$. The rest of the path connects $a$ to $x$ within $G\setminus\{v\}$, showing that $x$ lies in the same connected component as $a$, i.e., $x\in V(G_{1})$.
Since $x\in V(\mathcal{H})\setminus\{v\}$ was arbitrary, we conclude that $V(\mathcal{H})\setminus\{v\}\subseteq V(G_{1}),$ as required. ∎
We now extend Theorem 4.3 to establish a lower bound in the case when $G$ is a forest. Let $G$ be a tree, and let $\mathcal{H}$ be a $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$ for some $r\geq 1$. Consider the cut vertex $v$ of $G$ as given by Lemma 4.6, and let $G_{1}$ be the component of $G\setminus\{v\}$ that contains $V(\mathcal{H})\setminus\{v\}$. Define $C_{v}=V(G)\setminus(V(G_{1})\cup\{v\})$, ensuring that $C_{v}\cap V(\mathcal{H})=\emptyset$.
We now state the following generalization:
Theorem 4.7.
Let $G$ be a forest. Then, for all $r\geq n-5$, the hypergraph $\mathcal{C}_{r}(G)$ is $\operatorname{W}$-chordal.
Proof.
Since forest graphs form a hierarchy, and by Proposition 3.1, it suffices to prove that every $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$ contains a simplicial vertex when $G$ is a connected graph. Let $\mathcal{H}$ be a $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$, and use the notation from Set-Up 3.6. If $\ell=1$, then by Lemma 3.2, $\mathcal{H}$ contains a simplicial vertex. Hence, assume that $\ell\neq 1$. Define the set of vertices
[…]
Let $a_{k}\in S$ be such that $|C_{a_{k}}|=\min\{\,|C_{a_{i}}|\mid i=1,\dots,m\,\}.$
We claim that $a_{k}$ is a simplicial vertex of $\mathcal{H}$. Suppose that $a_{k}\in e_{1}\cap e_{2}$ for distinct $e_{1},e_{2}\in E(\mathcal{H})$. By an argument similar to Lemma 3.8, we can choose $f_{1},f_{2}\in E(\mathcal{C}_{r}(G))$ such that
$C_{a_{k}}\subseteq f_{i}\quad\text{and}\quad e_{i}=f_{i}\cap V(\mathcal{H})\quad\text{for }i=1,2.$
Next, choose $x_{1}\in e_{1}\setminus e_{2}\quad\text{and}\quad x_{2}\in e_{2}\setminus e_{1}$ such that
$\operatorname{d}(x_{1},a_{k})\geq\operatorname{d}(x,a_{k})\quad\text{for all }x\in e_{1}\setminus e_{2},\quad\operatorname{d}(x_{2},a_{k})\geq\operatorname{d}(y,a_{k})\quad\text{for all }y\in e_{2}\setminus e_{1}.$
Since $a_{k}\in f_{1}$ and $x_{1}\in f_{1}$, there exists a path $a_{k}y_{1}\cdots y_{p}x_{1}$ in $G$ with $y_{i}\in f_{1}$. Similarly, there exists a path $a_{k}z_{1}\cdots z_{q}x_{2}$ with $z_{j}\in f_{2}$. If $y_{1}\neq z_{1}$, then $x_{1}$ and $x_{2}$ lie in different connected components of $G\setminus a_{k}$. However, this contradicts the fact that $x_{1}$ and $x_{2}$ belong to the same connected component of $G\setminus a_{k}$. Therefore, we must have $y_{1}=z_{1}$, and we denote this common vertex by $a$. This implies $a\in f_{1}\cap f_{2}$.
We now claim that either $x_{1}$ or $x_{2}$ belongs to $S$. Suppose, for contradiction, that neither $x_{1}$ nor $x_{2}$ belongs to $S$. For $i=1,2$, define
$D_{i}=\{\,y\in(V(G)\setminus x_{i})\mid\text{the induced path }P(y,a_{k})\text{ contains }x_{i}\text{ as an internal vertex}\,\}.$
Note that $D_{1}\cap f_{2}=\emptyset$ and $D_{2}\cap f_{1}=\emptyset$; otherwise, $x_{1}\in e_{2}$ or $x_{2}\in e_{1}$, contradicting the choice of $x_{1}$ and $x_{2}$. If $D_{i}\cap V(\mathcal{H})=\emptyset$ for both $i=1,2$, then, since $x_{1}$ and $x_{2}$ are cut vertices satisfying Lemma 4.6, we must have $x_{1},x_{2}\in S$, a contradiction. Hence, $D_{i}\cap V(\mathcal{H})\neq\emptyset$ for both $i$. Let $x_{i}^{\prime}\in D_{i}\cap V(\mathcal{H})$ for $i=1,2$. Moreover, $x_{1}^{\prime}\notin f_{1}$ by the maximality condition on $x_{1}$. Since $x_{i}^{\prime}\in V(\mathcal{H})$ and $\ell\neq 1$, there exists $x_{i}^{\prime\prime}\in N_{G}(x_{i}^{\prime})$ such that
$\operatorname{d}(x_{i}^{\prime\prime},a_{k})>\operatorname{d}(x_{i}^{\prime},a_{k})\text{ for }i=1,2.$
Also, $x_{2}^{\prime},x_{2}^{\prime\prime}\in D_{2}$, and since $D_{2}\cap f_{1}=\emptyset$, we have $x_{2}^{\prime},x_{2}^{\prime\prime}\notin f_{1}$. If $x_{1}^{\prime\prime}\in f_{1}$, then $x_{1}^{\prime}\in f_{1}$, a contradiction. Thus, $x_{1}^{\prime\prime}\notin f_{1}$. This implies that $f_{1}$ misses at least five vertices: $x_{1}^{\prime},x_{1}^{\prime\prime},x_{2},x_{2}^{\prime},x_{2}^{\prime\prime}\notin f_{1}$, contradicting the assumption $|f_{1}|=r+1\geq n-4$. Therefore, either $x_{1}\in S$ or $x_{2}\in S$. Without loss of generality, assume $x_{1}\in S$.
Let $D_{1}=C_{x_{1}}$. Define $D=f_{2}\setminus(C_{a_{k}}\cup\{a_{k}\}).$ By construction of $C_{a_{k}}$, we have $D\subseteq G_{1}$, where $G_{1}$ is the connected component of $G\setminus a_{k}$ containing $V(\mathcal{H})\setminus\{a_{k}\}$. Now, we prove that the set
$S^{\prime}=D\cup\{a,y_{2},\ldots,y_{p},x_{1}\}\cup D_{1}$
is connected. Let $\alpha,\beta\in D$. Then there exists a path in $G_{1}$, hence in $G$, hence in $f_{2}$, and no internal vertex lies in $C_{a_{k}}\cup\{a_{k}\}$, so $D$ is connected. Let $c\in D_{1}$ and $d\in D$. Then there exists a path $P(d,a)$ in $f_{2}\setminus(C_{a_{k}}\cup\{a_{k}\})$, and a path $ay_{2}\cdots y_{p}x_{1}r_{1}\cdots r_{t}c$ where $r_{i}\in D_{1}$. Thus, $S^{\prime}$ is connected.
Observe that $f_{2}=D\cup C_{a_{k}}\cup\{a_{k}\}$, so
$r+1=|D|+|C_{a_{k}}|+1.$
Also, $|S^{\prime}|\geq r+1$, so we can choose $f\in E(\mathcal{C}_{r}(G))$ such that $f\subseteq S^{\prime}$. Set $e=f\cap V(\mathcal{H})$. Clearly, $e\subseteq e_{1}\cup e_{2}\setminus\{a_{k}\}$. If $e\in E(\mathcal{H})$, we are done. Otherwise, there exists $e_{s}\in E(\mathcal{H})$ such that $e_{s}\subseteq e_{1}\cup e_{2}\setminus\{a_{k}\}$. This proves that $a_{k}$ is a simplicial vertex. ∎
Remark 4.8.
If $G$ is a forest with $|V(G)|=n$ and $r<n-5$ (for example, even when $r=n-6$), the hypergraph $\mathcal{C}_{r}(G)$ may fail to be $\operatorname{W}$-chordal. As an illustration, consider the construction presented in Proposition 4.1, where $|V(H_{i})|=1$ for each $i=1,2,3$ and $t=1$. In this case, $G_{t}$ is a tree with $|V(G_{t})|=10$, and it can be verified that $\mathcal{C}_{4}(G_{t})$ is not $\operatorname{W}$-chordal.
5. $\operatorname{W}$-Chordality in Block Graphs
In this section, we establish sufficient conditions for the $\operatorname{W}$-chordality of $\mathcal{C}_{r}(G)$ when $G$ is a block graph. Our first result demonstrates that the hypergraph $\mathcal{C}_{2}(G)$ associated with a block graph is inherently $\operatorname{W}$-chordal, underscoring the structural robustness of block graphs in hypergraph settings.
Theorem 5.1.
If $G$ is a block graph, then $\mathcal{C}_{2}(G)$ is a $\operatorname{W}$-chordal hypergraph.
Proof.
Block graphs admit a hierarchical structure. By Proposition 3.1, it suffices to show that every $c^{\prime}$-minor of $\mathcal{C}_{2}(G)$ contains a simplicial vertex, assuming $G$ is connected. Let $\mathcal{H}$ be a $c^{\prime}$-minor of $\mathcal{C}_{2}(G)$. Without loss of generality, assume $G$ is connected and adopt the notation from Set-up 3.6. If $1\leq n\leq 2$, then $G$ is a star-clique graph, and by Corollary 3.3, $\mathcal{H}$ has a simplicial vertex. Now assume $n\geq 3$. If $\ell=1$, then Lemma 3.2 implies that $\mathcal{H}$ has a simplicial vertex. So suppose $\ell\neq 1$. In particular, $x_{0}\notin V(\mathcal{H})$, where $x_{0}\in V(B_{1})\cap V_{1}$. By Lemma 3.7, $x_{1}\in V(\mathcal{H})$, and since $x_{2}\in N_{G}(x_{1})\setminus\{x_{0}\}$, we also have $x_{2}\in V(\mathcal{H})$.
We now show that $|B_{1}|=2$. If $|B_{1}|\geq 3$, then $\{x_{1}\}$ would be an edge of $\mathcal{H}$, contradicting the definition of a $c^{\prime}$-minor. So $|B_{1}|=2$. If $V(B_{2})=\{x_{1},x_{2}\}$, then $x_{1}$ appears in only one hyperedge of $\mathcal{H}$, hence is simplicial. Otherwise, suppose $V(B_{2})=\{x_{1},x_{2},\beta_{1},\dots,\beta_{p}\}$ for some $p\geq 1$. Then for each $i$, we have $\beta_{i}\in V_{2}\cap V(\mathcal{H})$ and $|N_{G}(\beta_{i})\cap V_{1}|=1$. Let $N_{G}(\beta_{i})\cap V_{1}=\{z_{i}\}$.
Now consider $x_{1}\in e_{1}\cap e_{2}$, where $e_{1}\neq e_{2}$ and $e_{i}\in E(\mathcal{H})$ for $i=1,2$. Since $x_{1}\in V_{2}$, Observation 3.5 implies $S_{x_{1}}=\{x_{0}\}$. By Lemma 3.8, there exist $f_{1},f_{2}\in E(\mathcal{C}_{2}(G))$ such that $S_{x_{1}}\subseteq f_{i}$ and the set $f_{1}\cup f_{2}\setminus\{x_{0},x_{1}\}$ is connected. Each $f_{i}$ is either $\{x_{0},x_{1},x_{2}\}$ or $\{x_{0},x_{1},\beta_{j}\}$ for some $j$. Let $x_{1}^{\prime}\in e_{1}\setminus e_{2}$ and $x_{2}^{\prime}\in e_{2}\setminus e_{1}$. Without loss of generality, assume $x_{1}^{\prime}=\beta_{j}$ for some $j$. Then $|f_{1}\cup f_{2}\setminus\{x_{0},x_{1}\}|=2<r+1=3.$ Define $C=\{z_{j}\}$. Then $C$ satisfies the conditions in (3.1). Hence, by Lemma 3.10, $x_{1}$ is a simplicial vertex of $\mathcal{H}$. ∎
By extending Theorem 5.1, the following result shows that block graphs with bounded diameter retain the $\operatorname{W}$-chordality of the hypergraph $\mathcal{C}_{r}(G)$ for all $r\geq 2$, thereby highlighting a strong connection between graph diameter and the $\operatorname{W}$-chordality of associated hypergraphs.
Theorem 5.2.
If $G$ is a block graph with $\operatorname{diam}(G)\leq 4$, then $\mathcal{C}_{r}(G)$ is a $\operatorname{W}$-chordal hypergraph for all $r\geq 2$.
Proof.
Let $\mathcal{H}$ be a $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$ for some $r\geq 2$. We claim that $\mathcal{H}$ contains a simplicial vertex. To prove this, assume that $G$ is connected and adopt the notation from Set-up 3.6. We consider cases based on the number of blocks $n$.
Case 1: $1\leq n\leq 2$. In this case, $G$ is a star-clique graph. By Corollary 3.3, $\mathcal{H}$ contains a simplicial vertex.
Case 2: $3\leq n\leq 4$. If $r=2$, then by Theorem 5.1, $\mathcal{C}_{2}(G)$ is $\operatorname{W}$-chordal, so $\mathcal{H}$ contains a simplicial vertex.
Now assume $r\geq 3$. We analyze based on the value of $\ell$.
Subcase: $\ell=1$. Then, by Lemma 3.2, $\mathcal{H}$ has a simplicial vertex.
Subcase: $\ell=2$. Let $V_{2}\cap V(\mathcal{H})=\{v_{1},\dots,v_{s}\}$. By Observation 3.5, define
$|S_{v_{k}}|=\min\left\{|S_{v_{i}}|:1\leq i\leq s\right\}.$
We claim that $v_{k}$ is simplicial in $\mathcal{H}$. Let $e_{1},e_{2}\in E(\mathcal{H})$ be distinct edges such that $v_{k}\in e_{1}\cap e_{2}$. By Lemma 3.8, there exist $f_{1},f_{2}\in E(\mathcal{C}_{r}(G))$ such that
$S_{v_{k}}\subseteq f_{i},\quad f_{i}\cap V(\mathcal{H})=e_{i}\quad\text{for }i=1,2,$
and the set $f_{1}\cup f_{2}\setminus(S_{v_{k}}\cup\{v_{k}\})$ is connected.
We analyze this for different values of $n$:
Case $n=3$. Here, $V_{2}=B_{2}\setminus V_{1}$. Let $x_{1}^{\prime}\in e_{1}\setminus e_{2}$ and $x_{2}^{\prime}\in e_{2}\setminus e_{1}$. Since $x_{1}^{\prime},x_{2}^{\prime}\in V_{2}$, there are two possibilities:
  • If $\left|f_{1}\cup f_{2}\setminus(S_{v_{k}}\cup\{v_{k}\})\right|\geq r+1$, then by Lemma 3.10, $v_{k}$ is simplicial.
  • If $\left|f_{1}\cup f_{2}\setminus(S_{v_{k}}\cup\{v_{k}\})\right|\leq r$, then $x_{1}^{\prime}=v_{j}$ for some $j$, and we let $C=S_{v_{j}}$. Then $C$ satisfies the conditions in (3.1), and again by Lemma 3.10, $v_{k}$ is simplicial.
Case $n=4$. Here, $V_{2}=N_{G}(x_{2})\setminus V_{1}$. Let $x_{1}^{\prime}\in e_{1}\setminus e_{2}$ and $x_{2}^{\prime}\in e_{2}\setminus e_{1}$. If $x_{1}^{\prime},x_{2}^{\prime}\notin V_{2}$, then they belong to $V_{t}$ for $t\geq 3$, which is impossible since $V_{3}=\{x_{2}\}$ and $V_{t}=\emptyset$ for $t\geq 4$. So either $x_{1}^{\prime}$ or $x_{2}^{\prime}$ lies in $V_{2}$, and we can proceed as in the $n=3$ case to construct an edge $e_{3}\subseteq(e_{1}\cup e_{2})\setminus\{v_{k}\}$, confirming that $v_{k}$ is simplicial.
Note that $\ell=3$ does not occur. For $n=3$, we have $V_{3}=\emptyset$, and for $n=4$, $V_{3}=\{x_{2}\}$ contradicts the definition of a $c^{\prime}$-minor. Also, $V_{t}=\emptyset$ for all $t\geq 4$.
Thus, every $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$ contains a simplicial vertex. Since every induced subgraph of $G$ is a block graph with diameter at most $4$, Proposition 3.1 implies that $\mathcal{C}_{r}(G)$ is a $\operatorname{W}$-chordal hypergraph for all $r\geq 2$. ∎
It is important to note that if a block graph has diameter $5$, then the hypergraph $\mathcal{C}_{r}(G)$ may not necessarily be $\operatorname{W}$-chordal. The following example illustrates that $\mathcal{C}_{3}(G)$ need not be $\operatorname{W}$-chordal when $G$ is a block graph with $\operatorname{diam}(G)=5$.
Example 5.3.
Let $G$ be the graph shown below, referred to as the $\operatorname{T}_{3}$-graph. Consider the hypergraph $\mathcal{C}_{3}(G)$ and define $\mathcal{H}=\mathcal{C}_{3}(G)/x_{1}/x_{7}/x_{9}.$
$x_{1}$$x_{2}$$x_{3}$$x_{4}$$x_{5}$$x_{6}$$x_{7}$$x_{8}$$x_{9}$
$\operatorname{T}_{3}$-graph
The vertex set of $\mathcal{H}$ is given by $V(\mathcal{H})=\{x_{2},x_{3},x_{4},x_{5},x_{6},x_{8}\},$ and the edge set is $E(\mathcal{H})=$$\{\{x_{2},x_{3},x_{4}\}$, $\{x_{3},x_{5},x_{6}\}$, $\{x_{2},x_{3},x_{5}\}$, $\{x_{8},x_{3},x_{4}\}$, $\{x_{4},x_{5},x_{6}\}$, $\{x_{8},x_{4},x_{5}\}\}.$ One can observe that $\mathcal{H}$ contains no simplicial vertex. Therefore, we conclude that $\mathcal{C}_{3}(G)$ is not a $\operatorname{W}$-chordal hypergraph.
We present a theorem that extends the notion of $\operatorname{W}$-chordality from block graphs to a broader class of graphs, namely forests, under a diameter constraint. As a precursor, we recall the definition of a lobster tree.
Let $G$ be a tree, and let $P:=y_{1}y_{2}\cdots y_{n}$ be a longest path in $G$, referred to as the central path, where $y_{1}$ and $y_{n}$ are leaves. Then:
Caterpillar Tree::
$G$ is a caterpillar if every vertex $z\notin V(P)$ is adjacent to an internal vertex of $P$; that is,
$\operatorname{d}(z,y_{i})=1\quad\text{for some }1<i<n.$
Lobster Tree::
$G$ is a lobster if either:
  1. (i):
    it is a caterpillar, or
  2. (ii):
    every vertex $z\notin V(P)$ satisfies
    $\operatorname{d}(z,y_{i})=2\quad\text{for some }i\in\{3,\ldots,n-2\}.$
Theorem 5.4.
If $G$ is a forest with $\operatorname{diam}(G)\leq 5$, then $\mathcal{C}_{r}(G)$ is a $\operatorname{W}$-chordal hypergraph for all $r\geq 2$.
Proof.
If $\operatorname{diam}(G)\leq 4$, then by Theorem 5.2, $\mathcal{C}_{r}(G)$ is $\operatorname{W}$-chordal for all $r\geq 2$. Now assume $\operatorname{diam}(G)=5$. By Proposition 3.1, it suffices to show that every $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$ has a simplicial vertex when $G$ is connected. Let $\mathcal{H}$ be a $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$. Since $G$ is a tree with diameter $5$, there exists a maximum-length path $x_{0}x_{1}x_{2}x_{3}x_{4}x_{5}$. Thus, $G$ is a lobster tree. Assume the notation from Set-up 3.6. If $\ell=1$, then by Lemma 3.2, $\mathcal{H}$ contains a simplicial vertex. Consider next the case $\ell=2$. Note that $V_{2}=A_{1}\cup A_{2}$, where
$A_{1}=N_{G}(x_{3})\setminus(V_{1}\cup\{x_{2}\}),\quad A_{2}=N_{G}(x_{2})\setminus(V_{1}\cup\{x_{3}\}).$
Let $V_{2}\cap V(\mathcal{H})=\{v_{1},\ldots,v_{s}\}$. Using the notation from Observation 3.5, define
$|S_{v_{k}}|=\min\left\{|S_{v_{i}}|\mid i=1,\ldots,s\right\}.$
We claim that $v_{k}$ is a simplicial vertex of $\mathcal{H}$. To prove this, consider two edges $e_{1},e_{2}\in E(\mathcal{H})$ such that $v_{k}\in e_{1}\cap e_{2}$ and $e_{1}\neq e_{2}$. By Lemma 3.8, there exist edges $f_{1},f_{2}\in E(\mathcal{C}_{r}(G))$ such that $S_{v_{k}}\subseteq f_{i},\quad f_{i}\cap V(\mathcal{H})=e_{i}\quad\text{for }i=1,2,\text{and }f_{1}\cup f_{2}\setminus\left(S_{v_{k}}\cup\{v_{k}\}\right)\text{ is connected.}$. Let $x_{1}^{\prime}\in e_{1}\setminus e_{2}$ and $x_{2}^{\prime}\in e_{2}\setminus e_{1}$. Since $v_{k}\in V_{2}$, we have $v_{k}\in A_{1}$ or $v_{k}\in A_{2}$. Without loss of generality, assume $v_{k}\in A_{1}$.
We claim that either $x_{1}^{\prime}\in V_{2}$ or $x_{2}^{\prime}\in V_{2}$. Suppose not. Since $V_{t}=\emptyset$ for all $t\geq 4$ and $\ell=2$, it follows that $x_{1}^{\prime},x_{2}^{\prime}\in V_{3}=\{x_{2},x_{3}\}$. Without loss of generality, let $x_{1}^{\prime}=x_{2}$ and $x_{2}^{\prime}=x_{3}$. Then $v_{k}x_{3}(=x_{2}^{\prime})x_{2}(=x_{1}^{\prime})$ is a unique path from $v_{k}$ to $x_{1}^{\prime}$. Since both $v_{k},x_{1}^{\prime}\in f_{1}$, and $x_{3}=x_{2}^{\prime}\in f_{1}$, we contradict the assumption that $x_{2}^{\prime}\in e_{2}\setminus e_{1}$. Hence, either $x_{1}^{\prime}\in V_{2}$ or $x_{2}^{\prime}\in V_{2}$.
Assume $x_{1}^{\prime}\in V_{2}$ and let $x_{1}^{\prime}=v_{j}$ for some $j$. If $\left|f_{1}\cup f_{2}\setminus\left(S_{v_{k}}\cup\{v_{k}\}\right)\right|\geq r+1,$ then by Lemma 3.10, $v_{k}$ is a simplicial vertex. Otherwise, suppose $\left|f_{1}\cup f_{2}\setminus\left(S_{v_{k}}\cup\{v_{k}\}\right)\right|\leq r.$ Define $C=S_{v_{j}}$. Then $C$ satisfies the conditions of (3.1). Since $|C|\geq|S_{v_{k}}|$, Lemma 3.10 implies $v_{k}$ is simplicial in $\mathcal{H}$.
Now consider $\ell=3$. Let $v\in V_{3}\cap V(\mathcal{H})$. Since $V_{3}=\{x_{2},x_{3}\}$ and $G\setminus(V_{1}\cup V_{2})=V_{3}$, any such vertex $v$ lies in a unique edge $\{x_{2},x_{3}\}$. Hence, $v$ is simplicial. Finally, observe that $\ell\geq 4$ is not possible since $V_{t}=\emptyset$ for all $t\geq 4$. Thus, every $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$ contains a simplicial vertex, completing the proof. ∎
Remark 5.5.
If $G$ is a forest with $\operatorname{diam}(G)\geq 6$, then $\mathcal{C}_{r}(G)$ need not be a $\operatorname{W}$-chordal hypergraph for all $r\geq 2$. For instance, consider the graph described in Proposition 4.1 with $|V(H_{i})|=1$ for each $i=1,2,3$ and $t=1$. In this case, the resulting graph $G_{t}$ is a tree with $\operatorname{diam}(G_{t})=6$, and it follows that $\mathcal{C}_{4}(G_{t})$ is not a $\operatorname{W}$-chordal hypergraph.
The following theorem characterizes $\operatorname{W}$-chordality of the hypergraph $\mathcal{C}_{3}(G)$ in terms of the absence of specific substructures (i.e., -freeness) in the block graph $G$.
Theorem 5.6.
Let $G$ be a block graph. Then, $G$ is $\operatorname{T}_{3}$-free if and only if $\mathcal{C}_{3}(G)$ is $\operatorname{W}$-chordal.
Proof.
By Proposition 3.1, it suffices to show that every $c^{\prime}$-minor of $\mathcal{C}_{3}(G)$ has a simplicial vertex when $G$ is connected. Let $\mathcal{H}$ be a $c^{\prime}$-minor of $\mathcal{C}_{3}(G)$, following the notation from Set-up 3.6.
If $1\leq n\leq 4$, then by Theorem 5.2, $\mathcal{H}$ has a simplicial vertex. Now assume $n\geq 5$. Suppose there exists a vertex $y\in V(B_{3})\setminus\{x_{2},x_{3}\}$ such that $d(z,y)=2$, i.e., there is a path $yaz$ for some $z\in V(G)$. Since $a\in V(B_{3})$, if $a\notin V(B_{3})$, then $G$ contains a $\operatorname{T}_{3}$, contradicting the assumption that $G$ is $\operatorname{T}_{3}$-free.
If $\ell=1$, then by Lemma 3.2, $\mathcal{H}$ contains a simplicial vertex. We now consider the case $\ell\neq 1$.
Case 1: Suppose $(N_{G}(x_{2})\setminus N_{G}(x_{3}))\cap V_{2}\cap V(\mathcal{H})\neq\emptyset.$
This forces $\ell=2$. Let $(N_{G}(x_{2})\setminus N_{G}(x_{3}))\cap V_{2}\cap V(\mathcal{H})=\{v_{1},\ldots,v_{q}\}.$ Using the notation from Observation 3.5, define
$|S_{v_{k}}|=\min\left\{|S_{v_{i}}|\mid i\in\{1,\ldots,q\}\right\}.$
We claim that $v_{k}$ is a simplicial vertex of $\mathcal{H}$. To see this, let $e_{1},e_{2}\in E(\mathcal{H})$ be such that $v_{k}\in e_{1}\cap e_{2}$ and $e_{1}\neq e_{2}$. By Lemma 3.8, there exist edges $f_{1},f_{2}\in E(\mathcal{C}_{3}(G))$ such that $S_{v_{k}}\subseteq f_{i}$ and $f_{i}\cap V(\mathcal{H})=e_{i}$ for $i=1,2$. Moreover, the set $f_{1}\cup f_{2}\setminus\left(S_{v_{k}}\cup\{v_{k}\}\right)$ is connected. Let $x_{1}^{\prime}\in e_{1}\setminus e_{2}$ and $x_{2}^{\prime}\in e_{2}\setminus e_{1}$. Note that $|S_{v_{k}}|\leq 2$; otherwise, $\{v_{k}\}\in E(\mathcal{H})$, contradicting the assumption that $\mathcal{H}$ is a $c^{\prime}$-minor.
Subcase 1: $|S_{v_{k}}|=1$. Let $S_{v_{k}}=\{z_{0}\}$. We first claim that either $x_{1}^{\prime}$ or $x_{2}^{\prime}$ must belong to $V_{2}$. Since $|f_{i}|=4$ and $S_{v_{k}}\subseteq f_{i}$, it follows that $x_{1}^{\prime},x_{2}^{\prime}\in\{x_{2},x_{3}\}$. As $x_{1}^{\prime}\neq x_{2}^{\prime}$, we may assume without loss of generality that $x_{1}^{\prime}=x_{2}$ and $x_{2}^{\prime}=x_{3}$. Since $v_{k},x_{3}\in f_{2}$, there exists a shortest path $P(v_{k},x_{3})$ in $f_{2}$, with all internal vertices in $f_{2}$. Given $z_{0}\in f_{2}$ and $|f_{2}|=4$, and using the fact that $v_{k}\in(N_{G}(x_{2})\setminus N_{G}(x_{3}))\cap V_{2}\cap V(\mathcal{H})$, we conclude that $P(v_{k},x_{3}):=v_{k}x_{2}x_{3}.$ This is a contradiction since $x_{2}=x_{1}^{\prime}\in f_{2}$ but $x_{1}^{\prime}\in e_{1}\setminus e_{2}$. Therefore, either $x_{1}^{\prime}\in V_{2}$ or $x_{2}^{\prime}\in V_{2}$. Without loss of generality, assume $x_{1}^{\prime}\in V_{2}$, and let $C=S_{x_{1}^{\prime}}$. Then $C$ satisfies the conditions in (3.1). Since $|C|\geq|S_{v_{k}}|$, Lemma 3.10 implies that $v_{k}$ is simplicial.
Subcase 2: $|S_{v_{k}}|=2$. We claim that at least one of $x_{1}^{\prime}$ or $x_{2}^{\prime}$ lies in $\{v_{1},\ldots,v_{q}\}$. Suppose not. Then $x_{1}^{\prime},x_{2}^{\prime}\notin\{v_{1},\ldots,v_{q}\}$, which would force $x_{1}^{\prime}=x_{2}^{\prime}=x_{2}$, a contradiction. Hence, without loss of generality, assume $x_{1}^{\prime}=v_{j}$ for some $j$, and let $C=S_{v_{j}}$. Then $C$ satisfies conditions (i)–(iv) in (3.1). Furthermore, $|f_{1}\cup f_{2}\setminus(S_{v_{k}}\cup\{v_{k}\})|=2<3.$ So by Lemma 3.10, $v_{k}$ is simplicial.
Case 2: Suppose $(N_{G}(x_{2})\setminus N_{G}(x_{3}))\cap V_{2}\cap V(\mathcal{H})=\emptyset.$
Case A: Assume $(N_{G}(x_{3})\setminus N_{G}(x_{4}))\cap V_{2}\cap V(\mathcal{H})\neq\emptyset.$ This implies $\ell=2$. Let
$(N_{G}(x_{3})\setminus N_{G}(x_{4}))\cap V_{2}\cap V(\mathcal{H})=\{v_{1}^{\prime},\ldots,v_{p}^{\prime}\}.$
Using Observation 3.5, choose $v_{t}^{\prime}$ such that $|S_{v_{t}^{\prime}}|=\min\{|S_{v_{j}^{\prime}}|\mid 1\leq j\leq p\}.$ Take two distinct edges $e_{1},e_{2}\in E(\mathcal{H})$ with $v_{t}^{\prime}\in e_{1}\cap e_{2}$. By Lemma 3.8, there exist edges $f_{1},f_{2}\in E(\mathcal{C}_{3}(G))$ such that $S_{v_{t}^{\prime}}\subseteq f_{i}$ for each $i$, $f_{i}\cap V(\mathcal{H})=e_{i}$, and $f_{1}\cup f_{2}\setminus(S_{v_{t}^{\prime}}\cup\{v_{t}^{\prime}\})$ is connected. Let $x_{1}^{\prime}\in e_{1}\setminus e_{2}$ and $x_{2}^{\prime}\in e_{2}\setminus e_{1}$. Note $|S_{v_{t}^{\prime}}|\leq 2$; otherwise $\{v_{t}^{\prime}\}\in E(\mathcal{H})$, contradicting $\mathcal{H}$ being a $c^{\prime}$-minor.
Suppose there exists another maximal clique path
$\operatorname{CP}(A_{3},A_{2},A_{1},B_{4},\ldots,B_{n})$
where each $A_{i}$ is a clique ($1\leq i\leq 3$) with connecting vertices $b,a,x_{3},\ldots,x_{n-1}$.
  • If $(N_{G}(a)\setminus N_{G}(x_{3}))\cap V_{2}\cap V(\mathcal{H})\neq\emptyset$, then the conditions of Case 1 are satisfied, and consequently there exists a simplicial vertex in $\mathcal{H}$.
  • Otherwise ($(N_{G}(a)\setminus N_{G}(x_{3}))\cap V_{2}\cap V(\mathcal{H})=\emptyset$), we have $b\notin V(\mathcal{H})$.
Subcase 1: $|S_{v_{t}^{\prime}}|=1$.
Either $x_{1}^{\prime}$ or $x_{2}^{\prime}$ belongs to $V_{2}\cup\{x_{2},a\}$. Without loss of generality, assume $x_{1}^{\prime}\in V_{2}\cup\{x_{2},a\}$. Define
$C=\begin{cases}S_{x_{1}^{\prime}}&\text{if }x_{1}^{\prime}\in V_{2},\\ \{x_{1}\}&\text{if }x_{1}^{\prime}=x_{2},\\ \{b\}&\text{if }x_{1}^{\prime}=a.\end{cases}$
Then $C$ satisfies (3.1) and $|C|\geq|S_{v_{t}^{\prime}}|$, so by Lemma 3.10, $v_{t}^{\prime}$ is simplicial.
Subcase 2: $|S_{v_{t}^{\prime}}|=2$.
Similarly, either $x_{1}^{\prime}$ or $x_{2}^{\prime}$ belongs to $\{v_{1}^{\prime},\ldots,v_{p}^{\prime}\}\cup\{x_{2}\}$. Let $x_{0}\in V(B_{1})\cap V_{1}$ and note that
$|f_{1}\cup f_{2}\setminus(S_{v_{t}^{\prime}}\cup\{v_{t}^{\prime}\})|=2<4.$
Assuming $x_{1}^{\prime}\in\{v_{1}^{\prime},\ldots,v_{p}^{\prime}\}\cup\{x_{2}\}$, define
$C=\begin{cases}S_{x_{j}^{\prime}}&\text{if }x_{1}^{\prime}=v_{j}^{\prime},\\ \{x_{0},x_{1}\}&\text{if }x_{1}^{\prime}=x_{2}.\end{cases}$
Again $C$ satisfies (3.1) and $|C|\geq|S_{v_{t}^{\prime}}|$, proving $v_{t}^{\prime}$ is simplicial via Lemma 3.10.
If no such maximal clique path $\operatorname{CP}(A_{3},A_{2},A_{1},B_{4},\ldots,B_{n})$ exists, analogous arguments show $v_{t}^{\prime}$ remains simplicial in $\mathcal{H}$ Case B: Suppose $(N_{G}(x_{3})\setminus N_{G}(x_{4}))\cap V_{2}\cap V(\mathcal{H})=\emptyset.$
Observe that $x_{0},x_{1}\notin V(\mathcal{H})$, where $x_{0}\in V(B_{1})\cap V_{1}$. By Lemma 3.7, it follows that $x_{2}\in V(\mathcal{H})$. Since $x_{3}\in N_{G}(x_{2})\setminus\{x_{1}\}$, Lemma 3.7 further implies that $x_{3}\in V(\mathcal{H})$. Given that
$(N_{G}(x_{2})\setminus N_{G}(x_{3}))\cap V_{2}\cap V(\mathcal{H})=\emptyset,$
we conclude that the only edge containing $x_{2}$ is $\{x_{2},x_{3}\}$, since $\{x_{0},x_{1},x_{2},x_{3}\}\cap V(\mathcal{H})=\{x_{2},x_{3}\}.$ Thus, $x_{2}$ is a simplicial vertex of $\mathcal{H}$.
The converse follows from Example 5.3 and the fact that for any $x\in V(G)$ and $r\geq 1$, $\mathcal{C}_{r}(G)\setminus x=\mathcal{C}_{r}(G\setminus x).$
As an immediate consequence of Theorem 5.6, we obtain the following result:
Corollary 5.7.
If $G$ is a forest, then the hypergraph $\mathcal{C}_{3}(G)$ is $\operatorname{W}$-chordal.
Remark 5.8.
It is worth emphasizing that even if a graph is both a block graph and $\operatorname{T}_{3}$-free, the associated hypergraph $\mathcal{C}_{4}(G)$ need not be $\operatorname{W}$-chordal. A concrete example is provided by the graph $G_{t}$ discussed in Remark 5.5. Although $G_{t}$ satisfies both properties—being a block graph and $\operatorname{T}_{3}$-free—the corresponding hypergraph $\mathcal{C}_{4}(G_{t})$ is not $\operatorname{W}$-chordal.
We now introduce the notion of $\mathcal{T}_{1}$-graphs, which serve as a natural generalization of caterpillar tree graphs. The following definition formalizes this concept.
Definition 5.9.
A connected graph $H$ is called a $\mathcal{T}_{1}$-graph if it satisfies the following conditions:
  1. (1)
    $H$ is a block graph.
  2. (2)
    $H$ contains a maximum-length clique path subgraph, denoted by $\operatorname{CP}(B_{1},\dots,B_{n})$ for some $n\geq 1$, with connecting vertices $x_{1},\dots,x_{n-1}$.
  3. (3)
    For every vertex $z\in S$, where $S=V(H)\setminus V(\operatorname{CP}(B_{1},\dots,B_{n}))$, at least one of the following holds:
    1. (i)
      $z$ is adjacent to a connecting vertex $x_{i}$ for some $11\leq i\leq n-$ 1;
    2. (ii)
      if $n\geq 3$, then $z$ is adjacent to a vertex in $B_{2}$ or $B_{n-1}$;
    3. (iii)
      if $n\geq 4$, then there exists a connecting vertex $x_{i}$ with $i\in\{2,n-2\}$ and a vertex $y\in S$ such that $y$ is adjacent to both $x_{i}$ and $z$.
A graph $G$ is called a $\mathcal{T}_{1}$-graph if each of its connected components is a $\mathcal{T}_{1}$-graph.
To illustrate Definition 5.9, consider the following example.
$x_{1}$$x_{2}$$x_{3}$$x_{4}$$x_{5}$
Figure 1. Graph $H$
Let $H$ be the graph shown in the figure, containing a $6$-clique path subgraph $\operatorname{CP}(B_{1},B_{2},\dots,B_{6})$. The cliques in this path are: $B_{1}$, a clique on 2 vertices; $B_{2}$, a clique on 4 vertices; $B_{3}$, a clique on 3 vertices; $B_{4}$, a clique on 2 vertices; $B_{5}$, a clique on 4 vertices; and $B_{6}$, a clique on 2 vertices. The $6$-clique path subgraph is highlighted in the figure using dotted lines, with connecting vertices labeled $x_{1},x_{2},x_{3},x_{4},x_{5}$.
For each vertex $z\in S$, where $S=V(H)\setminus V(\operatorname{CP}(B_{1},B_{2},\dots,B_{6}))$, one of the following conditions holds:
  1. (i)
    $z$ is adjacent to a connecting vertex $x_{i}$ for some $1\leq i\leq 5$;
  2. (ii)
    $z$ is adjacent to a vertex in $B_{2}$ or $B_{5}$; or
  3. (iii)
    there exists a connecting vertex $x_{i}$ ($i=2$ or $i=4$) and a vertex $y\in S$ such that $y$ is adjacent to both $x_{i}$ and $z$.
Since $H$ satisfies all the conditions in Definition 5.9, it is a $\mathcal{T}_{1}$-graph.
Next, we introduce the concept of a $\mathcal{T}_{2}$-graph, which generalizes the notion of a lobster tree. The following definition formalizes this concept.
Definition 5.10.
A connected graph $H$ is called a $\mathcal{T}_{2}$-graph if it satisfies conditions (1), (2), and (3)(i)–(ii) of Definition 5.9, along with the following generalized condition that replaces (3)(iii):
  1. (iii)
    If $n\geq 4$, there exists a connecting vertex $x_{i}$ ($22\leq i\leq n-$ 2) and a vertex $y\in S$ such that $y$ is adjacent to both $x_{i}$ and $z$.
A graph $G$ is called a $\mathcal{T}_{2}$-graph if every connected component of $G$ is a $\mathcal{T}_{2}$-graph.
To illustrate Definition 5.10, consider the graph $H$ shown in the figure.
$x_{1}$$x_{2}$$x_{3}$$x_{4}$$x_{5}$$x_{6}$
Graph $H$
It contains a $7$-clique path subgraph $\operatorname{CP}(B_{1},B_{2},\dots,B_{7})$, where the cliques are: $B_{1}$, a $2$-vertex clique; $B_{2}$, a $5$-vertex clique; $B_{3}$, a $3$-vertex clique; $B_{4}$, a $4$-vertex clique; $B_{5}$, a $2$-vertex clique; $B_{6}$, a $6$-vertex clique; and $B_{7}$, a $3$-vertex clique. The cliques in the $7$-clique path subgraph are indicated by dotted lines in the figure. For every vertex $z\in S$, where $S=V(H)\setminus V(\operatorname{CP}(B_{1},B_{2},\dots,B_{7}))$, one of the following conditions holds:
  1. (i)
    $z$ is adjacent to a connecting vertex $x_{i}$ ($1\leq i\leq 6$),
  2. (ii)
    $z$ is adjacent to a vertex in $B_{2}$ or $B_{6}$, or
  3. (iii)
    there exists a connecting vertex $x_{i}$ ($2\leq i\leq 5$) and a vertex $y\in S$ such that $y$ is adjacent to both $x_{i}$ and $z$.
Since $H$ satisfies all the conditions of Definition 5.10, it is a $\mathcal{T}_{2}$-graph.
The following observation illustrates that both $\mathcal{T}_{1}$-graphs and $\mathcal{T}_{2}$-graphs form a hierarchy.
Observation 5.11.
Let $G$ be a connected $\mathcal{T}_{1}$-graph. Then, $G$ contains a maximum-size $n$-clique path subgraph $\operatorname{CP}(B_{1},\dots,B_{n})$ for some $n\geq 1$. Since $G$ is a block graph, every vertex in $G$ is either a simplicial vertex or a cut vertex.
Consider a vertex $x\in V(G)$. Suppose $x$ is a simplicial vertex. In this case, $G\setminus x$ remains connected. Assume $x\in B_{1}$ or $x\in B_{n}$; without loss of generality, let $x\in B_{1}$. If $\deg(x)=1$, we can select a different maximum-size clique path that satisfies all the conditions in Definition 5.9. If $\deg(x)>1$, we can retain the same maximum-size $n$-clique path $\operatorname{CP}(B_{1},\dots,B_{n})$ while ensuring that all conditions in Definition 5.9 remain satisfied. A similar argument holds if $x\in B_{n}$. If $x\notin B_{1}\cup B_{n}$, we can still retain the same maximum-size $n$-clique path $\operatorname{CP}(B_{1},\dots,B_{n})$ while preserving the conditions in Definition 5.9.
Now, suppose $x$ is a cut vertex. In this case, the removal of $x$ results in connected components, each of which contains a maximum-size clique path satisfying the conditions of Definition 5.9. Therefore, the structure of a $\mathcal{T}_{1}$-graph is preserved, establishing a hierarchy.
A similar argument shows that the structure of a $\mathcal{T}_{2}$-graph also forms a hierarchy.
We now present a key result concerning the structure of $\mathcal{T}_{2}$-graphs and their associated hypergraphs. Specifically, we show that the hypergraph $\mathcal{C}_{4}(G)$ of a $\mathcal{T}_{2}$-graph $G$ is $\operatorname{W}$-chordal.
Theorem 5.12.
If $G$ is a $\mathcal{T}_{2}$-graph, then $\mathcal{C}_{4}(G)$ is a $\operatorname{W}$-chordal hypergraph.
Proof.
By Proposition 3.1 and Observation 5.11, it suffices to show that every $c^{\prime}$-minor of $\mathcal{C}_{4}(G)$ has a simplicial vertex, assuming $G$ is connected. Let $\mathcal{H}$ be a $c^{\prime}$-minor of $\mathcal{C}_{4}(G)$, and let us follow the notation introduced in Set-up 3.6.
If $1\leq n\leq 4$, then by Theorem 5.2, the hypergraph $\mathcal{H}$ contains a simplicial vertex. Now suppose $n\geq 5$. If $\ell=1$, then Lemma 3.2 guarantees that $\mathcal{H}$ has a simplicial vertex. It remains to consider the case when $\ell\neq 1$.
Case 1: Suppose $N_{G}(x_{2})\cap V_{2}\cap V(\mathcal{H})\neq\emptyset$. This forces $\ell=2$. Let $N_{G}(x_{2})\cap V_{2}\cap V(\mathcal{H})=\{v_{1},\ldots,v_{q}\}.$ Using the notation from Observation 3.5, define $v_{k}$ such that
$|S_{v_{k}}|=\min\left\{|S_{v_{i}}|\mid i\in\{1,\ldots,q\}\right\}.$
We claim that $v_{k}$ is a simplicial vertex of $\mathcal{H}$. To establish this, consider any two edges $e_{1},e_{2}\in E(\mathcal{H})$ with $v_{k}\in e_{1}\cap e_{2}$ and $e_{1}\neq e_{2}$. By Lemma 3.8, there exist edges $f_{1},f_{2}\in E(\mathcal{C}_{4}(G))$ such that $S_{v_{k}}\subseteq f_{i}$ and $f_{i}\cap V(\mathcal{H})=e_{i}$ for $i=1,2$. Moreover, the set $f_{1}\cup f_{2}\setminus\left(S_{v_{k}}\cup\{v_{k}\}\right)$ is connected.
Let $x_{1}^{\prime}\in e_{1}\setminus e_{2}$ and $x_{2}^{\prime}\in e_{2}\setminus e_{1}$. Note that $|S_{v_{k}}|\leq 3$; otherwise, if $|S_{v_{k}}|\geq 4$, then $\{v_{k}\}\in E(\mathcal{H})$, contradicting the assumption that $\mathcal{H}$ is a $c^{\prime}$-minor. Now, following the analysis in Case 1 of Theorem 5.6, we observe: If $|S_{v_{k}}|=1$, then either $x_{1}^{\prime}$ or $x_{2}^{\prime}$ must belong to $V_{2}$. If $2\leq|S_{v_{k}}|\leq 3$, then either $x_{1}^{\prime}$ or $x_{2}^{\prime}$ must lie in $\{v_{1},\ldots,v_{q}\}$. Without loss of generality, assume:
$x_{1}^{\prime}\in\begin{cases}V_{2},&\text{if }|S_{v_{k}}|=1,\\ \{v_{1},\ldots,v_{q}\},&\text{if }2\leq|S_{v_{k}}|\leq 3.\end{cases}$
In the second case, let $x_{1}^{\prime}=v_{j}$ for some $j\in\{1,\ldots,q\}$. Define
$C=\begin{cases}S_{x_{1}^{\prime}},&\text{if }x_{1}^{\prime}\in V_{2},\\ S_{v_{j}},&\text{if }x_{1}^{\prime}=v_{j}.\end{cases}$
Then $C$ satisfies all the conditions listed in (3.1). Since $|C|\geq|S_{v_{k}}|$, Lemma 3.10 implies that $v_{k}$ is a simplicial vertex of $\mathcal{H}$.
Case 2: Suppose $N_{G}(x_{2})\cap V_{2}\cap V(\mathcal{H})=\emptyset$.
Subcase 1: Suppose $N_{G}(x_{3})\cap V_{2}\cap V(\mathcal{H})\neq\emptyset$, which forces $\ell=2$. Let $N_{G}(x_{3})\cap V_{2}\cap V(\mathcal{H})=\{v_{1}^{\prime},\ldots,v_{s}^{\prime}\}.$ Using the notation from Observation 3.5, define
$|S_{v_{k}^{\prime}}|=\min\left\{|S_{v_{i}^{\prime}}|\mid i\in\{1,\ldots,s\}\right\}.$
We claim that $v_{k}^{\prime}$ is a simplicial vertex of $\mathcal{H}$. To prove this, consider two distinct edges $e_{1},e_{2}\in E(\mathcal{H})$ with $v_{k}^{\prime}\in e_{1}\cap e_{2}$. By Lemma 3.8, there exist edges $f_{1},f_{2}\in E(\mathcal{C}_{4}(G))$ such that $S_{v_{k}^{\prime}}\subseteq f_{i}$ and $f_{i}\cap V(\mathcal{H})=e_{i}$ for $i=1,2$. Moreover, the set $f_{1}\cup f_{2}\setminus\left(S_{v_{k}^{\prime}}\cup\{v_{k}^{\prime}\}\right)$ is connected. Let $x_{1}^{\prime}\in e_{1}\setminus e_{2}$ and $x_{2}^{\prime}\in e_{2}\setminus e_{1}$.
We analyze the three possibilities for $|S_{v_{k}^{\prime}}|$:
  1. (1)
    If $|S_{v_{k}^{\prime}}|=1$: As in Case 2 of Theorem 5.6, one of $x_{1}^{\prime}$ or $x_{2}^{\prime}$ must belong to $V_{2}\cup\{x_{2}\}$. Without loss of generality, assume $x_{1}^{\prime}\in V_{2}\cup\{x_{2}\}$. Define
    $C=\begin{cases}S_{x_{1}^{\prime}}&\text{if }x_{1}^{\prime}\in V_{2},\\ \{x_{1}\}&\text{if }x_{1}^{\prime}=x_{2}.\end{cases}$
  2. (2)
    If $|S_{v_{k}^{\prime}}|=2$: A similar argument as above shows that $x_{1}^{\prime}\in\{v_{1}^{\prime},\ldots,v_{s}^{\prime}\}\cup\{x_{2}\}$. Define
    $C=\begin{cases}S_{v_{j}^{\prime}}&\text{if }x_{1}^{\prime}=v_{j}^{\prime},\\ \{x_{0},x_{1}\}&\text{if }x_{1}^{\prime}=x_{2},\end{cases}$
    where $x_{0}\in V(B_{1})\cap V_{1}$.
  3. (3)
    If $|S_{v_{k}^{\prime}}|=3$: Then either $x_{1}^{\prime}$ or $x_{2}^{\prime}$ belongs to $\{v_{1}^{\prime},\ldots,v_{s}^{\prime}\}$. Without loss of generality, assume $x_{1}^{\prime}=v_{j}^{\prime}$ and set $C=S_{v_{j}^{\prime}}$.
In all cases, $C$ satisfies conditions (i)–(iv) in (3.1). Hence, by Lemma 3.10, $v_{k}^{\prime}$ is a simplicial vertex of $\mathcal{H}$.
Subcase 2: Suppose $N_{G}(x_{3})\cap V_{2}\cap V(\mathcal{H})=\emptyset$.
First, assume $x_{2}\notin V(\mathcal{H})$. Then, by Lemma 3.7, it follows that $x_{3},x_{4}\in V(\mathcal{H})$. Since both
$N_{G}(x_{2})\cap V_{2}\cap V(\mathcal{H})=\emptyset\quad\text{and}\quad N_{G}(x_{3})\cap V_{2}\cap V(\mathcal{H})=\emptyset,$
the only edge in $\mathcal{H}$ containing $x_{3}$ is $\{x_{3},x_{4}\}$, implying that $x_{3}$ is a simplicial vertex.
Assume $x_{2}\in V(\mathcal{H})$. Since $\mathcal{H}$ is a $c^{\prime}$-minor and cannot contain isolated edges, at least one of $x_{3}$ or $x_{4}$ must also be in $V(\mathcal{H})$. Otherwise, $\{x_{2}\}$ would be an isolated edge, violating the $c^{\prime}$-minor condition. Define $f=\{x_{0},x_{1},x_{2},x_{3},x_{4}\}$, where $x_{0}\in V(B_{1})\cap V_{1}$. The intersection $f\cap V(\mathcal{H})$ is:
$f\cap V(\mathcal{H})=\begin{cases}\{x_{2},x_{3}\}&\text{if }x_{3}\in V(\mathcal{H}),\ x_{4}\notin V(\mathcal{H}),\\ \{x_{2},x_{4}\}&\text{if }x_{3}\notin V(\mathcal{H}),\ x_{4}\in V(\mathcal{H}),\\ \{x_{2},x_{3},x_{4}\}&\text{if }x_{3},x_{4}\in V(\mathcal{H}).\end{cases}$
Given that
$N_{G}(x_{2})\cap V_{2}\cap V(\mathcal{H})=\emptyset\quad\text{and}\quad N_{G}(x_{3})\cap V_{2}\cap V(\mathcal{H})=\emptyset,$
no additional vertices from $V_{2}$ are adjacent to $x_{2}$ or $x_{3}$ in $\mathcal{H}$. Thus, the only edge containing $x_{2}$ is $f\cap V(\mathcal{H})$, making $x_{2}$ a simplicial vertex in $\mathcal{H}$. Thus, in all cases, $\mathcal{H}$ contains a simplicial vertex, completing the proof. ∎
The following corollary characterizes lobster tree graphs in terms of the combinatorial properties of their associated hypergraphs $\mathcal{C}_{4}(G)$.
Corollary 5.13.
Let $G$ be a tree. Then $G$ is a lobster tree graph if and only if $\mathcal{C}_{4}(G)$ is $\operatorname{W}$-chordal.
Proof.
Suppose that $G$ is a lobster tree graph. Then, by Theorem 5.12, the associated hypergraph $\mathcal{C}_{4}(G)$ is $\operatorname{W}$-chordal.
Conversely, assume that $G$ is not a lobster tree graph. Then the diameter of $G$ is at least 6, i.e., $\operatorname{diam}(G)\geq 6$. Since $G$ is a tree, it contains a longest path of the form $x_{0}x_{1}\cdots x_{d}$ with $d\geq 6$. Moreover, there exists a path of the form $x_{p}y_{1}\cdots y_{m}$ for some $33\leq p\leq d-$ 3 and $m\geq 3$, such that $y_{1}$ is not equal to either $x_{p-1}$ or $x_{p+1}$, i.e., $y_{1}\neq x_{p-1}$ and $y_{1}\neq x_{p+1}$. Let $S^{\prime}=\{x_{p-3},x_{p-2},x_{p-1},x_{p},x_{p+1},x_{p+2},x_{p+3},y_{1},y_{2},y_{3}\}$, and $S=V(G)\setminus S^{\prime}$. Consider the minor $\mathcal{H}$ of $\mathcal{C}_{4}(G)$ obtained by deleting the vertices in $S$ and contracting $x_{p-3}$, $x_{p+3}$, and $y_{3}$. The vertex set of $\mathcal{H}$ is: $V(\mathcal{H})=\{x_{p-2},x_{p-1},x_{p},x_{p+1},x_{p+2},y_{1},y_{2}\},$ and the edge set is:
$E(\mathcal{H})=\left\{\begin{aligned} &\{x_{p-2},x_{p-1},x_{p},x_{p+1}\},\quad\{x_{p-2},x_{p-1},x_{p},y_{1}\},\quad\{x_{p+2},x_{p+1},x_{p},x_{p-1}\},\\ &\{x_{p+2},x_{p+1},x_{p},y_{1}\},\quad\{y_{2},y_{1},x_{p},x_{p+1}\},\quad\{y_{2},y_{1},x_{p},x_{p-1}\}\end{aligned}\right\}.$
In the hypergraph $\mathcal{H}$, none of the vertices are simplicial. Hence, $\mathcal{H}$ is not $\operatorname{W}$-chordal, and thus $\mathcal{C}_{4}(G)$ is not $\operatorname{W}$-chordal. This contradicts the assumption that $\mathcal{C}_{4}(G)$ is $\operatorname{W}$-chordal.
Therefore, $G$ must be a lobster tree graph. ∎
Theorem 5.12 asserts that if $G$ is a $\mathcal{T}_{2}$ graph, then $\mathcal{C}_{4}(G)$ is a $\operatorname{W}$-chordal hypergraph. However, the following example illustrates that this result does not necessarily hold for $\mathcal{C}_{5}(G)$ when $G$ is a $\mathcal{T}_{2}$-graph.
Example 5.14.
Let $G$ be a graph as shown in figure. Define $\mathcal{H}=\mathcal{C}_{5}(G)/x_{1}/x_{2}/x_{7}/x_{11}/x_{14}/x_{15}$.
$x_{2}$$x_{3}$$x_{4}$$x_{5}$$x_{1}$$x_{6}$$x_{7}$$x_{8}$$x_{9}$$x_{10}$$x_{11}$$x_{12}$$x_{13}$$x_{15}$$x_{14}$
Then, the vertex set of $\mathcal{H}$ is $V(\mathcal{H})$ = $\{x_{3}$, $x_{4}$, $x_{5}$, $x_{6}$, $x_{8}$, $x_{9}$, $x_{10}$, $x_{12}$, $x_{13}\}$, and its edge set is $E(\mathcal{H})$ = $\Big{\{}\{x_{3},x_{4},x_{5},x_{6}\}$, $\{x_{5},x_{6},x_{8},x_{9},x_{10}\}$, $\{x_{9},x_{10},x_{12},x_{13}\}$, $\{x_{5},x_{8},x_{9},x_{10},x_{12}\},$
$\{x_{4},x_{5},x_{8},x_{9},x_{10}\}$, $\{x_{8},x_{9},x_{12},x_{13}\}$, $\{x_{5},x_{6},x_{8},x_{9},x_{12}\}$, $\{x_{4},x_{5},x_{6},x_{8},x_{9}\}$, $\{x_{8},x_{3},x_{4},x_{5}\}\Big{\}}.$ It is clear that $\mathcal{H}$ contains no simplicial vertex. Therefore, $\mathcal{C}_{5}(G)$ is not a $\operatorname{W}$-chordal hypergraph.
We prove that for $\mathcal{T}_{1}$-graphs, the hypergraph $\mathcal{C}_{r}(G)$ is $\operatorname{W}$-chordal for all $r\geq 5$, thereby extending previous results to a wider range of $r$.
Theorem 5.15.
If $G$ is a $\mathcal{T}_{1}$-graph, then $\mathcal{C}_{r}(G)$ is a $\operatorname{W}$-chordal hypergraph for all $r\geq 5$.
Proof.
By Proposition 3.1 and Observation 5.11, it suffices to show that every $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$ contains a simplicial vertex whenever $G$ is connected and $r\geq 5$. Let $\mathcal{H}$ be such a $c^{\prime}$-minor, following the notation from Set-up 3.6. If $1\leq n\leq 4$, then Theorem 5.2 guarantees that $\mathcal{H}$ has a simplicial vertex. Now consider $n\geq 5$. If $\ell=1$, then Lemma 3.2 applies. Suppose $\ell=2$. Let $V_{2}\cap V(\mathcal{H})=\{v_{1},\ldots,v_{s}\}$. Define $|S_{v_{k}}|=\min\left\{|S_{v_{i}}|\mid i\in\{1,\ldots,s\}\right\}.$
We claim that $v_{k}$ is a simplicial vertex of $\mathcal{H}$. To show this, take any two distinct edges $e_{1},e_{2}\in E(\mathcal{H})$ with $v_{k}\in e_{1}\cap e_{2}$. By Lemma 3.8, there exist edges $f_{1},f_{2}\in E(\mathcal{C}_{r}(G))$ such that
$S_{v_{k}}\subseteq f_{i},\quad f_{i}\cap V(\mathcal{H})=e_{i}\quad\text{for }i=1,2,$
and $f_{1}\cup f_{2}\setminus(S_{v_{k}}\cup\{v_{k}\})$ is connected. Let $x_{1}^{\prime}\in e_{1}\setminus e_{2}$ and $x_{2}^{\prime}\in e_{2}\setminus e_{1}$.
We first claim that at least one of $x_{1}^{\prime},x_{2}^{\prime}$ lies in $V_{2}$. Suppose not; then both lie in $\{x_{2},\ldots,x_{n-2}\}$. Since
$V_{2}=A_{1}\cup A_{2},\quad A_{1}=N_{G}(x_{2})\setminus(V_{1}\cup\{x_{3}\}),\quad A_{2}=N_{G}(x_{n-2})\setminus(V_{1}\cup\{x_{n-3}\}),$
assume $v_{k}\in A_{1}$, $x_{1}^{\prime}=x_{p}$, $x_{2}^{\prime}=x_{q}$ with $p<q$. Since $v_{k},x_{2}^{\prime}\in f_{2}$, there is a shortest path $P(v_{k},x_{2}^{\prime})$ in $f_{2}$ with all internal vertices in $f_{2}$. Then
$P(v_{k},x_{2}^{\prime})=v_{k}x_{2}\cdots x_{q},\quad\text{so }x_{i}\in f_{2}\text{ for }i=2,\ldots,q.$
Since $x_{p}\in\{x_{2},\ldots,x_{q-1}\}$, we get $x_{p}=x_{1}^{\prime}\in e_{2}$, contradicting $x_{1}^{\prime}\in e_{1}\setminus e_{2}$. Thus, at least one of $x_{1}^{\prime},x_{2}^{\prime}$ lies in $V_{2}$. Assume $x_{1}^{\prime}=v_{j}\in V_{2}$, and define $C=S_{v_{j}}$. Then $C$ satisfies the conditions in (3.1). Since $|C|\geq|S_{v_{j}}|$, Lemma 3.10 ensures that $v_{j}$ is simplicial.
Now consider $\ell=3$. Then $V_{3}=\{x_{2},x_{n-2}\}$. Suppose $x_{2}\in V_{3}\cap V(\mathcal{H})$. If $x_{2}$ lies in two edges $e_{1}\neq e_{2}$ of $\mathcal{H}$, pick $x_{1}^{\prime}\in e_{1}\setminus e_{2}$, $x_{2}^{\prime}\in e_{2}\setminus e_{1}$ with $x_{1}^{\prime}=x_{p}$, $x_{2}^{\prime}=x_{q}$ and $p<q$. Then $P(x_{2},x_{q})=x_{2}x_{3}\cdots x_{q}$, so $x_{p}\in\{x_{3},\ldots,x_{q-1}\}\subseteq f_{2}$, contradicting the assumption. Thus, only one edge contains $x_{2}$.
A similar argument applies for $\ell\geq 4$, completing the proof. ∎
We define the hypergraph $\overline{\mathcal{C}_{r}(G)}$, the complement of the hypergraph $\mathcal{C}_{r}(G)$. This hypergraph is constructed on the same vertex set $V(G)$ as $\mathcal{C}_{r}(G)$, but its edges are given by the set
$\Big{\{}e\subseteq V(G)\mid|e|=r+1\text{ and }e\text{ is not an edge of }\mathcal{C}_{r}(G)\Big{\}}.$
Alternatively, $\overline{\mathcal{C}_{r}(G)}$ can be characterized as
$\overline{\mathcal{C}_{r}(G)}=\Big{\{}e\subseteq V(G)\mid|e|=r+1\text{ and each connected component of }$
$G[e]\text{ has at most }r\text{ vertices}\Big{\}}.$
For the special case $r=1$, the hypergraph $\overline{\mathcal{C}_{1}(G)}$ reduces to the complement of the graph $G$. This follows directly from the definition: $\overline{\mathcal{C}_{1}(G)}$ consists of all 2-element subsets of $V(G)$ that are not edges of $G$, which corresponds precisely to the edge set of the complement of $G$.
We are now ready to prove one of the main results of the paper.
Theorem 5.16.
Let $G=(V(G),E(G))$ be a graph with $|V(G)|=n\geq 3$. For each of the following cases, the independence complex $\mathcal{I}_{r}(G)$ is shellable and the Alexander dual $\mathcal{I}(\overline{\mathcal{C}_{r}(G)})^{\vee}$ is vertex decomposable:
  1. (1)
    For arbitrary graphs when $n-2\leq r\leq n-1$
  2. (2)
    For forests when $r\geq n-5$
  3. (3)
    For block graphs with $\operatorname{diam}(G)\leq 4$ when $r\geq 2$
  4. (4)
    For forests with $\operatorname{diam}(G)\leq 5$ when $r\geq 2$
  5. (5)
    For block graphs when $r=2$
  6. (6)
    For $T_{3}$-free block graphs (including all forests) when $r=3$
  7. (7)
    For $\mathcal{T}_{2}$-graphs when $r=4$
  8. (8)
    For $\mathcal{T}_{1}$-graphs when $r\geq 5$
Consequently, in all cases above, the edge ideal $I(\overline{\mathcal{C}_{r}(G)})$ has linear quotients and therefore a linear resolution.
Proof.
The assertion follows from Theorems 4.3, 4.7, 5.1, 5.2, 5.4, 5.6, 5.12, and 5.15, along with Theorem 2.2 and [35, Corollary 5.4, Theorem 6.9]. ∎
6. Shellability of Higher Independence Complexes
This section investigates the shellability of $r$-independence complexes $\mathcal{I}_{r}(G)$ for three fundamental graph constructions. Collectively, these results not only expand the known families of graphs with shellable independence complexes but also reveal new connections between graph operations, numerical parameters, and topological combinatorics.
Let $H$ be a graph. A subset $S\subseteq V(H)$ is called a vertex cover of $H$ if every edge of $H$ has at least one endpoint in $S$. Fix an integer $r\geq 1$. The graph $\operatorname{CCG}(H,S,r)$ is defined as the graph obtained from $H$ by attaching a star-clique graph $\operatorname{SC}(x)$ to each vertex $x\in S$, where $|V(\operatorname{SC}(x))|\geq r+1$.
Theorem 6.1.
Let $G=\operatorname{CCG}(H,S,r)$, where $r\geq 1$. Then $\mathcal{I}_{r}(G)$ is shellable.
Proof.
By Theorem 2.2(1), it suffices to show that every $c$-minor of $\mathcal{C}_{r}(G)$ contains a simplicial vertex. Let $\mathcal{H}$ be a $c$-minor of $\mathcal{C}_{r}(G)$. We consider two cases based on the intersection of $V(\mathcal{H})$ with the set of simplicial vertices of $G$.
Let $V_{1}$ denote the set of all simplicial vertices of $G$. Suppose first that $V_{1}\cap V(\mathcal{H})\neq\emptyset$. If there exists a vertex $v\in V_{1}\cap V(\mathcal{H})$ such that $\{v\}\notin E(\mathcal{H})$, then by Lemma 3.2, $v$ is a simplicial vertex of $\mathcal{H}$, and we are done. Assume now that $\{v\}\in E(\mathcal{H})$ for every $v\in V_{1}\cap V(\mathcal{H})$. We claim that if $a\in S\cap V(\mathcal{H})$, then $\{a\}\in E(\mathcal{H})$. Consider such a vertex $a$.
  • If $(\operatorname{SC}(a)\setminus\{a\})\cap V(\mathcal{H})\neq\emptyset$, then choose $c\in(\operatorname{SC}(a)\setminus\{a\})\cap V(\mathcal{H})$. Since $\{c\}\in E(\mathcal{H})$ by assumption, we can construct a connected set $\{c_{1},\ldots,c_{r},c\}$ such that $c_{i}\notin V(\mathcal{H})$ and $c_{i}\neq a$ for all $i$. Then $\{c_{1},\ldots,c_{r},a\}$ is connected, implying $\{a\}\in E(\mathcal{H})$.
  • If $(\operatorname{SC}(a)\setminus\{a\})\cap V(\mathcal{H})=\emptyset$, then since $|\operatorname{SC}(a)|\geq r+1$, we can form a connected set of size $r+1$ containing $a$ and $r$ elements from $\operatorname{SC}(a)\setminus\{a\}$. Hence, $\{a\}\in E(\mathcal{H})$.
We now claim that every edge $e\in E(\mathcal{H})$ satisfies $|e|=1$. Suppose for contradiction that some $e\in E(\mathcal{H})$ has $|e|\geq 2$. Let $a\in e$, and choose $b\in e\setminus\{a\}$ such that $\operatorname{d}(a,b)\leq\operatorname{d}(a,x)$ for all $x\in e\setminus\{a\}$.
  • If $a$ and $b$ are adjacent in $H$, then since $S$ is a vertex cover, either $a\in S$ or $b\in S$. Without loss of generality, assume $a\in S$. Then $\{a\}\in E(\mathcal{H})$ by the earlier claim, contradicting $|e|\geq 2$.
  • If $a$ and $b$ are not adjacent in $H$, then $e=f\cap V(\mathcal{H})$ for some $f\in E(\mathcal{C}_{r}(G))$, where $a,b\in f$. There exists a shortest path $ay_{1}\cdots y_{k}b$ with $k\geq 1$, $y_{i}\in f$ for all $i$, and $y_{i}\notin V(\mathcal{H})$. Since $a\notin S$ (otherwise $\{a\}\in E(\mathcal{H})$), and $y_{1}$ is adjacent to $a$, the vertex cover property implies $y_{1}\in S$.
    Suppose $\{c^{\prime}\}\in E(\mathcal{H})$ for some $c^{\prime}\in\operatorname{SC}(y_{1})\setminus\{y_{1}\}$. Then there exists a connected set $\{c_{1}^{\prime},\ldots,c_{r}^{\prime},c^{\prime}\}$ with $c_{i}^{\prime}\notin V(\mathcal{H})$ for all $i$. The set $\{c_{1}^{\prime},\ldots,c_{r}^{\prime},y_{1},a\}$ is connected with size at least $r+2$, implying $\{a\}\in E(\mathcal{H})$, a contradiction.
    Therefore, $(\operatorname{SC}(y_{1})\setminus\{y_{1}\})\cap V(\mathcal{H})=\emptyset$. But then $\operatorname{SC}(y_{1})\cup\{a\}$ is a connected set of size at least $r+1$, again implying $\{a\}\in E(\mathcal{H})$—another contradiction.
Thus, no such edge $e$ with $|e|\geq 2$ exists, and all edges in $\mathcal{H}$ are singletons. If $V_{1}\cap V(\mathcal{H})=\emptyset$, then again every edge of $\mathcal{H}$ is of size 1. In all cases, $\mathcal{H}$ contains a simplicial vertex. This completes the proof. ∎
The following result extends Theorem 6.1 by demonstrating the shellability of $\mathcal{I}_{r}(G)$ for a broader class of chordal graphs. Specifically, for a chordal graph $H$ and $G=\operatorname{CCG}(H,V(H),t)$, we show that $\mathcal{I}_{r}(G)$ is shellable for all $r\leq 2t+1$, thus relaxing the earlier bound on $r$.
Theorem 6.2.
Let $H$ be a chordal graph and let $G=\operatorname{CCG}(H,V(H),t)$ for some $t\geq 1$. If $r\leq 2t+1$, then $\mathcal{I}_{r}(G)$ is shellable.
Proof.
By Theorem 2.2(1), it suffices to show that every $c$-minor of $\mathcal{C}_{r}(G)$ has a simplicial vertex. Let $\mathcal{H}$ be a $c$-minor of $\mathcal{C}_{r}(G)$, i.e., $\mathcal{H}=\mathcal{C}_{r}(G)/V_{c}$ for some $V_{c}\subseteq V(G)$. Since $G$ is chordal, we adopt the notation from Observation 3.4.
Case 1: $V_{1}\cap V(\mathcal{H})\neq\emptyset$.
Subcase 1.1: There exists $v\in V_{1}\cap V(\mathcal{H})$ such that $\{v\}\notin E(\mathcal{H})$. By Lemma 3.2, $v$ is simplicial in $\mathcal{H}$.
Subcase 1.2: $\{v\}\in E(\mathcal{H})$ for all $v\in V_{1}\cap V(\mathcal{H})$. This implies that for every $x\in V(G)\setminus V(H)$, either $\{x\}\in E(\mathcal{H})$ or $x\in V_{c}$.
Claim 1: If $y\in V(\mathcal{H})\cap V(H)$ and $\{y\}\notin E(\mathcal{H})$, then $(\operatorname{SC}(y)\setminus\{y\})\cap V(\mathcal{H})=\emptyset.$
Proof of Claim 1: Suppose, for contradiction, that there exists $c\in(\operatorname{SC}(y)\setminus\{y\})\cap V(\mathcal{H})$. Since $\{c\}\in E(\mathcal{H})$, there exists an edge $f\in E(\mathcal{C}_{r}(G))$ of the form $f=\{b_{1},\ldots,b_{r},c\},$ where $f\cap V(\mathcal{H})=\{c\}$. However, $\{b_{1},\ldots,b_{r},y\}$ is connected in $G$, which would force $\{y\}\in E(\mathcal{H})$, a contradiction.
Claim 2: If $e\in E(\mathcal{H})$ and $|e|\neq 1$, then $e\in E(H)$.
Proof of Claim 2: Let $e\in E(\mathcal{H})$ with $|e|\geq 2$, and suppose $e=f\cap V(\mathcal{H})$ for some $f\in E(\mathcal{C}_{r}(G))$.
  • If $|e|\geq 3$, take $a\in e$ and choose $x\in e\setminus\{a\}$ minimizing $\operatorname{d}(a,x)$.
  • If $a$ and $x$ were non-adjacent in $G$, there would exist a shortest path $P(a,x)=(a=y_{0},y_{1},\ldots,y_{t}=x)$ in $f$ with all $y_{i}\notin V(\mathcal{H})$ for $11\leq i\leq t-$ 1 (since any $y_{i}\in V(\mathcal{H})$ would yield $\operatorname{d}(a,y_{i})<\operatorname{d}(a,x)$).
  • Considering $\mathrm{SC}(y_{1})$, any $c^{\prime}\in\mathrm{SC}(y_{1})\setminus\{y_{1}\}$ with $\{c^{\prime}\}\in E(\mathcal{H})$ would imply $\{a\}\in E(\mathcal{H})$, contradicting $\{a\}\notin E(\mathcal{H})$. Thus, $\mathrm{SC}(y_{1})\setminus\{y_{1}\}$ must be disjoint from $V(\mathcal{H})$.
  • The set $Q^{\prime}=\mathrm{SC}(a)\cup\mathrm{SC}(y_{1})$ is connected with $Q^{\prime}\cap V(\mathcal{H})=\{a\}$, forcing $\{a\}\in E(\mathcal{H})$, another contradiction.
Hence, $a$ and $x$ must be adjacent. Since $|\mathrm{SC}(a)\cup\mathrm{SC}(x)|\geq r+1$, there exists a connected set $g\subset\mathrm{SC}(a)\cup\mathrm{SC}(x)$ with $|g|=r+1$ containing $\{a,x\}$, yielding $g\cap V(\mathcal{H})=\{a,x\}\in E(\mathcal{H})$. This contradicts $|e|\geq 3$, so we must have $e=\{a,x\}\in E(H)$.
Thus, the edge set of $\mathcal{H}$ can be written as $E(\mathcal{H})=E(L)\cup\{\text{singleton edges}\},$ where $L$ is an induced subgraph of $H$. Since $L$ is chordal, it has a simplicial vertex, and therefore $\mathcal{H}$ also has a simplicial vertex.
Case 2: $V_{1}\cap V(\mathcal{H})=\emptyset$.
In this case, $E(\mathcal{H})=E(L^{\prime})\cup\{\text{singleton edges}\}$ for some induced subgraph $L^{\prime}$ of $H$. Since $L^{\prime}$ is chordal, it admits a simplicial vertex, and thus so does $\mathcal{H}$.
In both cases, $\mathcal{H}$ has a simplicial vertex, proving the theorem. ∎
The following example demonstrates that the complex $\mathcal{I}_{r}(G)$ may fail to be shellable when $r>2t+1$, even if $G=\operatorname{CCG}(H,V(H),t)$ for a chordal graph $H$ and $t\geq 1$.
Example 6.3.
Let $H$ be a chordal graph with vertex set $V(H)=\{a_{1},\ldots,a_{n},b_{1},\ldots,b_{n},c,d\}$ for some integer $n\geq 2$, and edge set
$E(H)=$$\big{\{}\{a_{i},a_{i+1}\}\mid 1\leq i\leq n-1\big{\}}\cup\big{\{}\{b_{i},b_{i+1}\}\mid 1\leq i\leq n-1\big{\}}$
$\cup\ \big{\{}\{a_{1},c\},\{a_{1},d\},\{b_{1},c\},\{b_{1},d\},\{c,d\}\big{\}}.$
Let $G=\operatorname{CCG}(H,V(H),t)$ be the graph obtained by attaching a star-clique graph $\operatorname{SC}(x)$ of size $t+1$ to each vertex $x\in V(H)$. We claim that $\mathcal{I}_{r}(G)$ is not shellable for $r=nt+n$. Assume, for contradiction, that $\mathcal{I}_{r}(G)$ is shellable. Define the set
$S=\left(\bigcup_{i=1}^{n}\operatorname{SC}(a_{i})\cup\bigcup_{j=1}^{n}\operatorname{SC}(b_{j})\right)\setminus\{a_{1},b_{1}\}.$
By repeatedly applying [27, Proposition 2.3] and [35, Lemma 2.2(1)], it follows that $\mathcal{I}(\mathcal{C}_{r}(G)/S)$ is shellable. However, one can verify that $\mathcal{C}_{r}(G)/S$ is isomorphic to the $4$-cycle $C_{4}$. By [34, Theorem 10], the complex $\mathcal{I}(C_{4})$ is not shellable, yielding a contradiction. Hence, $\mathcal{I}_{r}(G)$ is not shellable when $r=nt+n>2t+1$.
In [8], Cook II and Nagel introduced the notion of a clique whiskering of a graph. We generalize their construction as follows. Let $G$ be a graph. A clique vertex-partition of $G$ is a collection $\pi=\{W_{1},\ldots,W_{k}\}$ of pairwise disjoint cliques (some of which may be empty) such that $\bigcup_{i=1}^{k}W_{i}=V(G).$ Given a graph $G$, a clique vertex-partition $\pi=\{W_{1},\ldots,W_{k}\}$, and an integer $r\geq 1$, the $r$-clique whiskering of $G$ with respect to $\pi$, denoted $G^{\pi}_{r}$, is defined as follows. For each $1\leq i\leq k$, let $t_{i}\geq r$ be an integer. The vertex set of $G^{\pi}_{r}$ is
$V(G^{\pi}_{r})=V(G)\cup\bigcup_{i=1}^{k}\{x_{i,1},x_{i,2},\ldots,x_{i,t_{i}}\},$
where each $x_{i,j}$ is a new vertex added to $G$. The edge set of $G^{\pi}_{r}$ is
$E(G^{\pi}_{r})=E(G)\cup\bigcup_{i=1}^{k}\left\{\{a,b\}\mid a\neq b,\ a,b\in W_{i}\cup\{x_{i,1},x_{i,2},\ldots,x_{i,t_{i}}\}\right\}.$
In other words, for each clique $W_{i}$ in $\pi$, we introduce $t_{i}$ new vertices and form a clique on the vertex set $W_{i}\cup\{x_{i,1},\ldots,x_{i,t_{i}}\}$.
Note that a graph $G$ may admit multiple clique vertex-partitions. In particular, every graph has at least one such partition, namely the trivial partition
$\pi=\{\{x_{1}\},\ldots,\{x_{n}\}\},\quad\text{where }V(G)=\{x_{1},\ldots,x_{n}\}.$
Theorem 6.4.
Let $G$ be a graph, and let $\pi=\{W_{1},\ldots,W_{k}\}$ be a clique vertex-partition of $G$. Then $\mathcal{I}_{r}(G^{\pi}_{r})$ is shellable.
Proof.
By Theorem 2.2(1), it suffices to show that every $c$-minor of $\mathcal{C}_{r}(G^{\pi}_{r})$ contains a simplicial vertex. Let $\mathcal{H}$ be a $c$-minor of $\mathcal{C}_{r}(G^{\pi}_{r})$, and let $V_{1}$ denote the set of all simplicial vertices in $G^{\pi}_{r}$.
First, suppose that $V_{1}\cap V(\mathcal{H})\neq\emptyset$. If there exists a vertex $v\in V_{1}\cap V(\mathcal{H})$ such that $\{v\}\notin E(\mathcal{H})$, then by Lemma 3.2, $v$ is simplicial in $\mathcal{H}$. Otherwise, assume $\{v\}\in E(\mathcal{H})$ for all $v\in V_{1}\cap V(\mathcal{H})$. We claim that in this case, $\{a\}\in E(\mathcal{H})$ for every vertex $a\in V(\mathcal{H})$. If $a\in V_{1}\cap V(\mathcal{H})$, then $\{a\}\in E(\mathcal{H})$ by assumption. Now suppose $a\notin V_{1}$. Then $a\in V(G)$. Since $\pi=\{W_{1},\ldots,W_{k}\}$ is a clique vertex-partition of $G$, $a\in W_{i}$ for some $1\leq i\leq k$.
Consider two subcases:
  • Suppose $\{c\}\in E(\mathcal{H})$ for some $c\in\{x_{i,1},\ldots,x_{i,t_{i}}\}$. Then there exists a connected set $\{c_{1},\ldots,c_{r},c\}$ in $G^{\pi}_{r}$ such that $c_{i}\notin V(\mathcal{H})$ for all $1\leq i\leq r$. Since $W_{i}\cup\{x_{i,1},\ldots,x_{i,t_{i}}\}$ is a clique, it follows that $\{c_{1},\ldots,c_{r},a\}$ is also a connected set, so $\{a\}\in E(\mathcal{H})$.
  • Suppose $\{x_{i,1},\ldots,x_{i,t_{i}}\}\cap V(\mathcal{H})=\emptyset$. Since $t_{i}\geq r$, we can select distinct vertices $x_{i,1},\ldots,x_{i,r}$ such that $\{x_{i,1},\ldots,x_{i,r},a\}$ is connected in $G^{\pi}_{r}$, hence $\{a\}\in E(\mathcal{H})$.
Thus, in either case, $\{a\}\in E(\mathcal{H})$ for all $a\in V(\mathcal{H})$, so $\mathcal{H}$ is totally disconnected (i.e., all its edges are singleton sets).
Now suppose $V_{1}\cap V(\mathcal{H})=\emptyset$. Then every vertex of $\mathcal{H}$ lies in $V(G)$. Repeating the same argument as above, we again conclude that $\mathcal{H}$ is totally disconnected.
In both cases, $\mathcal{H}$ is totally disconnected, hence contains a simplicial vertex (every vertex is simplicial). This completes the proof. ∎
We introduce the notion of a clique cycle graph, which generalizes the classical concept of cycles. An $n$-clique cycle, denoted by $\operatorname{CC}(B_{1},\ldots,B_{n})$, is a graph formed by a cyclic sequence of $n$ cliques $B_{1},\ldots,B_{n}$, where $n\geq 3$. Each consecutive pair of cliques $B_{i}$ and $B_{i+1}$ (indices modulo $n$) intersects in exactly one vertex, denoted $x_{i}$. In particular, $B_{n}$ and $B_{1}$ share the vertex $x_{n}$, thereby completing the cycle. The vertices $x_{1},\ldots,x_{n}$ are called the connecting vertices of the $n$-clique cycle.
We define a graph $G(r)$, for a given integer $r\geq 1$, as follows: start with an $n$-clique cycle $\operatorname{CC}(B_{1},\ldots,B_{n})$ for some $n\geq 3$, and select a non-empty subset $S\subseteq\{x_{1},\ldots,x_{n}\}$ of the connecting vertices. To each vertex $x\in S$, attach a star-clique graph $\operatorname{SC}(x)$, such that $|V(\operatorname{SC}(x))|\geq r+1$ for at least one $x\in S$.
Note that $G(r)$ is not necessarily a block graph, and it may fail to be chordal when $n\geq 4$.
Theorem 6.5.
Let $G=G(r)$ for some $r\geq 1$. Then $\mathcal{I}_{r}(G)$ is shellable.
Proof.
By Theorem 2.2(1), it suffices to show that every $c$-minor of $\mathcal{C}_{r}(G)$ contains a simplicial vertex. Let $\mathcal{H}$ be a $c$-minor of $\mathcal{C}_{r}(G)$, i.e., $\mathcal{H}=\mathcal{C}_{r}(G)/S$ for some subset $S\subseteq V(G)$. Since $G=G(r)$ is constructed from an $n$-clique cycle with attached star-clique graphs, there exists a connecting vertex $x\in\{x_{1},\ldots,x_{n}\}$ such that $|V(\operatorname{SC}(x))|\geq r+1$. Let $V(\operatorname{SC}(x))=\{z_{1},\ldots,z_{m},x\},~{}\text{with }m\geq r.$
We consider the following two cases:
Case 1: $\{z_{1},\ldots,z_{m}\}\cap V(\mathcal{H})=\emptyset$.
We further divide this case: Suppose $x\in V(\mathcal{H})$. Then, since $x$ forms part of a star-clique of size at least $r+1$ and its neighbors are excluded from $\mathcal{H}$, we have $\{x\}\in E(\mathcal{H})$. By Proposition 3.1,
$\mathcal{H}\setminus x=(\mathcal{C}_{r}(G)/S)\setminus x=\mathcal{C}_{r}(G\setminus x)/S.$
Observe that $G\setminus x$ is a $\mathcal{T}_{1}$-graph. By Theorem 5.15, the complex $\mathcal{C}_{r}(G\setminus x)$ has a simplicial vertex, and hence so does $\mathcal{H}\setminus x$. Therefore, $\mathcal{H}$ itself has a simplicial vertex.
Suppose $x\notin V(\mathcal{H})$. Choose a vertex $c\in V(\mathcal{H})\cap\{x_{1},\ldots,x_{n}\}$ such that
$\operatorname{d}(c,z_{m})\leq\operatorname{d}(a,z_{m})\quad\text{for all }a\in V(\mathcal{H})\cap\{x_{1},\ldots,x_{n}\}.$
That is, $c$ is the connecting vertex closest to $z_{m}$ among those present in $\mathcal{H}$. Therefore $\{c\}\in E(\mathcal{H})$. Repeating the same argument as in above, we conclude that $\mathcal{H}$ contains a simplicial vertex. Case 2: $\{z_{1},\dots,z_{m}\}\cap V(\mathcal{H})\neq\emptyset$.
Let $y\in\{z_{1},\dots,z_{m}\}\cap V(\mathcal{H})$. If $y$ is not a singleton edge in $\mathcal{H}$, then Lemma 3.2 implies that $\mathcal{H}$ contains a simplicial vertex. If $y$ is a singleton edge in $\mathcal{H}$, then by an argument analogous to Case 1, $\mathcal{H}$ must also contain a simplicial vertex.
In all cases, the $c$-minor $\mathcal{H}$ has a simplicial vertex. Thus, by Theorem 2.2(1), $\mathcal{I}_{r}(G)$ is shellable. ∎
The following example illustrates that the conclusions of Theorems 6.1, 6.4, and 6.5 may fail when $\operatorname{CCG}(H,S,t)$, $G^{\pi}_{t}$, or $G(t)$ is considered with $t<r$.
Example 6.6.
Let $H$ be the $4$-cycle on vertices $\{x_{1},x_{2},x_{3},x_{4}\}$, and let $G$ be the graph obtained by attaching a whisker to each vertex of $H$. Clearly, $G$ lies in the class of graphs considered in Theorems 6.1, 6.4, and 6.5. Now, set $r=2$. Let $K$ denote the set of newly added whisker vertices, i.e., $K=V(G)\setminus\{x_{1},x_{2},x_{3},x_{4}\}$. Using [27, Proposition 2.3] and [35, Lemma 2.2(1)], we obtain
$\operatorname{link}_{\mathcal{I}_{r}(G)}(K)=\mathcal{I}(\mathcal{C}_{r}(G)/K)=\mathcal{I}(C_{4}).$
By [34, Theorem 10], the independence complex $\mathcal{I}(C_{4})$ is not shellable. Hence, $\mathcal{I}_{r}(G)$ is not shellable either.
Acknowledgments: We are deeply grateful to Priyavrat Deshpande, Anurag Singh and Russ Woodroofe for their invaluable insights and clarifications regarding their results, which greatly aided our work.
Data availability statement. Data sharing not applicable to this article as no datasets were generated or analysed during the current study.