### v3

# Shellability of Higher Independence Complexes of Graphs

Abstract.
This paper investigates the shellability of $r$-independence complexes $\mathcal{I}_{r}(G)$, a generalization of classical independence complexes introduced by Paolini and Salvetti. For a graph $G$, a subset $A\subseteq V(G)$ is $r$-independent if every connected component of the induced subgraph $G[A]$ has at most $r$ vertices. The associated simplicial complex $\mathcal{I}_{r}(G)$ has been the subject of significant interest due to its connections to combinatorial topology and commutative algebra. We address the classification problem for shellable $r$-independence complexes, focusing on block graphs, trees, and related families. Our main results establish sufficient conditions for shellability based on structural graph parameters such as diameter and forbidden subgraphs. Furthermore, we develop constructive techniques for generating shellable complexes through graph operations, including star-clique attachments, clique whiskering, and clique cycle constructions. These results extend and refine earlier work on classical independence complexes and provide a framework for understanding the topological and algebraic properties of higher independence complexes in structured graph families.
Key words and phrases:
$r$-independence complexes, shellability, block graphs, chordal hypergraphs, graph modifications.
AMS Classification 2020: 05E45, 13F55, 05E40
1. Introduction
Let $G=(V(G),E(G))$ be a simple (no loops, no multiple edges) undirected graph with vertex set $V(G)=\{x_{1},\dots,x_{n}\}$ and edge set $E(G)$. A subset $A\subseteq V(G)$ is called an independent set if no two vertices in $A$ are adjacent, equivalently, if the induced subgraph $G[A]$ contains no edges. The independence complex of $G$, denoted $\mathcal{I}(G)$, is the simplicial complex whose faces are the independent sets of $G$, that is, $\mathcal{I}(G)=\{A\subseteq V(G)\mid A\text{ is independent in }G\}.$ Independence complexes have been studied extensively due to their rich combinatorial, topological, and algebraic structure [2, 21, 22, 27, 30, 32]. A central theme in this area is the classification of graph families whose independence complexes exhibit desirable shellability properties [9, 16, 29, 30, 32, 33], and the construction of new graphs from given ones – via operations such as attaching whiskers, adding cliques, or other modifications – so that the resulting independence complexes are shellable [4, 8, 14, 15, 20, 29, 30, 32].
The concept of shellability has a rich history dating back to the 19th century, where it first appeared implicitly in the study of convex polytopes. A (not necessarily pure) simplicial complex $\Delta$ is said to be shellable if its facets can be ordered $F_{1},F_{2},\dots,F_{t}$ such that for each $k>1$, the simplicial complex $\langle F_{k}\rangle\cap\langle F_{1},\dots,F_{k-1}\rangle$ is pure and $(\dim F_{k}-1)$-dimensional, where $\langle\mathcal{F}\rangle$ denotes the simplicial complex generated by the family $\mathcal{F}$. Originally developed for pure complexes, the theory was significantly expanded in the 1990s when Björner, Lovász, and Yao [5] encountered nonpure complexes in complexity theory with topological properties resembling those of shellable complexes. This led Björner and Wachs [6, 7] to develop a comprehensive theory of shellability for nonpure complexes and posets. Shellability extends beyond simplicial complexes to various combinatorial structures and plays a crucial role in algebraic combinatorics and commutative algebra [19, 24, 27]. From an algebraic perspective, shellability provides a powerful combinatorial criterion for establishing the sequentially Cohen-Macaulay property of Stanley-Reisner rings. When a simplicial complex is shellable, its Stanley-Reisner ring is sequentially Cohen-Macaulay, with profound implications for both algebraic invariants and topological properties [24, 27]. Topologically, shellability guarantees strong connectivity properties: every pure $d$-dimensional shellable complex is homotopy equivalent to a wedge of $d$-spheres and is (d1)(d-1)-connected. Despite these elegant theoretical properties, the computational complexity of shellability is challenging – for every $d\geq 2$, determining whether a pure $d$-dimensional simplicial complex is shellable is NP-hard, hence NP-complete, [17].
Paolini and Salvetti [23] introduced the $r$-independence complex of a graph $G$, which generalizes the classical independence complex. A subset $A\subseteq V(G)$ is called $r$-independent if every connected component of the induced subgraph $G[A]$ contains at most $r$ vertices. The corresponding simplicial complex, denoted $\mathcal{I}_{r}(G)$, consists of all $r$-independent subsets of $V(G)$; in particular, $\mathcal{I}_{1}(G)=\mathcal{I}(G)$. These complexes have been studied in various contexts [1, 3, 10, 11, 12, 14, 28], with the shellability of $r$-independence complexes remaining a central question. In this work, we address the shellability of $r$-independence complexes through the framework of chordal hypergraph techniques introduced by Woodroofe [33]. In particular, we rely on the notion of chordal hypergraphs, which provides a powerful tool for analyzing shellability in generalized independence complexes.
Given a graph $G$ and an integer $r\geq 1$, we define the hypergraph $\mathcal{C}_{r}(G)$ associated with $G$ as follows: the vertex set of $\mathcal{C}_{r}(G)$ is $V(\mathcal{C}_{r}(G))=V(G)$, and the edge set consists of all subsets $S\subseteq V(G)$ of size $r+1$ such that the induced subgraph $G[S]$ is connected. Note that when $r=1$, the hypergraph $\mathcal{C}_{1}(G)$ coincides with the original graph $G$. The independence complex of this hypergraph coincides with the $r$-independence complex of $G$, i.e., $\mathcal{I}(\mathcal{C}_{r}(G))=\mathcal{I}_{r}(G)$ for all $r\geq 1$ [1, Proposition 2.6]. Recently, [1, Theorem 2.7] claimed that if $G$ is a tree, then $\mathcal{C}_{r}(G)$ is a chordal hypergraph for all $r\geq 2$, and consequently, by [33, Corollary 5.4], $\mathcal{I}_{r}(G)$ is shellable. However, using Proposition 4.1, we construct a class of counterexamples that show this claim does not hold in general. Hence, the shellability of $\mathcal{I}_{r}(G)$ for trees is not guaranteed and, to the best of our knowledge, no prior work has explored this direction.
We focus on block graphs, a classical family of graphs introduced by Harary [18]. Recall that a graph $G$ is a block graph if every block – that is, a maximal connected subgraph that remains connected upon the removal of any single vertex – is a complete graph. In this work, we establish several results connecting the structure of block graphs (and related tree-like graphs) to the shellability of their $r$-independence complexes.
Our main contributions are summarized as follows. First, for any graph $G$ with $n\geq 3$ vertices, we show that if $n-2\leq r\leq n-1$, then $\mathcal{I}_{r}(G)$ is shellable (Theorem 4.3). We then generalize this result to forests: for any forest $G$, we prove that $\mathcal{I}_{r}(G)$ is shellable for all $r\geq n-5$ (Theorem 4.5). Next, we establish that $\mathcal{I}_{2}(G)$ is shellable for any block graph $G$ (Theorem 4.6), and more generally, if a block graph has diameter at most $4$, then $\mathcal{I}_{r}(G)$ is shellable for all $r\geq 2$ (Theorem 4.7). For trees specifically, we prove that if $\operatorname{diam}(G)\leq 5$, then $\mathcal{I}_{r}(G)$ is shellable for all $r\geq 2$ (Theorem 4.8). We also explore shellability for larger independence parameters. If $G$ is a $\operatorname{T}_{3}$-free block graph, then $\mathcal{I}_{3}(G)$ is shellable, and as a corollary, the same holds for any forest (Theorem 4.9 and Corollary 4.10). Extending these ideas to more general tree-like structures, we show that if $G$ is a $\mathcal{T}_{2}$-graph (a generalization of lobster trees; see Definition 4.12), then $\mathcal{I}_{4}(G)$ is shellable (Theorem 4.14). Finally, if $G$ is a $\mathcal{T}_{1}$-graph (a generalization of caterpillar trees; see Definition 4.11), then $\mathcal{I}_{r}(G)$ is shellable for all $r\geq 5$ (Theorem 4.15).
The next focus of this paper is to investigate how various graph modifications – such as attaching whiskers, adding cliques, or performing other structural operations – affect the shellability of higher independence complexes. A foundational result by Villarreal [30] shows that for any graph $G$, the independence complex of its whiskered version $W(G)$ – formed by attaching a pendant vertex (whisker) to each vertex of $G$ – is Cohen–Macaulay. This was subsequently refined by Dochtermann and Engström [14] and independently by Woodroofe [32], who showed that $\mathcal{I}(W(G))$ is vertex decomposable. Since vertex decomposability implies shellability, this strengthens Villarreal’s result. Building on this, Cook and Nagel [8] introduced the vertex clique-whiskered graph $G^{\pi}$ and proved that $\mathcal{I}(G^{\pi})$ is both unmixed and vertex decomposable. Biermann et al. [4] further provided sufficient conditions on a subset $S\subseteq V(G)$ such that $\mathcal{I}(G\cup W(S))$ is vertex decomposable, where only the vertices in $S$ are whiskered.
Motivated by these developments, we study the shellability of $r$-independence complexes $\mathcal{I}_{r}(G)$ for a family of graphs constructed via star-clique attachments. Given a graph $H$ and a vertex cover $S\subseteq V(H)$, we define, for a fixed integer $r\geq 1$, the graph $\operatorname{CCG}(H,S,r)$ as the graph obtained by attaching a star-clique graph $\operatorname{SC}(x)$ with $|\operatorname{SC}(x)|\geq r+1$ to each vertex $x\in S$. Our main result in this setting is the following.
Theorem 5.1. Let $G=\operatorname{CCG}(H,S,r)$, where $r\geq 1$. Then $\mathcal{I}_{r}(G)$ is shellable.
This result extends to a broader class of chordal graphs. Specifically, if $H$ is chordal and $G=\operatorname{CCG}(H,V(H),t)$ for some integer $t\geq 1$, then $\mathcal{I}_{r}(G)$ is shellable whenever $r\leq 2t+1$ (Theorem 5.2). Moreover, we provide a counterexample demonstrating that this bound is sharp: the conclusion fails for $r>2t+1$ (Example 5.3). We also generalize the clique whiskering construction introduced by Cook and Nagel [8]. Given a graph $G$, a clique vertex-partition $\pi=\{W_{1},\ldots,W_{k}\}$, and an integer $r\geq 1$, the $r$-clique whiskering of $G$ with respect to $\pi$, denoted $G^{\pi}_{r}$, is obtained by adding $t_{i}\geq r$ new vertices to each clique $W_{i}$ and forming a clique on $W_{i}\cup\{x_{i,1},\ldots,x_{i,t_{i}}\}$. We prove the following.
Theorem 5.4. Let $G$ be a graph, and let $\pi=\{W_{1},\ldots,W_{k}\}$ be a clique vertex-partition of $G$. Then $\mathcal{I}_{r}(G^{\pi}_{r})$ is shellable for all $r\geq 1$.
Finally, we introduce clique cycle graphs, which generalize ordinary cycles by replacing edges with cliques. An $n$-clique cycle $\mathrm{CC}(B_{1},\ldots,B_{n})$ consists of a cyclic arrangement of $n\geq 3$ cliques in which each consecutive pair shares exactly one vertex. We consider graphs $G(r)$ constructed from an $n$-clique cycle by attaching a star-clique graph $\operatorname{SC}(x)$ to each vertex in a subset $A$ of the connecting vertices, requiring that $|V(\operatorname{SC}(x))|\geq r+1$ for at least one $x\in A$. Our final main result is the following.
Theorem 5.5. Let $G=G(r)$ for some $r\geq 1$. Then $\mathcal{I}_{r}(G)$ is shellable.
To conclude, we provide an example showing that the conclusions of Theorems 5.1, 5.4, and 5.5 may fail for $\operatorname{CCG}(H,S,t)$, $G^{\pi}_{t}$, or $G(t)$ when $t<r$ (Example 5.6).
Our paper is structured as follows. In Section 2, we introduce the necessary terminology and foundational results that underpin our work. Section 3 presents key technical lemmas and propositions essential for establishing our main results. In Section 4, we investigate the shellability of $\mathcal{I}_{r}(G)$, providing sufficient conditions for various classes of graphs. Finally, in Section 5, we examine the shellability of $r$-independence complexes in graph constructions.
2. Notation and Preliminaries
In this section, we establish the basic definitions and notation required for the main results. Let $G$ be a finite simple graph without isolated vertices. For a graph $G$, let $V(G)$ and $E(G)$ denote the vertex set and edge set of $G$, respectively. The degree of a vertex $x\in V(G)$, denoted by $\deg_{G}(x)$ (or simply $\deg(x)$), is the number of edges incident to $x$. A subgraph $H\subseteq G$ is called induced if for any two vertices $u,v\in V(H)$, the edge $\{u,v\}$ belongs to $E(H)$ if and only if $\{u,v\}\in E(G)$. For a subset $A\subseteq V(G)$, the induced subgraph of $G$ on $A$ is denoted by $G[A]$; that is, $G[A]$ is the subgraph of $G$ whose vertex set is $A$ and whose edge set consists of all edges of $G$ with both endpoints in $A$. For a subset ${u_{1},\ldots,u_{r}}\subseteq V(G)$, we define $N_{G}(u_{1},\ldots,u_{r})=\{v\in V(G)\mid\{u_{i},v\}\in E(G)\text{ for some }1\leq i\leq r\}$ and $N_{G}[u_{1},\ldots,u_{r}]=N_{G}(u_{1},\ldots,u_{r})\cup\{u_{1},\ldots,u_{r}\}$. For a subset $U\subseteq V(G)$, we denote by $G\setminus U$ the induced subgraph of $G$ on the vertex set $V(G)\setminus U$.
A path in $G$ is a sequence of distinct vertices $v_{0}v_{1}\cdots v_{k}$ such that $\{v_{i},v_{i+1}\}\in E(G)$ for all $0\leq i\leq k-1$. The length of a path is the number of edges it contains. A graph $G$ is called connected if for any two vertices $u,v\in V(G)$, there exists a path from $u$ to $v$. A connected component of $G$ is a maximal connected subgraph of $G$. The distance between two vertices $u$ and $v$ in $G$, denoted by $\operatorname{d}(u,v)$, is the length of the shortest path connecting $u$ and $v$. The diameter of $G$, denoted by $\operatorname{diam}(G)$, is the maximum distance between any pair of vertices in $G$. Formally, $\operatorname{diam}(G)=\max_{u,v\in V(G)}\operatorname{d}(u,v).$ A vertex $v\in V(G)$ is called a cut-vertex if the removal of $v$ from $G$ increases the number of connected components of $G$. A vertex $v\in V(G)$ is called a simplicial vertex if the subgraph induced by its neighbors is a complete graph (or clique); that is, for every pair of neighbors $u$ and $w$ of $v$, the edge $\{u,w\}\in E(G)$.
A simplicial complex $\Delta$ on a vertex set $V=\{x_{1},\ldots,x_{n}\}$ is a collection of subsets of $V$ satisfying two conditions: (i) every singleton $\{x_{i}\}$ belongs to $\Delta$ for each $x_{i}\in V$, and (ii) $\Delta$ is closed under taking subsets; that is, if $F\in\Delta$, then every subset $F^{\prime}\subseteq F$ also belongs to $\Delta$. The elements of $\Delta$ are called faces, and the maximal faces under inclusion are called facets. Several important constructions are associated with a face $F\in\Delta$: the link of $F$ is defined as $\operatorname{link}_{\Delta}(F)=\{F^{\prime}\subseteq V\mid F^{\prime}\cap F=\emptyset\text{ and }F^{\prime}\cup F\in\Delta\}$; the deletion of $F$ is $\Delta\setminus F=\{H\in\Delta\mid H\cap F=\emptyset\}$.
A hypergraph $\mathcal{H}$ is a pair $(V(\mathcal{H}),E(\mathcal{H}))$, where $V(\mathcal{H})$ is the vertex set and $E(\mathcal{H})\subseteq 2^{V(\mathcal{H})}$ is the edge set, consisting of subsets of $V(\mathcal{H})$ called the edges. A hypergraph is called simple if no edge is contained in another edge; that is, for any distinct edges $e,f\in E(\mathcal{H})$, we have $e\not\subseteq f$. In this paper, we consider all hypergraphs to be simple. A subset $S\subseteq V(\mathcal{H})$ is called independent if there is no $e\in E(\mathcal{H})$ such that $e\subseteq S$. The independence complex of a hypergraph $\mathcal{H}$, denoted $\mathcal{I}(\mathcal{H})$, is the simplicial complex whose simplices correspond to the independent subsets of $V(\mathcal{H})$. The deletion of the vertex $v$, denoted by $\mathcal{H}\setminus v$, is the hypergraph defined as: $V(\mathcal{H}\setminus v)=V(\mathcal{H})\setminus\{v\}$, $E(\mathcal{H}\setminus v)=\{e\in E(\mathcal{H})\mid v\notin e\}.$ The contraction of the vertex $v$, denoted by $\mathcal{H}/v$, is the hypergraph defined as:
$V(\mathcal{H}/v)=V(\mathcal{H})\setminus\{v\},\quad E(\mathcal{H}/v)=\min\bigl\{e\setminus\{v\}\mid e\in E(\mathcal{H})\bigr\},$
where $\min(S)$ denotes the inclusion-wise minimal subsets of a collection of sets $S$. Contractions and deletions of distinct vertices can be performed sequentially, and it is well known that the resulting hypergraph does not depend on the order of these operations [26, Lemma 3].
A hypergraph $\mathcal{H}^{\prime}$ obtained from $\mathcal{H}$ by deleting a subset of vertices $V_{d}\subseteq V(\mathcal{H})$ and contracting another subset $V_{c}\subseteq V(\mathcal{H})$, with $V_{d}\cap V_{c}=\emptyset$, is called a minor of $\mathcal{H}$, denoted by $\mathcal{H}\setminus V_{d}/V_{c}$. If $V_{d}=\emptyset$, the hypergraph $\mathcal{H}^{\prime}$ is called a $c$-minor of $\mathcal{H}$. Furthermore, if a $c$-minor $\mathcal{H}^{\prime}$ contains no singleton edges, it is called a $c^{\prime}$-minor of $\mathcal{H}$.
The concept of $\operatorname{W}$-chordality [^1] The notation honors Russ Woodroofe’s foundational work., introduced by Woodroofe, generalizes the notion of chordality from graphs to hypergraphs and provides a framework for studying shellability and linear resolutions.
Definition 2.1 ([33, Definition 4.2 and Definition 4.3]).
  1. (1)
    A vertex $v\in V(\mathcal{H})$ in a hypergraph $\mathcal{H}$ is called a simplicial vertex if for any two edges $e_{1},e_{2}\in E(\mathcal{H})$ containing $v$, there exists an edge $e_{3}\in E(\mathcal{H})$ such that $e_{3}\subseteq(e_{1}\cup e_{2})\setminus\{v\}.$
  2. (2)
    A hypergraph $\mathcal{H}$ is said to be $\operatorname{W}$-chordal if every minor of $\mathcal{H}$ contains a simplicial vertex.
These definitions lay the foundation for analyzing the shellability and sequentially Cohen–Macaulay properties of independence complexes.
Theorem 2.2 ([33, Corollary 5.3 and Corollary 5.4]).
  1. (1)
    Let $\mathcal{H}$ be a hypergraph such that every contraction of $\mathcal{H}$ has a simplicial vertex. Then $\mathcal{I}(\mathcal{H})$ is shellable, and therefore the Stanley–Reisner ideal of $\mathcal{I}(\mathcal{H})$ is sequentially Cohen–Macaulay.
  2. (2)
    If $\mathcal{H}$ is a $\operatorname{W}$-chordal hypergraph, then $\mathcal{I}(\mathcal{H})$ is shellable, and therefore the Stanley–Reisner ideal of $\mathcal{I}(\mathcal{H})$ is sequentially Cohen–Macaulay.
For any undefined terminology and further basic definitions and properties, we refer the reader to [19] and [31].
3. Technical Preliminaries and Key Results
In this section, we establish the technical groundwork necessary for the proof of our main theorem. We present a series of key lemmas and propositions that serve as the foundation for our arguments.
A collection $\mathcal{F}$ of graphs is called a hierarchy if, for every nonempty graph $G\in\mathcal{F}$, the graph $G\setminus u$ also belongs to $\mathcal{F}$ for every vertex $u\in V(G)$. The following proposition establishes a connection between the structural properties of a graph hierarchy and the $\operatorname{W}$-chordality of their associated hypergraphs. In particular, it provides a sufficient condition for $\mathcal{C}_{r}(G)$ to be $\operatorname{W}$-chordal, based on the existence of simplicial vertices in its minors.
Proposition 3.1.
Let $\mathcal{F}$ be a hierarchy of graphs and $r$ a positive integer. If every $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$ contains a simplicial vertex for any $G\in\mathcal{F}$, then $\mathcal{C}_{r}(G)$ is $\operatorname{W}$-chordal.
Proof.
We first show that every $c$-minor of $\mathcal{C}_{r}(G)$ contains a simplicial vertex for any $G\in\mathcal{F}$. Let $\mathcal{H}$ be a $c$-minor of $\mathcal{C}_{r}(G)$, i.e., $\mathcal{H}=\mathcal{C}_{r}(G)/V_{c}$ for some $V_{c}\subseteq V(G)$. We consider two cases based on the presence of singleton edges in $\mathcal{H}$: If $\mathcal{H}$ has no singleton edges, then $\mathcal{H}$ is a $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$. By assumption, $\mathcal{H}$ contains a simplicial vertex. Suppose $\mathcal{H}$ contains singleton edges. Let $S$ denote the set of all singleton edges in $\mathcal{H}$. Consider the hypergraph $\mathcal{H}\setminus S=(\mathcal{C}_{r}(G)/V_{c})\setminus S.$ By [26, Lemma 3], we have $\mathcal{H}\setminus S=(\mathcal{C}_{r}(G)\setminus S)/V_{c}.$ Additionally, for any vertex $x\in V(G)$ and $r\geq 1$, we observe that $\mathcal{C}_{r}(G)\setminus x=\mathcal{C}_{r}(G\setminus x).$ Extending this identity to the set $S$, we obtain $\mathcal{H}\setminus S=\mathcal{C}_{r}(G\setminus S)/V_{c}.$ Since $G\setminus S\in\mathcal{F}$ (as $\mathcal{F}$ is a hierarchy), and $\mathcal{H}\setminus S$ is a $c^{\prime}$-minor of $\mathcal{C}_{r}(G\setminus S)$, the hypothesis implies that $\mathcal{H}\setminus S$ contains a simplicial vertex. Thus, $\mathcal{H}$ must also contain a simplicial vertex. In both cases, we conclude that every $c$-minor of $\mathcal{C}_{r}(G)$ contains a simplicial vertex. Since any minor can be obtained through deletions and contractions, and deletions correspond to graphs in the hierarchy by definition, it follows that all minors of $\mathcal{C}_{r}(G)$ possess a simplicial vertex. Hence, $\mathcal{C}_{r}(G)$ is $\operatorname{W}$-chordal for every $G\in\mathcal{F}$. ∎
In [10, Theorem 4.3], Dao and Schweig proved that a vertex $v$ of a graph $G$ is simplicial if and only if it remains simplicial in $\mathcal{C}_{r}(G)$ for all $r\geq 1$. However, this equivalence does not hold for general graphs. For example, consider the cycle graph $C_{4}$: while no vertex in $C_{4}$ is simplicial, every vertex in $\mathcal{C}_{r}(C_{4})$ becomes simplicial for $r=2$ (or $r=3$ under their construction). This demonstrates that the claimed equivalence fails in general. Motivated by this observation, we examine the persistence of simplicial vertices in minors of the hypergraph $\mathcal{C}_{r}(G)$.
Lemma 3.2.
Let $G$ be a graph, and let $\mathcal{H}$ be a minor of the hypergraph $\mathcal{C}_{r}(G)$ for some $r\geq 1$. If $x$ is a simplicial vertex of $G$ and $x\in V(\mathcal{H})$, then $x$ remains a simplicial vertex in $\mathcal{H}$.
Proof.
Suppose first that $x$ belongs to exactly one edge of $\mathcal{H}$; then $x$ is simplicial by definition. Now consider the case where $x$ belongs to two distinct edges $e_{1},e_{2}\in E(\mathcal{H})$, with $e_{1}\neq e_{2}$. By construction of the minor, there exist edges $f_{1},f_{2}\in E(\mathcal{C}_{r}(G))$ such that $e_{i}=f_{i}\cap V(\mathcal{H})$ for $i=1,2.$ Since $x$ is simplicial in $G$, the induced subgraph of $G$ on $(f_{1}\cup f_{2})\setminus\{x\}$ is connected. Moreover, the fact that $f_{1}\neq f_{2}$ implies that $|f_{1}\cup f_{2}|\geq r+2,$ and hence, $|(f_{1}\cup f_{2})\setminus\{x\}|\geq r+1.$ Select a subset $f\subseteq(f_{1}\cup f_{2})\setminus\{x\}$ of size $r+1$ such that the induced subgraph on $f$ is connected (which is possible since the entire set is connected and of size at least $r+1$). Define $e_{3}:=f\cap V(\mathcal{H}).$ Clearly, $e_{3}\subseteq e_{1}\cup e_{2}\setminus\{x\}$. If $e_{3}\in E(\mathcal{H})$, we are done. Otherwise, there exists $e_{3}^{\prime}\in E(\mathcal{H})$ such that $e_{3}^{\prime}\subseteq e_{3}\subseteq e_{1}\cup e_{2}\setminus\{x\}$. This proves that $x$ is a simplicial vertex in $\mathcal{H}$. ∎
A graph is called a star-clique graph, denoted $\operatorname{SC}(x)$, if it is formed by attaching a collection of cliques $B_{1},\ldots,B_{t}$ to a central vertex $x$, where each $B_{i}$ is a complete graph (clique) that may vary in size.
Corollary 3.3.
If $G$ is a star-clique graph, then $\mathcal{C}_{r}(G)$ is a $\operatorname{W}$-chordal hypergraph for all $r\geq 1$.
Proof.
Let $V_{1}$ be the set of all simplicial vertices of $G$. Consider a minor $\mathcal{H}$ of the hypergraph $\mathcal{C}_{r}(G)$. If $V_{1}\cap V(\mathcal{H})\neq\emptyset$, then by Lemma 3.2, $\mathcal{H}$ contains a simplicial vertex. If $V_{1}\cap V(\mathcal{H})=\emptyset$, then all simplicial vertices of $G$ are excluded from $\mathcal{H}$, which can only happen if $\mathcal{H}$ has exactly one vertex. In that case, $\mathcal{H}$ trivially contains a simplicial vertex. Therefore, every minor $\mathcal{H}$ of $\mathcal{C}_{r}(G)$ has a simplicial vertex, which implies that $\mathcal{C}_{r}(G)$ is $\operatorname{W}$-chordal. ∎
The following observation highlights a key property of chordal graphs, leveraging the structure of simplicial vertices and their iterative removal.
Observation 3.4.
Let $G$ be a chordal graph with $|V(G)|\geq 2$. A fundamental result of Dirac [13] states that every chordal graph contains at least one simplicial vertex, and any chordal graph that is not complete contains at least two non-adjacent simplicial vertices.
Define $V_{1}$ as the set of all simplicial vertices of $G$. By Dirac’s theorem, $|V_{1}|\geq 2$ unless $G$ is complete (in which case $|V_{1}|=|V(G)|\geq 2$). For $i\geq 2$, recursively define $V_{i}$ as the set of simplicial vertices in the subgraph $G\setminus\bigcup_{j=1}^{i-1}V_{j}$. Since chordality is preserved under induced subgraphs, each $G\setminus\bigcup_{j=1}^{i-1}V_{j}$ remains chordal, and thus each $V_{i}$ is non-empty. This process terminates at some step $\kappa\geq 1$ when $G\setminus\bigcup_{j=1}^{\kappa}V_{j}$ becomes empty. Moreover, for all $1\leq i<\kappa$, we have $|V_{i}|\geq 2$ (unless the remaining subgraph is complete, in which case all remaining vertices are simplicial). The sets ${V_{i}}_{i=1}^{\kappa}$ form a partition of $V(G)$ into pairwise disjoint subsets.
The following observation explores the structure of block graphs, a subclass of chordal graphs. Using the framework established in Observation 3.4, we analyze the cliques associated with vertices in $V_{2}$ and derive a key property regarding their structure and size.
Observation 3.5.
Let $G$ be a block graph. Since every block graph is chordal, we apply the framework introduced in Observation 3.4. For each vertex $v\in V_{2}$, there exists a maximum-sized clique in $G$ that contains $v$, such that all other vertices in this clique belong to $V_{1}$. Denote these cliques by $C_{1},C_{2},\ldots,C_{m}$, where $m\geq 1$. Define $S_{v}=\bigcup_{i=1}^{m}\left(C_{i}\setminus\{v\}\right).$ As a consequence, we observe that $|S_{v}|\geq 1$.
We now define an $n$-clique path graph, denoted $\operatorname{CP}(B_{1},\ldots,B_{n})$, as a graph consisting of a sequence of $n$ cliques $B_{1},\ldots,B_{n}$, arranged linearly, with $n\geq 1$. Each consecutive pair of cliques $B_{i}$ and $B_{i+1}$ shares exactly one vertex, denoted $x_{i}$. The vertices $x_{1},\ldots,x_{n-1}$ are referred to as the connecting vertices of the $n$-clique path graph.
$B_{1}$$B_{2}$$B_{3}$$B_{4}$$B_{5}$$x_{1}$$x_{2}$$x_{3}$$x_{4}$
$\operatorname{CP}(B_{1},B_{2},B_{3},B_{4},B_{5})$
For example, the graph shown left is a 5-clique path graph $\operatorname{CP}(B_{1},B_{2},B_{3},B_{4},B_{5})$, where $B_{1}$ is the clique on 4 vertices, $B_{2}$ is the clique on 3 vertices, $B_{3}$ is the clique on 2 vertices, $B_{4}$ is the clique on 6 vertices, and $B_{5}$ is the clique on 3 vertices. The vertices $x_{1},x_{2},x_{3},x_{4}$ are the connecting vertices, which link each consecutive pair of cliques in this path.
The following setup introduces key notation and assumptions that will be used in subsequent results. Using Observation 3.4, we define a parameter $\ell$ that plays a crucial role in the analysis.
Set-up 3.6.
Let $G$ be a connected block graph. Then there exists a maximum-sized $n$-clique path, denoted $\operatorname{CP}(B_{1},\ldots,B_{n})$, as an induced subgraph with connecting vertices $x_{1},x_{2},\ldots,x_{n-1}$, where $n\geq 1$. Let $\mathcal{H}$ be a $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$ for some $r\geq 1$, and assume that $|E(\mathcal{H})|\geq 3$. Using Observation 3.4 and its associated notation, we define the parameter $\ell$ as: $\ell=\min\left\{i\mid V_{i}\cap V(\mathcal{H})\neq\emptyset\right\}.$
Lemma 3.7.
Assume the notation from Set-Up 3.6, and suppose $\ell\neq 1$ and $r<n$. If $x_{0},\ldots,x_{r-2}\notin V(\mathcal{H})$, where $x_{0}\in V(B_{1})\cap V_{1}$, then $x_{r-1}\in V(\mathcal{H})$. Furthermore, if $y\in N_{G}(x_{r-1})\setminus\{x_{r-2}\}$, then $y\in V(\mathcal{H})$.
Proof.
Suppose, for contradiction, that $x_{r-1}\notin V(\mathcal{H})$. Choose a vertex $c\in V(\mathcal{H})$ such that $\mathrm{d}(c,x_{r-1})=\min\{\mathrm{d}(z,x_{r-1})\mid z\in V(\mathcal{H})\}.$ Set $\mathrm{d}(c,x_{r-1})=t$. Then, there exists a shortest path $(x_{r-1}=y_{0})y_{1}\cdots y_{t-1}(y_{t}=c)$ from $x_{r-1}$ to $c$. Since $c$ is the closest vertex in $V(\mathcal{H})$, all intermediate vertices $y_{1},\ldots,y_{t-1}$ must lie outside $V(\mathcal{H})$. Define $S=\{x_{0},\ldots,x_{r-1},y_{1},\ldots,y_{t}\}.$ Clearly, $|S|\geq r+1$ and $S$ is connected. Moreover, $S\cap V(\mathcal{H})=\{c\}$. Since $c\in V(\mathcal{H})$, any $f\in E(\mathcal{C}_{r}(G))$ such that $f\subseteq S$ and $c\in f$ would imply $f\cap V(\mathcal{H})=\{c\}\in E(\mathcal{H})$, contradicting the assumption that $\mathcal{H}$ is a $c^{\prime}$-minor (i.e., no edge of $\mathcal{C}_{r}(G)$ intersects $V(\mathcal{H})$ in exactly one vertex). Hence, we must have $x_{r-1}\in V(\mathcal{H})$.
Now suppose $y\in N_{G}(x_{r-1})\setminus\{x_{r-2}\}$, and assume, for contradiction, that $y\notin V(\mathcal{H})$. Then the set $T=\{x_{0},\ldots,x_{r-1},y\}$ is connected. Since $x_{r-1}\in V(\mathcal{H})$ and all other elements of $T$ lie outside $V(\mathcal{H})$, we have $T\cap V(\mathcal{H})=\{x_{r-1}\}.$ Thus, $T$ defines an edge in $\mathcal{C}_{r}(G)$ whose intersection with $V(\mathcal{H})$ is exactly $\{x_{r-1}\}\in E(\mathcal{H})$, again contradicting the $c^{\prime}$-minor condition. Therefore, $y\in V(\mathcal{H})$. ∎
Lemma 3.8.
Assume the setup in 3.6 and the notation introduced in Observation 3.5. Let $\ell=2$, and suppose that $v_{k}\in V_{2}\cap V(\mathcal{H})$ and $v_{k}\in e_{1}\cap e_{2}$, where $e_{1},e_{2}\in E(\mathcal{H})$ and $e_{1}\neq e_{2}$. Then, there exist edges $f_{1},f_{2}\in E(\mathcal{C}_{r}(G))$ such that the following conditions hold:
$S_{v_{k}}\subseteq f_{i},\quad f_{i}\cap V(\mathcal{H})=e_{i}\quad\text{for }i=1,2,$
and $f_{1}\cup f_{2}\setminus\left(S_{v_{k}}\cup\{v_{k}\}\right)$ is connected.
Proof.
Let $S_{v_{k}}=S$. We first claim that there exists $f_{1}\in E(\mathcal{C}_{r}(G))$ such that $S\subseteq f_{1}$ and $f_{1}\cap V(\mathcal{H})=e_{1}$. Since $\mathcal{H}$ is a $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$, there exists an edge $f^{\prime}\in E(\mathcal{C}_{r}(G))$ such that $f^{\prime}\cap V(\mathcal{H})=e_{1}$. If $S\subseteq f^{\prime}$, the claim is true. Otherwise, since every vertex of $S$ is connected to $v_{k}$, and $v_{k}\in f^{\prime}$, the set $S\cup f^{\prime}$ is connected, and we have $|S\cup f^{\prime}|\geq r+1$. We can then choose a connected subset $f_{1}$ of $S\cup f^{\prime}$ such that $S\subseteq f_{1}$ and $|f_{1}|=r+1$. Consequently, we have: $f_{1}\cap V(\mathcal{H})\subseteq f^{\prime}\cap V(\mathcal{H})=e_{1}.$ If $f_{1}\cap V(\mathcal{H})\subset e_{1}$, this would contradict the assumption that $e_{1}\in E(\mathcal{H})$, since $f_{1}\cap V(\mathcal{H})$ would then be a proper subset of $e_{1}$. Therefore, it must be that $f_{1}\cap V(\mathcal{H})=e_{1}$. Similarly, there exists $f_{2}\in E(\mathcal{C}_{r}(G))$ such that $S\subseteq f_{2}$ and $f_{2}\cap V(\mathcal{H})=e_{2}.$
Next, we claim that there exists $z\in f_{1}\setminus(S\cup\{v_{k}\})$ such that $z\notin V_{1}$ and $z\in N_{G}(v_{k})$. There must exist a vertex $t\in N_{G}(v_{k})$ such that $t\notin V_{1}$, because if every vertex in $N_{G}(v_{k})$ were in $V_{1}$, then $v_{k}$ would be in $V_{1}$, contradicting the assumption that $v_{k}\in V_{2}$. Since $e_{1}\neq e_{2}$, there exists $x_{1}^{\prime}\in e_{1}\setminus e_{2}$, and hence $x_{1}^{\prime}\in f_{1}$. Consider the shortest path $(v_{k}=a_{0})a_{1}\cdots a_{n-1}(a_{n}=x_{1}^{\prime})$ in $G$, where $a_{0},\dots,a_{n}\in f_{1}$. Then, $a_{1}\in N_{G}(v_{k})$, and if $a_{1}\in V_{1}$, $a_{1}$ would be a simplicial vertex, implying the path $v_{k}a_{2}\cdots a_{n-1}(a_{n}=x_{1}^{\prime})$ would be a shorter path, leading to a contradiction. Therefore, $a_{1}\notin V_{1}$, and hence $a_{1}\notin S\cup\{v_{k}\}$. This proves the claim. Similarly, there exists $z\in f_{2}\setminus(S\cup\{v_{k}\})$ such that $z\notin V_{1}$ and $z\in N_{G}(v_{k})$.
Finally, we claim that $f_{1}\cup f_{2}\setminus(S\cup\{v_{k}\})$ is connected. By the previous argument, there exist vertices $z_{1}\in f_{1}\setminus(S\cup\{v_{k}\})$ and $z_{2}\in f_{2}\setminus(S\cup\{v_{k}\})$, both of which are not in $V_{1}$ and are in $N_{G}(v_{k})$. If $z_{1}=z_{2}$, the claim follows immediately. Otherwise, if $z_{1}\neq z_{2}$, there must be an edge between $z_{1}$ and $z_{2}$ because $v_{k}\in V_{2}$, which ensures that $z_{1}$ and $z_{2}$ are connected. Therefore, $f_{1}\cup f_{2}\setminus(S\cup\{v_{k}\})$ is connected, and the proof is complete. ∎
Notation 3.9.
Assume the setup in 3.6 and the notation introduced in Observation 3.5. Let $\ell=2$, and suppose that $v_{k}\in V_{2}\cap V(\mathcal{H})$ and $v_{k}\in e_{1}\cap e_{2}$, where $e_{1},e_{2}\in E(\mathcal{H})$ and $e_{1}\neq e_{2}$. By Lemma 3.8, there exist edges $f_{1},f_{2}\in E(\mathcal{C}_{r}(G))$ such that:
$S_{v_{k}}\subseteq f_{i},\quad f_{i}\cap V(\mathcal{H})=e_{i},\quad\text{for }i=1,2,$
and the set $f_{1}\cup f_{2}\setminus(S_{v_{k}}\cup\{v_{k}\})$ induces a connected subgraph. Next, fix an element $x_{1}^{\prime}\in e_{1}\setminus e_{2}$. For each $1\leq i\leq\gamma$, where $\gamma\geq 1$, define a set $C_{i}=\{c_{i,1},\ldots,c_{i,q_{i}}\}\subseteq V(G),$ satisfying the following conditions:
[…]
Finally, define $C=\bigcup_{i=1}^{\gamma}C_{i}.$
Lemma 3.10.
Let the notation be as in Notation 3.9. Then $v_{k}$ is a simplicial vertex of $\mathcal{H}$ in either of the following cases: if $|f_{1}\cup f_{2}\setminus(S_{v_{k}}\cup\{v_{k}\})|\geq r+1$, or if $|f_{1}\cup f_{2}\setminus(S_{v_{k}}\cup\{v_{k}\})|\leq r$ and $|C|\geq|S_{v_{k}}|$.
Proof.
(1) Suppose that $\left|(f_{1}\cup f_{2})\setminus(S_{v_{k}}\cup\{v_{k}\})\right|\geq r+1$. Then, there exists a connected subset $f\subseteq(f_{1}\cup f_{2})\setminus(S_{v_{k}}\cup\{v_{k}\})$ with $|f|=r+1$. Let $e_{3}=f\cap V(\mathcal{H})$; since $f$ avoids $\{v_{k}\}$, we have $e_{3}\subseteq(e_{1}\cup e_{2})\setminus\{v_{k}\}$. If $e_{3}$ is itself an edge in $E(\mathcal{H})$, then $v_{k}$ must be a simplicial vertex. Otherwise, by the properties of the hypergraph $\mathcal{H}$, there exists a subset $e_{3}^{\prime}\in E(\mathcal{H})$ such that $e_{3}^{\prime}\subseteq e_{3}$, which again implies that $v_{k}$ is simplicial. In either case, the conclusion follows.
(2) Now assume $|A|\leq r$, where $A=f_{1}\cup f_{2}\setminus(S_{v_{k}}\cup\{v_{k}\})$. Let $S=S_{v_{k}}$ and define $D^{\prime}=f_{2}\setminus(S\cup\{v_{k}\})$. We claim that $D^{\prime}$ is a connected set. Let $\alpha,\beta\in D^{\prime}$. If $\alpha$ and $\beta$ are not adjacent, since $\alpha,\beta\in f_{2}$, there exists a shortest path $\alpha\ y_{1}\cdots y_{p}\ \beta$ within $f_{2}$, where each $y_{i}\in f_{2}$ for $1\leq i\leq p$. Since $y_{i}$ is not a simplicial vertex of $G$ for all $1\leq i\leq p$, we have $y_{i}\notin S\cup\{v_{k}\}$, ensuring that $D^{\prime}$ remains connected.
Next, we claim that $D^{\prime}\cap C=\emptyset$. Suppose, for contradiction, that there exists $c_{i,j}\in D^{\prime}\cap C$, where $c_{i,j}\in C_{i}$ for some $1\leq i\leq\gamma$, $1\leq j\leq q_{i}$. Since $v_{k},x_{1}^{\prime}\in f_{1}$, there exists a shortest path $(v_{k}=z_{0})z_{1}\cdots z_{t-1}(z_{t}=x_{1}^{\prime})$ in $G$, where $z_{i}\in f_{1}$ for all $0\leq i\leq t$. Let $B$ be the maximal clique containing $z_{t-1}$ and $x_{1}^{\prime}$. Similarly, since $v_{k},c_{i,j}\in f_{2}$, there exists a shortest path $P(v_{k},c_{i,j})$ within $f_{2}$. Since $c_{i,j}\in C_{i}$ and conditions (i) and (ii) from (3.1) hold, there exists a shortest path $x_{1}^{\prime}a_{1}\cdots a_{m}c_{i,j}$, where $a_{p}\in C_{i}$ for all $1\leq p\leq m$. We observe that $a_{1}\notin V(B)$, since otherwise, $\operatorname{d}(v_{k},a_{1})\leq\operatorname{d}(v_{k},x_{1}^{\prime})$, contradicting condition (iv) from (3.1). Since $x_{1}^{\prime}$ is a cut vertex, $a_{1}$ and $z_{t-1}$ belong to different components of $G\setminus\{x_{1}^{\prime}\}$. However, the path $z_{t-1}\cdots v_{k}\ P(v_{k},c_{i,j})\ a_{m}\cdots a_{1}$ remains in $G\setminus\{x_{1}^{\prime}\}$, which is a contradiction. Therefore, $D^{\prime}\cap C=\emptyset$.
Since $D^{\prime}\cap\{x_{1}^{\prime}\}=\emptyset$ (because $x_{1}^{\prime}\in e_{2}$), we have $|f_{2}|=r+1=|D^{\prime}|+|S|+1.$ Now, consider $D^{\prime}\cup R\cup C$, where $R=\{z_{1},\ldots,z_{t-1},z_{t}\}$. We can show that $D^{\prime}\cup R\cup C$ is connected. By computing its size, we obtain $|D^{\prime}\cup R\cup C|\geq r+1.$ Select a connected subset $f\subseteq D^{\prime}\cup R\cup C$ with $|f|=r+1$, and define $e=f\cap V(\mathcal{H}).$ This ensures that $e\subseteq e_{1}\cup e_{2}\setminus\{v_{k}\}$. If $e\in E(\mathcal{H})$, then $v_{k}$ is a simplicial vertex. Otherwise, there exists $e_{s}\in E(\mathcal{H})$ such that $e_{s}\subseteq e$, proving that $v_{k}$ is a simplicial vertex. ∎
4. Shellability of Higher Independence Complexes
This section establishes sufficient conditions on $r$ and on the graph $G$ under which the $r$-independence complex $\mathcal{I}_{r}(G)$ is shellable. We begin by examining a claim from [1], where it was asserted that for any tree $G$, the hypergraph $\mathcal{C}_{r}(G)$ is $W$-chordal for all $r\geq 2$ (see [1, Theorem 2.7]). However, the proof presented there contains a flaw. In fact, we exhibit an explicit family of trees for which $\mathcal{C}_{r}(G)$ is not $W$-chordal. Moreover, for each $r\geq 4$, we construct a family of graphs $G_{t}$ such that the hypergraph $\mathcal{C}_{r}(G_{t})$ fails to be $W$-chordal. The construction is given below.
Construction 1.
Let $H_{1}$, $H_{2}$, and $H_{3}$ be connected graphs with pairwise disjoint vertex sets, each satisfying $|V(H_{i})|\geq t$ for $1\leq i\leq 3$. Define a base graph $G$ with vertex set $V(G)=\{v_{1},v_{2},v_{3},v_{4},v_{5},v_{6},v_{7}\}$ and edge set
$E(G)=\{\{v_{1},v_{2}\},\{v_{2},v_{3}\},\{v_{1},v_{4}\},\{v_{4},v_{5}\},\{v_{1},v_{6}\},\{v_{6},v_{7}\}\}.$
$H_{1}$$H_{2}$$H_{3}$$v_{1}$$v_{2}$$v_{4}$$v_{6}$$v_{3}$$v_{5}$$v_{7}$$u_{1}$$u_{2}$$u_{3}$
$G_{t}$
The graph $G_{t}$ is formed by taking the disjoint union of $G$, $H_{1}$, $H_{2}$, and $H_{3}$, and then attaching each $H_{i}$ to $G$ via a single edge. Specifically, let $u_{1}\in V(H_{1})$, $u_{2}\in V(H_{2})$, and $u_{3}\in V(H_{3})$ be arbitrary vertices. Then the vertex and edge sets of $G_{t}$ are given by $V(G_{t})=V(G)\cup V(H_{1})\cup V(H_{2})\cup V(H_{3}),$ $E(G_{t})=E(G)\cup E(H_{1})\cup E(H_{2})\cup E(H_{3})\cup\{\{v_{3},u_{1}\},\{v_{5},u_{2}\},\{v_{7},u_{3}\}\}.$
Proposition 4.1.
If $t\geq r-3$ for some $r\geq 4$, then the hypergraph $\mathcal{C}_{r}(G_{t})$ is not $\operatorname{W}$-chordal.
Proof.
For each $1\leq i\leq 3$, choose a subset $S_{i}\subseteq V(H_{i})$ with $|S_{i}|=r-3$ such that the following subsets of $V(G_{t})$ belong to $E(\mathcal{C}_{r}(G_{t}))$: $S_{1}\cup\{v_{1},v_{2},v_{3},v_{4}\}$, $S_{1}\cup\{v_{1},v_{2},v_{3},v_{6}\}$, $S_{2}\cup\{v_{5},v_{4},v_{1},v_{6}\}$, $S_{2}\cup\{v_{5},v_{4},v_{1},v_{2}\}$, $S_{3}\cup\{v_{7},v_{6},v_{1},v_{4}\}$, and $S_{3}\cup\{v_{7},v_{6},v_{1},v_{2}\}$.
Let $S^{\prime}_{i}=V(H_{i})\setminus S_{i}$, and define $S=S^{\prime}_{1}\cup S^{\prime}_{2}\cup S^{\prime}_{3}$. Let $\mathcal{H}$ be the hypergraph obtained from $\mathcal{C}_{r}(G_{t})$ by deleting all vertices in $S$ and contracting the vertices in $S_{1}\cup S_{2}\cup S_{3}$. Since vertex deletion and contraction behave compatibly with $\mathcal{C}_{r}(G)$ (that is, $\mathcal{C}_{r}(G)\setminus x=\mathcal{C}_{r}(G\setminus x)$ for any vertex $x\in V(G)$), the resulting hypergraph $\mathcal{H}$ has vertex set $V(\mathcal{H})=\{v_{1},v_{2},v_{3},v_{4},v_{5},v_{6},v_{7}\}$ and edge set $E(\mathcal{H})=$ $\{\{v_{1},v_{2},v_{3},v_{4}\}$, $\{v_{1},v_{2},v_{3},v_{6}\}$, $\{v_{5},v_{4},v_{1},v_{6}\}$, $\{v_{5},v_{4},v_{1},v_{2}\}$, $\{v_{7},v_{6},v_{1},v_{4}\}$, $\{v_{7},v_{6},v_{1},v_{2}\}\}$. One can verify that $\mathcal{H}$ contains no simplicial vertices. Therefore, $\mathcal{H}$ is not $\operatorname{W}$-chordal, and consequently, $\mathcal{C}_{r}(G_{t})$ is not $\operatorname{W}$-chordal. ∎
Remark 4.2.
In [1], it was claimed that for any tree $G$, the hypergraph $\mathcal{C}_{r}(G)$ is $\operatorname{W}$-chordal for all $r\geq 2$ (see [1, Theorem 2.7]), and consequently that the $r$-independence complex $\mathcal{I}_{r}(G)$ is shellable. However, the argument provided there appears to be incomplete. Proposition 4.1 shows that for every $r\geq 4$, there exists a tree $G$ (constructed by taking each $H_{i}$ to be a tree for $i=1,2,3$) such that $\mathcal{C}_{r}(G)$ is not $\operatorname{W}$-chordal. This provides a counterexample to the claim and calls into question the corresponding shellability conclusion.
The following theorem provides a sufficient condition for the $r$-independence complex $\mathcal{I}_{r}(G)$ to be shellable when $r$ is close to the order of the graph.
Theorem 4.3.
Let $G=(V(G),E(G))$ be a graph on $n=|V(G)|\geq 3$ vertices. If $n-2\leq r\leq n-1$, then $\mathcal{I}_{r}(G)$ is shellable.
Proof.
By Theorem 2.2, it is enough to prove that $\mathcal{C}_{r}(G)$ is $\operatorname{W}$-chordal. If r=n1r=n-1, then each edge of $\mathcal{C}_{r}(G)$ has size $n$, which implies that there can be at most one such edge. Hence, $\mathcal{C}_{r}(G)$ is trivially $\operatorname{W}$-chordal. Now suppose r=n2r=n-2. Then each edge in $\mathcal{C}_{r}(G)$ has size n1n-1. If $|E(\mathcal{C}_{r}(G))|\leq 2$, then again the hypergraph is trivially $\operatorname{W}$-chordal. Assume $|E(\mathcal{C}_{r}(G))|\geq 3$. We first show that for any two distinct edges $f_{1},f_{2}\in\mathcal{C}_{r}(G)$, we have $|f_{1}\cap f_{2}|=n-2$. Since $|f_{1}|=|f_{2}|=n-1$, we have $|f_{1}\cup f_{2}|=2(n-1)-|f_{1}\cap f_{2}|.$ But $f_{1}\cup f_{2}\subseteq V(G)$, so $2(n-1)-|f_{1}\cap f_{2}|\leq n\Rightarrow|f_{1}\cap f_{2}|\geq n-2.$ Also, since $f_{1}\neq f_{2}$, we cannot have $|f_{1}\cap f_{2}|=n-1$. Hence, $|f_{1}\cap f_{2}|=n-2$, and so $f_{1}\cup f_{2}=V(G)$.
Now let $\mathcal{H}$ be any minor of $\mathcal{C}_{r}(G)$. We show that $\mathcal{H}$ contains a simplicial vertex. If $|E(\mathcal{H})|\leq 2$, the result is immediate. Assume $|E(\mathcal{H})|\geq 3$. Let $e_{1},e_{2}\in E(\mathcal{H})$ with $e_{1}\neq e_{2}$, and choose vertices $x_{1}^{\prime}\in e_{1}\setminus e_{2}$ and $x_{2}^{\prime}\in e_{2}\setminus e_{1}$. We claim that $x_{1}^{\prime}$ is a simplicial vertex in $\mathcal{H}$. Let $e_{1}^{\prime}$ and $e_{2}^{\prime}$ be two distinct edges in $E(\mathcal{H})$ such that their intersection $e_{1}^{\prime}\cap e_{2}^{\prime}$ contains the vertex $x_{1}^{\prime}$. Since $x_{1}^{\prime}\in e_{1}\setminus e_{2}$, it follows that $e_{1}^{\prime},e_{2}^{\prime}\neq e_{2}$. Let $f_{1}^{\prime},f_{2}^{\prime},g_{2}\in E(\mathcal{C}_{r}(G))$ be preimages of $e_{1}^{\prime},e_{2}^{\prime},e_{2}$ under the minor operation, so that $f_{i}^{\prime}\cap V(\mathcal{H})=e_{i}^{\prime}$ for $i=1,2$ and $g_{2}\cap V(\mathcal{H})=e_{2}$. From earlier, we know $f_{1}^{\prime}\cup f_{2}^{\prime}=V(G)$, and we have $e_{2}=g_{2}\cap V(\mathcal{H})\subseteq(f_{1}^{\prime}\cup f_{2}^{\prime})\cap V(\mathcal{H})=e_{1}^{\prime}\cup e_{2}^{\prime}\subseteq V(\mathcal{H}).$ Since $x_{1}^{\prime}\notin e_{2}$, this implies $e_{2}\subseteq(e_{1}^{\prime}\cup e_{2}^{\prime})\setminus\{x_{1}^{\prime}\}$. Hence, $x_{1}^{\prime}$ is a simplicial vertex of $\mathcal{H}$. Therefore, every minor of $\mathcal{C}_{r}(G)$ contains a simplicial vertex, and so $\mathcal{C}_{r}(G)$ is $\operatorname{W}$-chordal. ∎
In the next lemma, we establish a connection between a cut vertex of the tree $G$ and a $c^{\prime}$-minor of the graph $\mathcal{C}_{r}(G)$, where $r\geq 1$.
Lemma 4.4.
Let $G$ be a tree, and let $\mathcal{H}$ be a $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$ for some $r\geq 1$. Following the notation in Set-Up 3.6, assume that $\ell\neq 1$. Then there exists a cut vertex $v$ of $G$ such that $v\in V(\mathcal{H})$ and $V(\mathcal{H})\setminus\{v\}\subseteq V(G_{1})$, where $G_{1}$ is a connected component of $G\setminus\{v\}$.
Proof.
Let $v\in V_{\ell}\cap V(\mathcal{H})$. Since $\mathcal{H}$ is a $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$, the vertex $v$ cannot appear as a singleton edge in $\mathcal{H}$, i.e., $\{v\}\notin E(\mathcal{H})$. Hence, $v$ must be adjacent to some vertex in $G\setminus\bigcup_{i=1}^{\ell-1}V_{i}$. As $v$ is simplicial in the induced subgraph $G\setminus\bigcup_{i=1}^{\ell-1}V_{i}$, its neighborhood in this subgraph is a singleton: $N_{G\setminus\bigcup_{i=1}^{\ell-1}V_{i}}(v)=\{a\},$ for some $a\in V(G)$. Since $G$ is a tree and $v$ is a cut vertex, the deletion of $v$ produces at least two connected components. Let $G_{1}$ be the component containing $a$. Now consider any vertex $x\in V(\mathcal{H})\setminus\{v\}$. Because $G$ is a tree, there exists a unique path from $v$ to $x$, say $vz_{1}z_{2}\cdots z_{t}x.$ Since $v\in V_{\ell}$, there exists a path $a_{1}a_{2}\cdots a_{\ell-1}v$ with $a_{i}\in V_{i}$ for $1\leq i\leq\ell-1$. Moreover, as $x\in V(\mathcal{H})\setminus\{v\}$ and $x\in V_{j}$ for some $j\geq\ell$, there exists a path $xb_{1}b_{2}\cdots b_{r}$ with $r\geq\ell-1$.
Let $z_{i}$ be an internal vertex on the path from $v$ to $x$. Then both $a_{1}a_{2}\cdots a_{\ell-1}vz_{1}\cdots z_{i-1}z_{i}$ and $z_{i}z_{i+1}\cdots z_{t}xb_{1}\cdots b_{r}$ are paths of length at least $\ell$. Consequently, $z_{i}\notin V_{s}$ for $1\leq s\leq\ell$, which means all internal vertices lie outside $\bigcup_{s=1}^{\ell}V_{s}$. Since $a$ is the unique neighbor of $v$ in $G\setminus\bigcup_{s=1}^{\ell-1}V_{s}$, the first edge on the path from $v$ to $x$ must be $vz_{1}=va$, so $z_{1}=a$. The remainder of the path connects $a$ to $x$ within $G\setminus\{v\}$, which shows that $x$ lies in the same connected component as $a$, namely $G_{1}$. As $x\in V(\mathcal{H})\setminus\{v\}$ was arbitrary, we conclude that $V(\mathcal{H})\setminus\{v\}\subseteq V(G_{1}),$ as desired. ∎
We now extend Theorem 4.3 to establish a lower bound in the case when $G$ is a forest. Let $G$ be a tree, and let $\mathcal{H}$ be a $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$ for some $r\geq 1$. Consider the cut vertex $v$ of $G$ as given by Lemma 4.4, and let $G_{1}$ be the component of $G\setminus\{v\}$ that contains $V(\mathcal{H})\setminus\{v\}$. Define $C_{v}=V(G)\setminus(V(G_{1})\cup\{v\})$, ensuring that $C_{v}\cap V(\mathcal{H})=\emptyset$.
Theorem 4.5.
Let $G$ be a forest. Then, for all $r\geq n-5$, the complex $\mathcal{I}_{r}(G)$ is shelable.
Proof.
[…]
We now claim that either $x_{1}$ or $x_{2}$ belongs to $S$. Suppose, for contradiction, that neither $x_{1}$ nor $x_{2}$ belongs to $S$. For $i=1,2$, define
[…]
Note that $D_{1}\cap f_{2}=\emptyset$ and $D_{2}\cap f_{1}=\emptyset$; otherwise, $x_{1}\in e_{2}$ or $x_{2}\in e_{1}$, contradicting the choice of $x_{1}$ and $x_{2}$. If $D_{i}\cap V(\mathcal{H})=\emptyset$ for both $i=1,2$, then, since $x_{1}$ and $x_{2}$ are cut vertices satisfying Lemma 4.4, we must have $x_{1},x_{2}\in S$, a contradiction. Hence, $D_{i}\cap V(\mathcal{H})\neq\emptyset$ for both $i$. Let $x_{i}^{\prime}\in D_{i}\cap V(\mathcal{H})$ for $i=1,2$. Moreover, $x_{1}^{\prime}\notin f_{1}$ by the maximality condition on $x_{1}$. Since $x_{i}^{\prime}\in V(\mathcal{H})$ and $\ell\neq 1$, there exists $x_{i}^{\prime\prime}\in N_{G}(x_{i}^{\prime})$ such that $\operatorname{d}(x_{i}^{\prime\prime},a_{k})>\operatorname{d}(x_{i}^{\prime},a_{k})\text{ for }i=1,2.$ Also, $x_{2}^{\prime},x_{2}^{\prime\prime}\in D_{2}$, and since $D_{2}\cap f_{1}=\emptyset$, we have $x_{2}^{\prime},x_{2}^{\prime\prime}\notin f_{1}$. If $x_{1}^{\prime\prime}\in f_{1}$, then $x_{1}^{\prime}\in f_{1}$, a contradiction. Thus, $x_{1}^{\prime\prime}\notin f_{1}$. This implies that $f_{1}$ misses at least five vertices: $x_{1}^{\prime},x_{1}^{\prime\prime},x_{2},x_{2}^{\prime},x_{2}^{\prime\prime}\notin f_{1}$, contradicting the assumption $|f_{1}|=r+1\geq n-4$. Therefore, either $x_{1}\in S$ or $x_{2}\in S$. Without loss of generality, assume $x_{1}\in S$.
Let $D_{1}=C_{x_{1}}$. Define $D=f_{2}\setminus(C_{a_{k}}\cup\{a_{k}\}).$ By construction of $C_{a_{k}}$, we have $D\subseteq G_{1}$, where $G_{1}$ is the connected component of $G\setminus a_{k}$ containing $V(\mathcal{H})\setminus\{a_{k}\}$. Now, we prove that the set $S^{\prime}=D\cup\{a,y_{2},\ldots,y_{p},x_{1}\}\cup D_{1}$ is connected. Let $\alpha,\beta\in D$. Then there exists a path in $G_{1}$, hence in $G$, hence in $f_{2}$, and no internal vertex lies in $C_{a_{k}}\cup\{a_{k}\}$, so $D$ is connected. Let $c\in D_{1}$ and $d\in D$. Then there exists a path $P(d,a)$ in $f_{2}\setminus(C_{a_{k}}\cup\{a_{k}\})$, and a path $ay_{2}\cdots y_{p}x_{1}r_{1}\cdots r_{t}c$ where $r_{i}\in D_{1}$. Thus, $S^{\prime}$ is connected.
Observe that $f_{2}=D\cup C_{a_{k}}\cup\{a_{k}\}$, so $r+1=|D|+|C_{a_{k}}|+1.$ Also, $|S^{\prime}|\geq r+1$, so we can choose $f\in E(\mathcal{C}_{r}(G))$ such that $f\subseteq S^{\prime}$. Set $e=f\cap V(\mathcal{H})$. Clearly, $e\subseteq e_{1}\cup e_{2}\setminus\{a_{k}\}$. If $e\in E(\mathcal{H})$, we are done. Otherwise, there exists $e_{s}\in E(\mathcal{H})$ such that $e_{s}\subseteq e_{1}\cup e_{2}\setminus\{a_{k}\}$. This proves that $a_{k}$ is a simplicial vertex. ∎
We now turn to the case $r=2$, where a clean structural condition on $G$ guarantees the shellability of $\mathcal{I}_{2}(G)$.
Theorem 4.6.
If $G$ is a block graph, then $\mathcal{I}_{2}(G)$ is shellable.
Proof.
By Theorem 2.2, it is enough to prove that $\mathcal{C}_{2}(G)$ is $\operatorname{W}$-chordal. Block graphs admit a hierarchical structure. By Proposition 3.1, it suffices to show that every $c^{\prime}$-minor of $\mathcal{C}_{2}(G)$ contains a simplicial vertex, assuming $G$ is connected. Let $\mathcal{H}$ be a $c^{\prime}$-minor of $\mathcal{C}_{2}(G)$. Without loss of generality, assume $G$ is connected and adopt the notation from Set-up 3.6. If $1\leq n\leq 2$, then $G$ is a star-clique graph, and by Corollary 3.3, $\mathcal{H}$ has a simplicial vertex. Now assume $n\geq 3$. If $\ell=1$, then Lemma 3.2 implies that $\mathcal{H}$ has a simplicial vertex. So suppose $\ell\neq 1$. In particular, $x_{0}\notin V(\mathcal{H})$, where $x_{0}\in V(B_{1})\cap V_{1}$. By Lemma 3.7, $x_{1}\in V(\mathcal{H})$, and since $x_{2}\in N_{G}(x_{1})\setminus\{x_{0}\}$, we also have $x_{2}\in V(\mathcal{H})$.
We now show that $|B_{1}|=2$. If $|B_{1}|\geq 3$, then $\{x_{1}\}$ would be an edge of $\mathcal{H}$, contradicting the definition of a $c^{\prime}$-minor. So $|B_{1}|=2$. If $V(B_{2})=\{x_{1},x_{2}\}$, then $x_{1}$ appears in only one hyperedge of $\mathcal{H}$, hence is simplicial. Otherwise, suppose $V(B_{2})=\{x_{1},x_{2},\beta_{1},\dots,\beta_{p}\}$ for some $p\geq 1$. Then for each $i$, we have $\beta_{i}\in V_{2}\cap V(\mathcal{H})$ and $|N_{G}(\beta_{i})\cap V_{1}|=1$. Let $N_{G}(\beta_{i})\cap V_{1}=\{z_{i}\}$. Now consider $x_{1}\in e_{1}\cap e_{2}$, where $e_{1}\neq e_{2}$ and $e_{i}\in E(\mathcal{H})$ for $i=1,2$. Since $x_{1}\in V_{2}$, Observation 3.5 implies $S_{x_{1}}=\{x_{0}\}$. By Lemma 3.8, there exist $f_{1},f_{2}\in E(\mathcal{C}_{2}(G))$ such that $S_{x_{1}}\subseteq f_{i}$ and the set $f_{1}\cup f_{2}\setminus\{x_{0},x_{1}\}$ is connected. Each $f_{i}$ is either $\{x_{0},x_{1},x_{2}\}$ or $\{x_{0},x_{1},\beta_{j}\}$ for some $j$. Let $x_{1}^{\prime}\in e_{1}\setminus e_{2}$ and $x_{2}^{\prime}\in e_{2}\setminus e_{1}$. Without loss of generality, assume $x_{1}^{\prime}=\beta_{j}$ for some $j$. Then $|f_{1}\cup f_{2}\setminus\{x_{0},x_{1}\}|=2<r+1=3.$ Define $C=\{z_{j}\}$. Then $C$ satisfies the conditions in (3.1). Hence, by Lemma 3.10, $x_{1}$ is a simplicial vertex of $\mathcal{H}$. ∎
Beyond the case $r=2$, one can obtain shellability of $\mathcal{I}_{r}(G)$ for all $r\geq 2$ under an additional structural restriction on $G$, namely a bound on its diameter.
Theorem 4.7.
If $G$ is a block graph with $\operatorname{diam}(G)\leq 4$, then $\mathcal{I}_{r}(G)$ is shellable for all $r\geq 2$.
Proof.
By Theorem 2.2, it is enough to prove that $\mathcal{C}_{r}(G)$ is $\operatorname{W}$-chordal. Let $\mathcal{H}$ be a $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$ for some $r\geq 2$. We claim that $\mathcal{H}$ contains a simplicial vertex. To prove this, assume that $G$ is connected and adopt the notation from Set-up 3.6. We consider cases based on the number of blocks $n$.
Case 1: $1\leq n\leq 2$. In this case, $G$ is a star-clique graph. By Corollary 3.3, $\mathcal{H}$ contains a simplicial vertex.
Case 2: $3\leq n\leq 4$. If $r=2$, then by Theorem 4.6, $\mathcal{C}_{2}(G)$ is $\operatorname{W}$-chordal, so $\mathcal{H}$ contains a simplicial vertex.
Now assume $r\geq 3$. We analyze based on the value of $\ell$.
Subcase: $\ell=1$. Then, by Lemma 3.2, $\mathcal{H}$ has a simplicial vertex.
Subcase: $\ell=2$. Let $V_{2}\cap V(\mathcal{H})=\{v_{1},\dots,v_{s}\}$. By Observation 3.5, define $|S_{v_{k}}|=\min\left\{|S_{v_{i}}|:1\leq i\leq s\right\}.$ We claim that $v_{k}$ is simplicial in $\mathcal{H}$. Let $e_{1},e_{2}\in E(\mathcal{H})$ be distinct edges such that $v_{k}\in e_{1}\cap e_{2}$. By Lemma 3.8, there exist $f_{1},f_{2}\in E(\mathcal{C}_{r}(G))$ such that $S_{v_{k}}\subseteq f_{i}$, $f_{i}\cap V(\mathcal{H})=e_{i}$ for $i=1,2$, and the set $f_{1}\cup f_{2}\setminus(S_{v_{k}}\cup\{v_{k}\})$ is connected.
We analyze this for different values of $n$:
Case $n=3$. Here, $V_{2}=B_{2}\setminus V_{1}$. Let $x_{1}^{\prime}\in e_{1}\setminus e_{2}$ and $x_{2}^{\prime}\in e_{2}\setminus e_{1}$. Since $x_{1}^{\prime},x_{2}^{\prime}\in V_{2}$, there are two possibilities: If $\left|f_{1}\cup f_{2}\setminus(S_{v_{k}}\cup\{v_{k}\})\right|\geq r+1$, then by Lemma 3.10, $v_{k}$ is simplicial. If $\left|f_{1}\cup f_{2}\setminus(S_{v_{k}}\cup\{v_{k}\})\right|\leq r$, then $x_{1}^{\prime}=v_{j}$ for some $j$, and we let $C=S_{v_{j}}$. Then each component of $C$ satisfies the conditions in (3.1), and again by Lemma 3.10, $v_{k}$ is simplicial.
Case $n=4$. Here, $V_{2}=N_{G}(x_{2})\setminus V_{1}$. Let $x_{1}^{\prime}\in e_{1}\setminus e_{2}$ and $x_{2}^{\prime}\in e_{2}\setminus e_{1}$. If $x_{1}^{\prime},x_{2}^{\prime}\notin V_{2}$, then they belong to $V_{t}$ for $t\geq 3$, which is impossible since $V_{3}=\{x_{2}\}$ and $V_{t}=\emptyset$ for $t\geq 4$. So either $x_{1}^{\prime}$ or $x_{2}^{\prime}$ lies in $V_{2}$, and we can proceed as in the $n=3$ case to construct an edge $e_{3}\subseteq(e_{1}\cup e_{2})\setminus\{v_{k}\}$, confirming that $v_{k}$ is simplicial.
Note that $\ell=3$ does not occur. For $n=3$, we have $V_{3}=\emptyset$, and for $n=4$, $V_{3}=\{x_{2}\}$ contradicts the definition of a $c^{\prime}$-minor. Also, $V_{t}=\emptyset$ for all $t\geq 4$.
Thus, every $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$ contains a simplicial vertex. Since every induced subgraph of $G$ is a block graph with diameter at most $4$, Proposition 3.1 implies that $\mathcal{C}_{r}(G)$ is a $\operatorname{W}$-chordal hypergraph for all $r\geq 2$. ∎
Let $G$ be a tree, and let $P:=y_{1}y_{2}\cdots y_{n}$ be a longest path in $G$, called the central path, where $y_{1}$ and $y_{n}$ are leaves. The tree $G$ is a caterpillar if every vertex $z\notin V(P)$ is adjacent to an internal vertex of $P$, i.e., $\mathrm{d}(z,y_{i})=1$ for some $i\in\{2,\ldots,n-1\}$. More generally, $G$ is a lobster if every vertex $z\notin V(P)$ either satisfies $\mathrm{d}(z,y_{i})=1$ for some $i\in\{2,\ldots,n-1\}$ or $\mathrm{d}(z,y_{i})=2$ for some $i\in\{3,\ldots,n-2\}$.
The following result shows that, more generally, every forest of diameter at most five has a shellable $r$-independence complex for all $r\geq 2$.
Theorem 4.8.
If $G$ is a tree with $\operatorname{diam}(G)\leq 5$, then $\mathcal{I}_{r}(G)$ is shellable for all $r\geq 2$.
Proof.
By Theorem 2.2, it is enough to prove that $\mathcal{C}_{r}(G)$ is $\operatorname{W}$-chordal. If $\operatorname{diam}(G)\leq 4$, then by Theorem 4.7, $\mathcal{C}_{r}(G)$ is $\operatorname{W}$-chordal for all $r\geq 2$. Now assume $\operatorname{diam}(G)=5$. By Proposition 3.1, it suffices to show that every $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$ has a simplicial vertex when $G$ is connected. Let $\mathcal{H}$ be a $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$. Since $G$ is a tree with diameter $5$, there exists a maximum-length path $x_{0}x_{1}x_{2}x_{3}x_{4}x_{5}$. Thus, $G$ is a lobster tree. Assume the notation from Set-up 3.6. If $\ell=1$, then by Lemma 3.2, $\mathcal{H}$ contains a simplicial vertex. Consider next the case $\ell=2$. Note that $V_{2}=A_{1}\cup A_{2}$, where $A_{1}=N_{G}(x_{3})\setminus(V_{1}\cup\{x_{2}\})$, $A_{2}=N_{G}(x_{2})\setminus(V_{1}\cup\{x_{3}\})$. Let $V_{2}\cap V(\mathcal{H})=\{v_{1},\ldots,v_{s}\}$. Using the notation from Observation 3.5, define $|S_{v_{k}}|=\min\left\{|S_{v_{i}}|\mid i=1,\ldots,s\right\}.$ We claim that $v_{k}$ is a simplicial vertex of $\mathcal{H}$. To prove this, consider two edges $e_{1},e_{2}\in E(\mathcal{H})$ such that $v_{k}\in e_{1}\cap e_{2}$ and $e_{1}\neq e_{2}$. By Lemma 3.8, there exist edges $f_{1},f_{2}\in E(\mathcal{C}_{r}(G))$ such that $S_{v_{k}}\subseteq f_{i},\quad f_{i}\cap V(\mathcal{H})=e_{i}\quad\text{for }i=1,2,\text{and }f_{1}\cup f_{2}\setminus\left(S_{v_{k}}\cup\{v_{k}\}\right)\text{ is connected.}$ Let $x_{1}^{\prime}\in e_{1}\setminus e_{2}$ and $x_{2}^{\prime}\in e_{2}\setminus e_{1}$. Since $v_{k}\in V_{2}$, we have $v_{k}\in A_{1}$ or $v_{k}\in A_{2}$. Without loss of generality, assume $v_{k}\in A_{1}$.
We claim that either $x_{1}^{\prime}\in V_{2}$ or $x_{2}^{\prime}\in V_{2}$. Suppose not. Since $V_{t}=\emptyset$ for all $t\geq 4$ and $\ell=2$, it follows that $x_{1}^{\prime},x_{2}^{\prime}\in V_{3}=\{x_{2},x_{3}\}$. Without loss of generality, let $x_{1}^{\prime}=x_{2}$ and $x_{2}^{\prime}=x_{3}$. Then $v_{k}x_{3}(=x_{2}^{\prime})x_{2}(=x_{1}^{\prime})$ is a unique path from $v_{k}$ to $x_{1}^{\prime}$. Since both $v_{k},x_{1}^{\prime}\in f_{1}$, and $x_{3}=x_{2}^{\prime}\in f_{1}$, we contradict the assumption that $x_{2}^{\prime}\in e_{2}\setminus e_{1}$. Hence, either $x_{1}^{\prime}\in V_{2}$ or $x_{2}^{\prime}\in V_{2}$. Assume $x_{1}^{\prime}\in V_{2}$ and let $x_{1}^{\prime}=v_{j}$ for some $j$. If $\left|f_{1}\cup f_{2}\setminus\left(S_{v_{k}}\cup\{v_{k}\}\right)\right|\geq r+1,$ then by Lemma 3.10, $v_{k}$ is a simplicial vertex. Otherwise, suppose $\left|f_{1}\cup f_{2}\setminus\left(S_{v_{k}}\cup\{v_{k}\}\right)\right|\leq r.$ Define $C=S_{v_{j}}$. Then each component of $C$ satisfies the conditions of (3.1). Since $|C|\geq|S_{v_{k}}|$, Lemma 3.10 implies $v_{k}$ is simplicial in $\mathcal{H}$.
Now consider $\ell=3$. Let $v\in V_{3}\cap V(\mathcal{H})$. Since $V_{3}=\{x_{2},x_{3}\}$ and $G\setminus(V_{1}\cup V_{2})=V_{3}$, any such vertex $v$ lies in a unique edge $\{x_{2},x_{3}\}$. Hence, $v$ is simplicial. Finally, observe that $\ell\geq 4$ is not possible since $V_{t}=\emptyset$ for all $t\geq 4$. Thus, every $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$ contains a simplicial vertex, completing the proof. ∎
Let $G$ be the graph shown below, referred to as the $\operatorname{T}_{3}$-graph. The vertex set is $V(G)=\{x_{1},x_{2},x_{3},x_{4},x_{5},x_{6},x_{7},x_{8},x_{9}\}$, and the edge set is
$E(G)=\{\{x_{1},x_{2}\},\{x_{2},x_{3}\},\{x_{3},x_{4}\},\{x_{3},x_{5}\},\{x_{4},x_{5}\},\{x_{5},x_{6}\},\{x_{6},x_{7}\},\{x_{4},x_{8}\},\{x_{8},x_{9}\}\}.$
$x_{1}$$x_{2}$$x_{3}$$x_{4}$$x_{5}$$x_{6}$$x_{7}$$x_{8}$$x_{9}$
For $r=3$, the shellability of $\mathcal{I}_{3}(G)$ for block graphs can be characterized by the absence of certain forbidden configurations. In particular, excluding $\operatorname{T}_{3}$ as an induced subgraph guarantees shellability.
Theorem 4.9.
Let $G$ be a block graph. If $G$ is $\operatorname{T}_{3}$-free, then the complex $\mathcal{I}_{3}(G)$ is shellable.
Proof.
By Theorem 2.2, it is enough to prove that $\mathcal{C}_{3}(G)$ is $\operatorname{W}$-chordal. By Proposition 3.1, it suffices to show that every $c^{\prime}$-minor of $\mathcal{C}_{3}(G)$ has a simplicial vertex when $G$ is connected. Let $\mathcal{H}$ be a $c^{\prime}$-minor of $\mathcal{C}_{3}(G)$, following the notation from Set-up 3.6. If $1\leq n\leq 4$, then by Theorem 4.7, $\mathcal{H}$ has a simplicial vertex. Now assume $n\geq 5$. Suppose there exists a vertex $y\in V(B_{3})\setminus\{x_{2},x_{3}\}$ such that $d(z,y)=2$, i.e., there is a path yazyaz for some $z\in V(G)$. Clearly $a\in V(B_{3})$, if $a\notin V(B_{3})$, then $G$ contains a $\operatorname{T}_{3}$, contradicting the assumption that $G$ is $\operatorname{T}_{3}$-free. If $\ell=1$, then by Lemma 3.2, $\mathcal{H}$ contains a simplicial vertex. We now consider the case $\ell\neq 1$.
Suppose $(N_{G}(x_{2})\setminus N_{G}(x_{3}))\cap V_{2}\cap V(\mathcal{H})\neq\emptyset.$ This forces $\ell=2$. Let $(N_{G}(x_{2})\setminus N_{G}(x_{3}))\cap V_{2}\cap V(\mathcal{H})=\{v_{1},\ldots,v_{q}\}.$ Using the notation from Observation 3.5, define $|S_{v_{k}}|=\min\left\{|S_{v_{i}}|\mid i\in\{1,\ldots,q\}\right\}.$ We claim that $v_{k}$ is a simplicial vertex of $\mathcal{H}$. To see this, let $e_{1},e_{2}\in E(\mathcal{H})$ be such that $v_{k}\in e_{1}\cap e_{2}$ and $e_{1}\neq e_{2}$. By Lemma 3.8, there exist edges $f_{1},f_{2}\in E(\mathcal{C}_{3}(G))$ such that $S_{v_{k}}\subseteq f_{i}$ and $f_{i}\cap V(\mathcal{H})=e_{i}$ for $i=1,2$. Moreover, the set $f_{1}\cup f_{2}\setminus\left(S_{v_{k}}\cup\{v_{k}\}\right)$ is connected. Let $x_{1}^{\prime}\in e_{1}\setminus e_{2}$ and $x_{2}^{\prime}\in e_{2}\setminus e_{1}$. Note that $|S_{v_{k}}|\leq 2$; otherwise, $\{v_{k}\}\in E(\mathcal{H})$, contradicting the assumption that $\mathcal{H}$ is a $c^{\prime}$-minor.
Case 1: $|S_{v_{k}}|=1$. Let $S_{v_{k}}=\{z_{0}\}$. We first claim that either $x_{1}^{\prime}$ or $x_{2}^{\prime}$ must belong to $V_{2}$. Suppose, for contradiction, that neither $x_{1}^{\prime}$ nor $x_{2}^{\prime}$ is in $V_{2}$. Since $|f_{i}|=4$ and $S_{v_{k}}\subseteq f_{i}$, it follows that $x_{1}^{\prime},x_{2}^{\prime}\in\{x_{2},x_{3}\}$. As $x_{1}^{\prime}\neq x_{2}^{\prime}$, we may assume without loss of generality that $x_{1}^{\prime}=x_{2}$ and $x_{2}^{\prime}=x_{3}$. Since $v_{k},x_{3}\in f_{2}$, there exists a shortest path $P(v_{k},x_{3})$ in $f_{2}$, with all internal vertices in $f_{2}$. Given $z_{0}\in f_{2}$ and $|f_{2}|=4$, and using the fact that $v_{k}\in(N_{G}(x_{2})\setminus N_{G}(x_{3}))\cap V_{2}\cap V(\mathcal{H})$, we conclude that $P(v_{k},x_{3}):=v_{k}x_{2}x_{3}.$ This is a contradiction since $x_{2}=x_{1}^{\prime}\in f_{2}$ but $x_{1}^{\prime}\in e_{1}\setminus e_{2}$. Therefore, either $x_{1}^{\prime}\in V_{2}$ or $x_{2}^{\prime}\in V_{2}$. Without loss of generality, assume $x_{1}^{\prime}\in V_{2}$, and let $C=S_{x_{1}^{\prime}}$. Then each component of $C$ satisfies the conditions in (3.1). Since $|C|\geq|S_{v_{k}}|$, Lemma 3.10 implies that $v_{k}$ is simplicial.
Case 2: $|S_{v_{k}}|=2$. We claim that at least one of $x_{1}^{\prime}$ or $x_{2}^{\prime}$ lies in $\{v_{1},\ldots,v_{q}\}$. Suppose not. Then $x_{1}^{\prime},x_{2}^{\prime}\notin\{v_{1},\ldots,v_{q}\}$, which would force $x_{1}^{\prime}=x_{2}^{\prime}=x_{2}$, a contradiction. Hence, without loss of generality, assume $x_{1}^{\prime}=v_{j}$ for some $j$, and let $C=S_{v_{j}}$. Then each component of $C$ satisfies conditions (i)–(iv) in (3.1). Furthermore, $|f_{1}\cup f_{2}\setminus(S_{v_{k}}\cup\{v_{k}\})|=2<3.$ So by Lemma 3.10, $v_{k}$ is simplicial.
Suppose $(N_{G}(x_{2})\setminus N_{G}(x_{3}))\cap V_{2}\cap V(\mathcal{H})=\emptyset.$
Assume $(N_{G}(x_{3})\setminus N_{G}(x_{4}))\cap V_{2}\cap V(\mathcal{H})\neq\emptyset.$ This implies $\ell=2$. Let $(N_{G}(x_{3})\setminus N_{G}(x_{4}))\cap V_{2}\cap V(\mathcal{H})=\{v_{1}^{\prime},\ldots,v_{p}^{\prime}\}.$ Using Observation 3.5, choose $v_{t}^{\prime}$ such that $|S_{v_{t}^{\prime}}|=\min\{|S_{v_{j}^{\prime}}|\mid 1\leq j\leq p\}.$ Take two distinct edges $e_{1},e_{2}\in E(\mathcal{H})$ with $v_{t}^{\prime}\in e_{1}\cap e_{2}$. By Lemma 3.8, there exist edges $f_{1},f_{2}\in E(\mathcal{C}_{3}(G))$ such that $S_{v_{t}^{\prime}}\subseteq f_{i}$ for each $i$, $f_{i}\cap V(\mathcal{H})=e_{i}$, and $f_{1}\cup f_{2}\setminus(S_{v_{t}^{\prime}}\cup\{v_{t}^{\prime}\})$ is connected. Let $x_{1}^{\prime}\in e_{1}\setminus e_{2}$ and $x_{2}^{\prime}\in e_{2}\setminus e_{1}$. Note $|S_{v_{t}^{\prime}}|\leq 2$; otherwise $\{v_{t}^{\prime}\}\in E(\mathcal{H})$, contradicting $\mathcal{H}$ being a $c^{\prime}$-minor. Suppose there exists another maximal clique path $\operatorname{CP}(A_{3},A_{2},A_{1},B_{4},\ldots,B_{n})$ where each $A_{i}$ is a clique ($1\leq i\leq 3$) with connecting vertices $b,a,x_{3},\ldots,x_{n-1}$. If $(N_{G}(a)\setminus N_{G}(x_{3}))\cap V_{2}\cap V(\mathcal{H})\neq\emptyset$, then the conditions of previous case are satisfied, and consequently there exists a simplicial vertex in $\mathcal{H}$. Otherwise ($(N_{G}(a)\setminus N_{G}(x_{3}))\cap V_{2}\cap V(\mathcal{H})=\emptyset$), we have $b\notin V(\mathcal{H})$.
Case 1: $|S_{v_{t}^{\prime}}|=1$. Either $x_{1}^{\prime}$ or $x_{2}^{\prime}$ belongs to $V_{2}\cup\{x_{2},a\}$. Without loss of generality, assume $x_{1}^{\prime}\in V_{2}\cup\{x_{2},a\}$. Define
$C=\begin{cases}S_{x_{1}^{\prime}}&\text{if }x_{1}^{\prime}\in V_{2},\\ \{x_{1}\}&\text{if }x_{1}^{\prime}=x_{2},\\ \{b\}&\text{if }x_{1}^{\prime}=a.\end{cases}$
Then each component of $C$ satisfies (3.1) and $|C|\geq|S_{v_{t}^{\prime}}|$, so by Lemma 3.10, $v_{t}^{\prime}$ is simplicial.
Case 2: $|S_{v_{t}^{\prime}}|=2$. Similarly, either $x_{1}^{\prime}$ or $x_{2}^{\prime}$ belongs to $\{v_{1}^{\prime},\ldots,v_{p}^{\prime}\}\cup\{x_{2}\}$. Let $x_{0}\in V(B_{1})\cap V_{1}$ and note that $|f_{1}\cup f_{2}\setminus(S_{v_{t}^{\prime}}\cup\{v_{t}^{\prime}\})|=2<4.$ Assuming $x_{1}^{\prime}\in\{v_{1}^{\prime},\ldots,v_{p}^{\prime}\}\cup\{x_{2}\}$, define
$C=\begin{cases}S_{x_{j}^{\prime}}&\text{if }x_{1}^{\prime}=v_{j}^{\prime},\\ \{x_{0},x_{1}\}&\text{if }x_{1}^{\prime}=x_{2}.\end{cases}$
Again each component of $C$ satisfies (3.1) and $|C|\geq|S_{v_{t}^{\prime}}|$, proving $v_{t}^{\prime}$ is simplicial via Lemma 3.10. If no such maximal clique path $\operatorname{CP}(A_{3},A_{2},A_{1},B_{4},\ldots,B_{n})$ exists, analogous arguments show $v_{t}^{\prime}$ remains simplicial in $\mathcal{H}$
Suppose $(N_{G}(x_{3})\setminus N_{G}(x_{4}))\cap V_{2}\cap V(\mathcal{H})=\emptyset.$ Observe that $x_{0},x_{1}\notin V(\mathcal{H})$, where $x_{0}\in V(B_{1})\cap V_{1}$. By Lemma 3.7, it follows that $x_{2}\in V(\mathcal{H})$. Since $x_{3}\in N_{G}(x_{2})\setminus\{x_{1}\}$, Lemma 3.7 further implies that $x_{3}\in V(\mathcal{H})$. Given that $(N_{G}(x_{2})\setminus N_{G}(x_{3}))\cap V_{2}\cap V(\mathcal{H})=\emptyset,$ we conclude that the only edge containing $x_{2}$ is $\{x_{2},x_{3}\}$, since $\{x_{0},x_{1},x_{2},x_{3}\}\cap V(\mathcal{H})=\{x_{2},x_{3}\}.$ Thus, $x_{2}$ is a simplicial vertex of $\mathcal{H}$. ∎
As an immediate consequence of Theorem 4.9, we obtain the following result:
Corollary 4.10.
If $G$ is a forest, then the complex $\mathcal{I}_{3}(G)$ is shellable.
We now introduce a special class of block graphs defined in terms of their clique-path structure.
Definition 4.11.
A connected graph $H$ is called a $\mathcal{T}_{1}$-graph if it is a block graph and contains a clique-path subgraph $P=\operatorname{CP}(B_{1},B_{2},\dots,B_{n})$, $n\geq 1$, of maximum length, with connecting vertices $x_{i}\in B_{i}\cap B_{i+1}$ for $1\leq i\leq n-1$, such that every clique $K$ of $H$ not belonging to $P$ intersects $P$ in exactly one of these connecting vertices, that is, $V(K)\cap V(P)=\{x_{i}\}$ for some $i$ with $1\leq i\leq n-1$. A graph $G$ is called a $\mathcal{T}_{1}$-graph if each of its connected components is a $\mathcal{T}_{1}$-graph.
We next extend this notion by allowing a broader attachment of cliques beyond the clique-path, leading to the following definition.
Definition 4.12.
A connected graph $H$ is called a $\mathcal{T}_{2}$-graph if it is a block graph and contains a clique-path subgraph $P=\operatorname{CP}(B_{1},B_{2},\dots,B_{n})$, $n\geq 1$, of maximum length, with connecting vertices $x_{i}\in B_{i}\cap B_{i+1}$ for $1\leq i\leq n-1$, such that every clique $K$ of $H$ not belonging to $P$ satisfies one of the following conditions:
  1. (1)
    $V(K)\cap V(P)=\{x_{i}\}$ for some $i$ with $1\leq i\leq n-1$, or
  2. (2)
    $V(K)\cap V(P)=\emptyset$ and there exists a clique $K^{\prime}$ of $H$ satisfying (a) such that $V(K)\cap V(K^{\prime})=\{y\}$ for some vertex $y$.
A graph $G$ is called a $\mathcal{T}_{2}$-graph if each of its connected components is a $\mathcal{T}_{2}$-graph.
The following observation illustrates that both $\mathcal{T}_{1}$-graphs and $\mathcal{T}_{2}$-graphs form a hierarchy.
Observation 4.13.
Let $G$ be a connected $\mathcal{T}_{1}$-graph, and let $\operatorname{CP}(B_{1},\dots,B_{n})$, $n\geq 1$, be a maximum-size clique path subgraph of $G$. Since $G$ is a block graph, every vertex is either simplicial or a cut vertex. If $x$ is simplicial, then $G\setminus x$ remains connected. When $x\in B_{1}$ or $x\in B_{n}$, if $\deg(x)=1$ we may choose another maximum-size clique path, while if $\deg(x)>1$ the same clique path $\operatorname{CP}(B_{1},\dots,B_{n})$ remains valid; if $x\notin B_{1}\cup B_{n}$, the path $\operatorname{CP}(B_{1},\dots,B_{n})$ still satisfies all conditions of Definition 4.11. If $x$ is a cut vertex, then $G\setminus x$ decomposes into connected components, each containing a maximum-size clique path that satisfies Definition 4.11. Thus the structure of a $\mathcal{T}_{1}$-graph is preserved, and hence $\mathcal{T}_{1}$-graphs form a hierarchy. A similar argument shows that $\mathcal{T}_{2}$-graphs also form a hierarchy.
The structural restrictions imposed by $\mathcal{T}_{2}$-graphs allow us to extend shellability results to higher independence complexes. In particular, we have the following.
Theorem 4.14.
If $G$ is a $\mathcal{T}_{2}$-graph, then $\mathcal{I}_{4}(G)$ is shellable.
Proof.
By Theorem 2.2, Proposition 3.1 and Observation 4.13, it suffices to show that every $c^{\prime}$-minor of $\mathcal{C}_{4}(G)$ has a simplicial vertex, assuming $G$ is connected. Let $\mathcal{H}$ be a $c^{\prime}$-minor of $\mathcal{C}_{4}(G)$, and let us follow the notation introduced in Set-up 3.6. If $1\leq n\leq 4$, then by Theorem 4.7, the hypergraph $\mathcal{H}$ contains a simplicial vertex. Now suppose $n\geq 5$. If $\ell=1$, then Lemma 3.2 guarantees that $\mathcal{H}$ has a simplicial vertex. It remains to consider the case when $\ell\neq 1$.
Case 1: Suppose $N_{G}(x_{2})\cap V_{2}\cap V(\mathcal{H})\neq\emptyset$. This forces $\ell=2$. Let $N_{G}(x_{2})\cap V_{2}\cap V(\mathcal{H})=\{v_{1},\ldots,v_{q}\}.$ Using the notation from Observation 3.5, define $v_{k}$ such that $|S_{v_{k}}|=\min\left\{|S_{v_{i}}|\mid i\in\{1,\ldots,q\}\right\}.$ We claim that $v_{k}$ is a simplicial vertex of $\mathcal{H}$. To establish this, consider any two edges $e_{1},e_{2}\in E(\mathcal{H})$ with $v_{k}\in e_{1}\cap e_{2}$ and $e_{1}\neq e_{2}$. By Lemma 3.8, there exist edges $f_{1},f_{2}\in E(\mathcal{C}_{4}(G))$ such that $S_{v_{k}}\subseteq f_{i}$ and $f_{i}\cap V(\mathcal{H})=e_{i}$ for $i=1,2$. Moreover, the set $f_{1}\cup f_{2}\setminus\left(S_{v_{k}}\cup\{v_{k}\}\right)$ is connected. Let $x_{1}^{\prime}\in e_{1}\setminus e_{2}$ and $x_{2}^{\prime}\in e_{2}\setminus e_{1}$. Note that $|S_{v_{k}}|\leq 3$; otherwise, if $|S_{v_{k}}|\geq 4$, then $\{v_{k}\}\in E(\mathcal{H})$, contradicting the assumption that $\mathcal{H}$ is a $c^{\prime}$-minor. Now, following the analysis in Case 1 of Theorem 4.9, we observe: If $|S_{v_{k}}|=1$, then either $x_{1}^{\prime}$ or $x_{2}^{\prime}$ must belong to $V_{2}$. If $2\leq|S_{v_{k}}|\leq 3$, then either $x_{1}^{\prime}$ or $x_{2}^{\prime}$ must lie in $\{v_{1},\ldots,v_{q}\}$. Without loss of generality, assume:
$x_{1}^{\prime}\in\begin{cases}V_{2},&\text{if }|S_{v_{k}}|=1,\\ \{v_{1},\ldots,v_{q}\},&\text{if }2\leq|S_{v_{k}}|\leq 3.\end{cases}$
In the second case, let $x_{1}^{\prime}=v_{j}$ for some $j\in\{1,\ldots,q\}$. Define
$C=\begin{cases}S_{x_{1}^{\prime}},&\text{if }x_{1}^{\prime}\in V_{2},\\ S_{v_{j}},&\text{if }x_{1}^{\prime}=v_{j}.\end{cases}$
Then each component of $C$ satisfies all the conditions listed in (3.1). Since $|C|\geq|S_{v_{k}}|$, Lemma 3.10 implies that $v_{k}$ is a simplicial vertex of $\mathcal{H}$.
Case 2: Suppose $N_{G}(x_{2})\cap V_{2}\cap V(\mathcal{H})=\emptyset$.
Subcase 1: Suppose $N_{G}(x_{3})\cap V_{2}\cap V(\mathcal{H})\neq\emptyset$, which forces $\ell=2$. Let $N_{G}(x_{3})\cap V_{2}\cap V(\mathcal{H})=\{v_{1}^{\prime},\ldots,v_{s}^{\prime}\}.$ Using the notation from Observation 3.5, define $|S_{v_{k}^{\prime}}|=\min\left\{|S_{v_{i}^{\prime}}|\mid i\in\{1,\ldots,s\}\right\}.$ We claim that $v_{k}^{\prime}$ is a simplicial vertex of $\mathcal{H}$. To prove this, consider two distinct edges $e_{1},e_{2}\in E(\mathcal{H})$ with $v_{k}^{\prime}\in e_{1}\cap e_{2}$. By Lemma 3.8, there exist edges $f_{1},f_{2}\in E(\mathcal{C}_{4}(G))$ such that $S_{v_{k}^{\prime}}\subseteq f_{i}$ and $f_{i}\cap V(\mathcal{H})=e_{i}$ for $i=1,2$. Moreover, the set $f_{1}\cup f_{2}\setminus\left(S_{v_{k}^{\prime}}\cup\{v_{k}^{\prime}\}\right)$ is connected. Let $x_{1}^{\prime}\in e_{1}\setminus e_{2}$ and $x_{2}^{\prime}\in e_{2}\setminus e_{1}$.
We analyze the three possibilities for $|S_{v_{k}^{\prime}}|$:
  1. (1)
    If $|S_{v_{k}^{\prime}}|=1$: As in the proof of Theorem 4.9, one of $x_{1}^{\prime}$ or $x_{2}^{\prime}$ must belong to $V_{2}\cup\{x_{2}\}$. Without loss of generality, assume $x_{1}^{\prime}\in V_{2}\cup\{x_{2}\}$. Define
    $C=\begin{cases}S_{x_{1}^{\prime}}&\text{if }x_{1}^{\prime}\in V_{2},\\ \{x_{1}\}&\text{if }x_{1}^{\prime}=x_{2}.\end{cases}$
  2. (2)
    If $|S_{v_{k}^{\prime}}|=2$: A similar argument as above shows that $x_{1}^{\prime}\in\{v_{1}^{\prime},\ldots,v_{s}^{\prime}\}\cup\{x_{2}\}$. Define
    $C=\begin{cases}S_{v_{j}^{\prime}}&\text{if }x_{1}^{\prime}=v_{j}^{\prime},\\ \{x_{0},x_{1}\}&\text{if }x_{1}^{\prime}=x_{2},\end{cases}$
    where $x_{0}\in V(B_{1})\cap V_{1}$.
  3. (3)
    If $|S_{v_{k}^{\prime}}|=3$: Then either $x_{1}^{\prime}$ or $x_{2}^{\prime}$ belongs to $\{v_{1}^{\prime},\ldots,v_{s}^{\prime}\}$. Without loss of generality, assume $x_{1}^{\prime}=v_{j}^{\prime}$ and set $C=S_{v_{j}^{\prime}}$.
In all cases, each component of $C$ satisfies conditions (i)–(iv) in (3.1). Hence, by Lemma 3.10, $v_{k}^{\prime}$ is a simplicial vertex of $\mathcal{H}$.
Subcase 2: Suppose $N_{G}(x_{3})\cap V_{2}\cap V(\mathcal{H})=\emptyset$. First, assume $x_{2}\notin V(\mathcal{H})$. Then, by Lemma 3.7, it follows that $x_{3},x_{4}\in V(\mathcal{H})$. Since $N_{G}(x_{2})\cap V_{2}\cap V(\mathcal{H})=\emptyset$ and $N_{G}(x_{3})\cap V_{2}\cap V(\mathcal{H})=\emptyset,$ the only edge in $\mathcal{H}$ containing $x_{3}$ is $\{x_{3},x_{4}\}$, implying that $x_{3}$ is a simplicial vertex in $\mathcal{H}$.
Now, assume $x_{2}\in V(\mathcal{H})$. Suppose $\operatorname{diam}(G)>5$. Since $\mathcal{H}$ is a $c^{\prime}$-minor and hence cannot contain isolated edges, at least one of $x_{3}$ or $x_{4}$ must be in $V(\mathcal{H})$; otherwise, $\{x_{2}\}$ would form an isolated edge, contradicting the definition of a $c^{\prime}$-minor. Define $f=\{x_{0},x_{1},x_{2},x_{3},x_{4}\}$, where $x_{0}\in V(B_{1})\cap V_{1}$. The intersection $f\cap V(\mathcal{H})$ is:
$f\cap V(\mathcal{H})=\begin{cases}\{x_{2},x_{3}\}&\text{if }x_{3}\in V(\mathcal{H}),\ x_{4}\notin V(\mathcal{H}),\\ \{x_{2},x_{4}\}&\text{if }x_{3}\notin V(\mathcal{H}),\ x_{4}\in V(\mathcal{H}),\\ \{x_{2},x_{3},x_{4}\}&\text{if }x_{3},x_{4}\in V(\mathcal{H}).\end{cases}$
Given that $N_{G}(x_{2})\cap V_{2}\cap V(\mathcal{H})=\emptyset\quad\text{and}\quad N_{G}(x_{3})\cap V_{2}\cap V(\mathcal{H})=\emptyset,$ no additional vertices from $V_{2}$ are adjacent to $x_{2}$ or $x_{3}$ in $\mathcal{H}$. Thus, the only edge in $\mathcal{H}$ containing $x_{2}$ is $f\cap V(\mathcal{H})$, making $x_{2}$ a simplicial vertex.
Now suppose $\operatorname{diam}(G)=5$. Then $x_{4}\notin V(\mathcal{H})$. Moreover, $x_{3}\in V(\mathcal{H})$; otherwise, $\{x_{2}\}\in E(\mathcal{H})$, contradicting the $c^{\prime}$-minor property. Hence, $\{x_{2},x_{3}\}\in E(\mathcal{H})$. Since $N_{G}(x_{i})\cap V_{2}\cap V(\mathcal{H})=\emptyset$ for $i=2,3$, it follows that $x_{2}$ is a simplicial vertex of $\mathcal{H}$.
Thus, in all cases, $\mathcal{H}$ contains a simplicial vertex, completing the proof. ∎
For $\mathcal{T}_{1}$-graphs, the structural conditions are strong enough to guarantee shellability uniformly once the independence parameter is sufficiently large. This yields the following general result.
Theorem 4.15.
If $G$ is a $\mathcal{T}_{1}$-graph, then $\mathcal{I}_{r}(G)$ is shellable for all $r\geq 5$.
Proof.
By Theorem 2.2, Proposition 3.1 and Observation 4.13, it suffices to show that every $c^{\prime}$-minor of $\mathcal{C}_{r}(G)$ contains a simplicial vertex whenever $G$ is connected and $r\geq 5$. Let $\mathcal{H}$ be such a $c^{\prime}$-minor, following the notation from Set-up 3.6. If $1\leq n\leq 4$, then Theorem 4.7 guarantees that $\mathcal{H}$ has a simplicial vertex. Now consider $n\geq 5$. If $\ell=1$, then Lemma 3.2 applies. Suppose $\ell=2$. Let $V_{2}\cap V(\mathcal{H})=\{v_{1},\ldots,v_{s}\}$. Define $|S_{v_{k}}|=\min\left\{|S_{v_{i}}|\mid i\in\{1,\ldots,s\}\right\}.$ We claim that $v_{k}$ is a simplicial vertex of $\mathcal{H}$. To show this, take any two distinct edges $e_{1},e_{2}\in E(\mathcal{H})$ with $v_{k}\in e_{1}\cap e_{2}$. By Lemma 3.8, there exist edges $f_{1},f_{2}\in E(\mathcal{C}_{r}(G))$ such that $S_{v_{k}}\subseteq f_{i}$, $f_{i}\cap V(\mathcal{H})=e_{i}$ for $i=1,2,$ and $f_{1}\cup f_{2}\setminus(S_{v_{k}}\cup\{v_{k}\})$ is connected. Let $x_{1}^{\prime}\in e_{1}\setminus e_{2}$ and $x_{2}^{\prime}\in e_{2}\setminus e_{1}$. We first claim that at least one of $x_{1}^{\prime},x_{2}^{\prime}$ lies in $V_{2}$. Suppose not; then both lie in $\{x_{2},\ldots,x_{n-2}\}$. Since $V_{2}=A_{1}\cup A_{2}$, where $A_{1}=N_{G}(x_{2})\setminus(V_{1}\cup\{x_{3}\})$, $A_{2}=N_{G}(x_{n-2})\setminus(V_{1}\cup\{x_{n-3}\})$, assume $v_{k}\in A_{1}$, $x_{1}^{\prime}=x_{p}$, $x_{2}^{\prime}=x_{q}$ with $p<q$. Since $v_{k},x_{2}^{\prime}\in f_{2}$, there is a shortest path $P(v_{k},x_{2}^{\prime})$ in $f_{2}$ with all internal vertices in $f_{2}$. Then $P(v_{k},x_{2}^{\prime})=v_{k}x_{2}\cdots x_{q},\quad\text{so }x_{i}\in f_{2}\text{ for }i=2,\ldots,q.$ Since $x_{p}\in\{x_{2},\ldots,x_{q-1}\}$, we get $x_{p}=x_{1}^{\prime}\in e_{2}$, contradicting $x_{1}^{\prime}\in e_{1}\setminus e_{2}$. Thus, at least one of $x_{1}^{\prime},x_{2}^{\prime}$ lies in $V_{2}$. Assume $x_{1}^{\prime}=v_{j}\in V_{2}$, and define $C=S_{v_{j}}$. Then each component of $C$ satisfies the conditions in (3.1). Since $|C|\geq|S_{v_{j}}|$, Lemma 3.10 ensures that $v_{j}$ is simplicial.
Now consider $\ell=3$. Then $V_{3}=\{x_{2},x_{n-2}\}$. Suppose $x_{2}\in V_{3}\cap V(\mathcal{H})$. If $x_{2}$ lies in two edges $e_{1}\neq e_{2}$ of $\mathcal{H}$, pick $x_{1}^{\prime}\in e_{1}\setminus e_{2}$, $x_{2}^{\prime}\in e_{2}\setminus e_{1}$ with $x_{1}^{\prime}=x_{p}$, $x_{2}^{\prime}=x_{q}$ and $p<q$. Then $P(x_{2},x_{q})=x_{2}x_{3}\cdots x_{q}$, so $x_{p}\in\{x_{3},\ldots,x_{q-1}\}\subseteq f_{2}$, contradicting the assumption. Thus, only one edge contains $x_{2}$. A similar argument applies for $\ell\geq 4$, completing the proof. ∎
5. Shellability via Graph Constructions
This section establishes the shellability of $r$-independence complexes for graphs built from several key operations. We first prove that attaching star-cliques to a vertex cover of any graph $H$ yields a shellable complex (Theorem 5.1), a result we then extend to a broader class of chordal graphs (Theorem 5.2). Furthermore, we demonstrate that shellability is preserved under a clique-based partitioning of a graph (Theorem 5.4). Finally, we introduce the concept of a clique cycle and prove that appending star-cliques to its structure also results in a shellable complex (Theorem 5.5). These constructive theorems significantly expand the known families of graphs with shellable higher independence complexes.
Let $H$ be a graph. A subset $S\subseteq V(H)$ is called a vertex cover of $H$ if every edge of $H$ has at least one endpoint in $S$. Fix an integer $r\geq 1$. The graph $\operatorname{CCG}(H,S,r)$ is defined as the graph obtained from $H$ by attaching a star-clique graph $\operatorname{SC}(x)$ to each vertex $x\in S$, where $|V(\operatorname{SC}(x))|\geq r+1$.
Theorem 5.1.
Let $G=\operatorname{CCG}(H,S,r)$, where $r\geq 1$. Then $\mathcal{I}_{r}(G)$ is shellable.
Proof.
By Theorem 2.2(1), it suffices to show that every $c$-minor of $\mathcal{C}_{r}(G)$ contains a simplicial vertex. Let $\mathcal{H}$ be a $c$-minor of $\mathcal{C}_{r}(G)$. We consider two cases based on the intersection of $V(\mathcal{H})$ with the set of simplicial vertices of $G$. Let $V_{1}$ denote the set of all simplicial vertices of $G$. Suppose first that $V_{1}\cap V(\mathcal{H})\neq\emptyset$. If there exists a vertex $v\in V_{1}\cap V(\mathcal{H})$ such that $\{v\}\notin E(\mathcal{H})$, then by Lemma 3.2, $v$ is a simplicial vertex of $\mathcal{H}$, and we are done. Assume now that $\{v\}\in E(\mathcal{H})$ for every $v\in V_{1}\cap V(\mathcal{H})$. We claim that if $a\in S\cap V(\mathcal{H})$, then $\{a\}\in E(\mathcal{H})$. Consider such a vertex $a$. If $(\operatorname{SC}(a)\setminus\{a\})\cap V(\mathcal{H})\neq\emptyset$, then choose $c\in(\operatorname{SC}(a)\setminus\{a\})\cap V(\mathcal{H})$. Since $\{c\}\in E(\mathcal{H})$ by assumption, we can construct a connected set $\{c_{1},\ldots,c_{r},c\}$ such that $c_{i}\notin V(\mathcal{H})$ and $c_{i}\neq a$ for all $i$. Then $\{c_{1},\ldots,c_{r},a\}$ is connected, implying $\{a\}\in E(\mathcal{H})$. If $(\operatorname{SC}(a)\setminus\{a\})\cap V(\mathcal{H})=\emptyset$, then since $|\operatorname{SC}(a)|\geq r+1$, we can form a connected set of size $r+1$ containing $a$ and $r$ elements from $\operatorname{SC}(a)\setminus\{a\}$. Hence, $\{a\}\in E(\mathcal{H})$.
We now claim that every edge $e\in E(\mathcal{H})$ satisfies $|e|=1$. Suppose for contradiction that some $e\in E(\mathcal{H})$ has $|e|\geq 2$. Let $a\in e$, and choose $b\in e\setminus\{a\}$ such that $\operatorname{d}(a,b)\leq\operatorname{d}(a,x)$ for all $x\in e\setminus\{a\}$. If $a$ and $b$ are adjacent in $H$, then since $S$ is a vertex cover, either $a\in S$ or $b\in S$. Without loss of generality, assume $a\in S$. Then $\{a\}\in E(\mathcal{H})$ by the earlier claim, contradicting $|e|\geq 2$. If $a$ and $b$ are not adjacent in $H$, then $e=f\cap V(\mathcal{H})$ for some $f\in E(\mathcal{C}_{r}(G))$, where $a,b\in f$. There exists a shortest path $ay_{1}\cdots y_{k}b$ with $k\geq 1$, $y_{i}\in f$ for all $i$, and $y_{i}\notin V(\mathcal{H})$. Since $a\notin S$ (otherwise $\{a\}\in E(\mathcal{H})$), and $y_{1}$ is adjacent to $a$, the vertex cover property implies $y_{1}\in S$. Suppose $\{c^{\prime}\}\in E(\mathcal{H})$ for some $c^{\prime}\in\operatorname{SC}(y_{1})\setminus\{y_{1}\}$. Then there exists a connected set $\{c_{1}^{\prime},\ldots,c_{r}^{\prime},c^{\prime}\}$ with $c_{i}^{\prime}\notin V(\mathcal{H})$ for all $i$. The set $\{c_{1}^{\prime},\ldots,c_{r}^{\prime},y_{1},a\}$ is connected with size at least $r+2$, implying $\{a\}\in E(\mathcal{H})$, a contradiction. Therefore, $(\operatorname{SC}(y_{1})\setminus\{y_{1}\})\cap V(\mathcal{H})=\emptyset$. But then $\operatorname{SC}(y_{1})\cup\{a\}$ is a connected set of size at least $r+1$, again implying $\{a\}\in E(\mathcal{H})$-another contradiction. Thus, no such edge $e$ with $|e|\geq 2$ exists, and all edges in $\mathcal{H}$ are singletons. If $V_{1}\cap V(\mathcal{H})=\emptyset$, then again every edge of $\mathcal{H}$ is of size 1. In all cases, $\mathcal{H}$ contains a simplicial vertex. This completes the proof. ∎
The following result extends Theorem 5.1 by demonstrating the shellability of $\mathcal{I}_{r}(G)$ for a broader class of chordal graphs. Specifically, for a chordal graph $H$ and $G=\operatorname{CCG}(H,V(H),t)$, we show that $\mathcal{I}_{r}(G)$ is shellable for all $r\leq 2t+1$, thus relaxing the earlier bound on $r$.
Theorem 5.2.
Let $H$ be a chordal graph and let $G=\operatorname{CCG}(H,V(H),t)$ for some $t\geq 1$. If $r\leq 2t+1$, then $\mathcal{I}_{r}(G)$ is shellable.
Proof.
By Theorem 2.2(1), it suffices to show that every $c$-minor of $\mathcal{C}_{r}(G)$ has a simplicial vertex. Let $\mathcal{H}$ be a $c$-minor of $\mathcal{C}_{r}(G)$, i.e., $\mathcal{H}=\mathcal{C}_{r}(G)/V_{c}$ for some $V_{c}\subseteq V(G)$. Since $G$ is chordal, we adopt the notation from Observation 3.4.
Case 1: $V_{1}\cap V(\mathcal{H})\neq\emptyset$. There exists $v\in V_{1}\cap V(\mathcal{H})$ such that $\{v\}\notin E(\mathcal{H})$. By Lemma 3.2, $v$ is simplicial in $\mathcal{H}$. Suppose $\{v\}\in E(\mathcal{H})$ for all $v\in V_{1}\cap V(\mathcal{H})$. This implies that for every $x\in V(G)\setminus V(H)$, either $\{x\}\in E(\mathcal{H})$ or $x\in V_{c}$.
Claim 1: If $y\in V(\mathcal{H})\cap V(H)$ and $\{y\}\notin E(\mathcal{H})$, then $(\operatorname{SC}(y)\setminus\{y\})\cap V(\mathcal{H})=\emptyset.$
Suppose, for contradiction, that there exists $c\in(\operatorname{SC}(y)\setminus\{y\})\cap V(\mathcal{H})$. Since $\{c\}\in E(\mathcal{H})$, there exists an edge $f\in E(\mathcal{C}_{r}(G))$ of the form $f=\{b_{1},\ldots,b_{r},c\},$ where $f\cap V(\mathcal{H})=\{c\}$. However, $\{b_{1},\ldots,b_{r},y\}$ is connected in $G$, which would force $\{y\}\in E(\mathcal{H})$, a contradiction.
Claim 2: If $e\in E(\mathcal{H})$ and $|e|\neq 1$, then $e\in E(H)$.
Let $e\in E(\mathcal{H})$ with $|e|\geq 2$, and suppose $e=f\cap V(\mathcal{H})$ for some $f\in E(\mathcal{C}_{r}(G))$. If $|e|\geq 3$, take $a\in e$ and choose $x\in e\setminus\{a\}$ minimizing $\operatorname{d}(a,x)$. If $a$ and $x$ were non-adjacent in $G$, there would exist a shortest path $P(a,x)=(a=y_{0},y_{1},\ldots,y_{t}=x)$ in $f$ with all $y_{i}\notin V(\mathcal{H})$ for $1\leq i\leq t-1$ (since any $y_{i}\in V(\mathcal{H})$ would yield $\operatorname{d}(a,y_{i})<\operatorname{d}(a,x)$). Considering $\mathrm{SC}(y_{1})$, any $c^{\prime}\in\mathrm{SC}(y_{1})\setminus\{y_{1}\}$ with $\{c^{\prime}\}\in E(\mathcal{H})$ would imply $\{a\}\in E(\mathcal{H})$, contradicting $\{a\}\notin E(\mathcal{H})$. Thus, $\mathrm{SC}(y_{1})\setminus\{y_{1}\}$ must be disjoint from $V(\mathcal{H})$. The set $Q^{\prime}=\mathrm{SC}(a)\cup\mathrm{SC}(y_{1})$ is connected with $Q^{\prime}\cap V(\mathcal{H})=\{a\}$, forcing $\{a\}\in E(\mathcal{H})$, another contradiction. Hence, $a$ and $x$ must be adjacent. Since $|\mathrm{SC}(a)\cup\mathrm{SC}(x)|\geq r+1$, there exists a connected set $g\subset\mathrm{SC}(a)\cup\mathrm{SC}(x)$ with $|g|=r+1$ containing $\{a,x\}$, yielding $g\cap V(\mathcal{H})=\{a,x\}\in E(\mathcal{H})$. This contradicts $|e|\geq 3$, so we must have $e=\{a,x\}\in E(H)$.
[…]
In both cases, $\mathcal{H}$ has a simplicial vertex, proving the theorem. ∎
The following example demonstrates that the complex $\mathcal{I}_{r}(G)$ may fail to be shellable when $r>2t+1$, even if $G=\operatorname{CCG}(H,V(H),t)$ for a chordal graph $H$ and $t\geq 1$.
Example 5.3.
Let $H$ be a chordal graph with vertex set $V(H)=\{a_{1},\ldots,a_{n},b_{1},\ldots,b_{n},c,d\}$ for some integer $n\geq 2$, and edge set
$E(H)=\$$\big\{\{a_{i},a_{i+1}\}\mid 1\leq i\leq n-1\big\}\cup\big\{\{b_{i},b_{i+1}\}\mid 1\leq i\leq n-1\big\}$
$\cup\ \big\{\{a_{1},c\},\{a_{1},d\},\{b_{1},c\},\{b_{1},d\},\{c,d\}\big\}.$
Let $G=\operatorname{CCG}(H,V(H),t)$ be the graph obtained by attaching a star-clique graph $\operatorname{SC}(x)$ of size $t+1$ to each vertex $x\in V(H)$. We claim that $\mathcal{I}_{r}(G)$ is not shellable for $r=nt+n$. Assume, for contradiction, that $\mathcal{I}_{r}(G)$ is shellable. Define the set
$S=\left(\bigcup_{i=1}^{n}\operatorname{SC}(a_{i})\cup\bigcup_{j=1}^{n}\operatorname{SC}(b_{j})\cup\operatorname{SC}(c)\cup\operatorname{SC}(d)\right)\setminus\{a_{1},b_{1},c,d\}.$
By repeatedly applying [25, Proposition 2.3] and [33, Lemma 2.2(1)], it follows that $\mathcal{I}(\mathcal{C}_{r}(G)/S)$ is shellable. However, one can verify that $\mathcal{C}_{r}(G)/S$ is isomorphic to the $4$-cycle $C_{4}$. By [32, Theorem 10], the complex $\mathcal{I}(C_{4})$ is not shellable, yielding a contradiction. Hence, $\mathcal{I}_{r}(G)$ is not shellable when $r=nt+n>2t+1$.
In [8], Cook II and Nagel introduced the notion of a clique whiskering of a graph. We generalize their construction as follows. Let $G$ be a graph. A clique vertex-partition of $G$ is a collection $\pi=\{W_{1},\ldots,W_{k}\}$ of pairwise disjoint cliques (some of which may be empty) such that $\bigcup_{i=1}^{k}W_{i}=V(G).$ Given a graph $G$, a clique vertex-partition $\pi=\{W_{1},\ldots,W_{k}\}$, and an integer $r\geq 1$, the $r$-clique whiskering of $G$ with respect to $\pi$, denoted $G^{\pi}_{r}$, is defined as follows. For each $1\leq i\leq k$, let $t_{i}\geq r$ be an integer. The vertex set of $G^{\pi}_{r}$ is
$V(G^{\pi}_{r})=V(G)\cup\bigcup_{i=1}^{k}\{x_{i,1},x_{i,2},\ldots,x_{i,t_{i}}\},$
where each $x_{i,j}$ is a new vertex added to $G$. The edge set of $G^{\pi}_{r}$ is
$E(G^{\pi}_{r})=E(G)\cup\bigcup_{i=1}^{k}\left\{\{a,b\}\mid a\neq b,\ a,b\in W_{i}\cup\{x_{i,1},x_{i,2},\ldots,x_{i,t_{i}}\}\right\}.$
In other words, for each clique $W_{i}$ in $\pi$, we introduce $t_{i}$ new vertices and form a clique on the vertex set $W_{i}\cup\{x_{i,1},\ldots,x_{i,t_{i}}\}$. Note that a graph $G$ may admit multiple clique vertex-partitions. In particular, every graph has at least one such partition, namely the trivial partition $\pi=\{\{x_{1}\},\ldots,\{x_{n}\}\}$, where $V(G)=\{x_{1},\ldots,x_{n}\}$.
Theorem 5.4.
Let $G$ be a graph, and let $\pi=\{W_{1},\ldots,W_{k}\}$ be a clique vertex-partition of $G$. Then $\mathcal{I}_{r}(G^{\pi}_{r})$ is shellable for all $r\geq 1$
Proof.
By Theorem 2.2(1), it suffices to show that every $c$-minor of $\mathcal{C}_{r}(G^{\pi}_{r})$ contains a simplicial vertex. Let $\mathcal{H}$ be a $c$-minor of $\mathcal{C}_{r}(G^{\pi}_{r})$, and let $V_{1}$ denote the set of all simplicial vertices in $G^{\pi}_{r}$. First, suppose that $V_{1}\cap V(\mathcal{H})\neq\emptyset$. If there exists a vertex $v\in V_{1}\cap V(\mathcal{H})$ such that $\{v\}\notin E(\mathcal{H})$, then by Lemma 3.2, $v$ is simplicial in $\mathcal{H}$. Otherwise, assume $\{v\}\in E(\mathcal{H})$ for all $v\in V_{1}\cap V(\mathcal{H})$. We claim that in this case, $\{a\}\in E(\mathcal{H})$ for every vertex $a\in V(\mathcal{H})$. If $a\in V_{1}\cap V(\mathcal{H})$, then $\{a\}\in E(\mathcal{H})$ by assumption. Now suppose $a\notin V_{1}$. Then $a\in V(G)$. Since $\pi=\{W_{1},\ldots,W_{k}\}$ is a clique vertex-partition of $G$, $a\in W_{i}$ for some $1\leq i\leq k$.
Consider two subcases: Suppose $\{c\}\in E(\mathcal{H})$ for some $c\in\{x_{i,1},\ldots,x_{i,t_{i}}\}$. Then there exists a connected set $\{c_{1},\ldots,c_{r},c\}$ in $G^{\pi}_{r}$ such that $c_{i}\notin V(\mathcal{H})$ for all $1\leq i\leq r$. Since $W_{i}\cup\{x_{i,1},\ldots,x_{i,t_{i}}\}$ is a clique, it follows that $\{c_{1},\ldots,c_{r},a\}$ is also a connected set, so $\{a\}\in E(\mathcal{H})$. Suppose $\{x_{i,1},\ldots,x_{i,t_{i}}\}\cap V(\mathcal{H})=\emptyset$. Since $t_{i}\geq r$, we can select distinct vertices $x_{i,1},\ldots,x_{i,r}$ such that $\{x_{i,1},\ldots,x_{i,r},a\}$ is connected in $G^{\pi}_{r}$, hence $\{a\}\in E(\mathcal{H})$.
Thus, in either case, $\{a\}\in E(\mathcal{H})$ for all $a\in V(\mathcal{H})$, so $\mathcal{H}$ is totally disconnected (i.e., all its edges are singleton sets).
Now suppose $V_{1}\cap V(\mathcal{H})=\emptyset$. Then every vertex of $\mathcal{H}$ lies in $V(G)$. Repeating the same argument as above, we again conclude that $\mathcal{H}$ is totally disconnected.
In both cases, $\mathcal{H}$ is totally disconnected, hence contains a simplicial vertex (every vertex is simplicial). This completes the proof. ∎
We introduce the notion of a clique cycle graph, which generalizes the classical concept of cycles. An $n$-clique cycle, denoted by $\operatorname{CC}(B_{1},\ldots,B_{n})$, is a graph formed by a cyclic sequence of $n$ cliques $B_{1},\ldots,B_{n}$, where $n\geq 3$. Each consecutive pair of cliques $B_{i}$ and $B_{i+1}$ (indices modulo $n$) intersects in exactly one vertex, denoted $x_{i}$. In particular, $B_{n}$ and $B_{1}$ share the vertex $x_{n}$, thereby completing the cycle. The vertices $x_{1},\ldots,x_{n}$ are called the connecting vertices of the $n$-clique cycle. We define a graph $G(r)$, for a given integer $r\geq 1$, as follows: start with an $n$-clique cycle $\operatorname{CC}(B_{1},\ldots,B_{n})$ for some $n\geq 3$, and select a non-empty subset $A\subseteq\{x_{1},\ldots,x_{n}\}$ of the connecting vertices. To each vertex $x\in A$, attach a star-clique graph $\operatorname{SC}(x)$, such that $|V(\operatorname{SC}(x))|\geq r+1$ for at least one $x\in A$.
Theorem 5.5.
Let $G=G(r)$ for some $r\geq 1$. Then $\mathcal{I}_{r}(G)$ is shellable.
Proof.
By Theorem 2.2(1), it suffices to show that every $c$-minor of $\mathcal{C}_{r}(G)$ contains a simplicial vertex. Let $\mathcal{H}$ be a $c$-minor of $\mathcal{C}_{r}(G)$, i.e., $\mathcal{H}=\mathcal{C}_{r}(G)/S$ for some subset $S\subseteq V(G)$. Since $G=G(r)$ is constructed from an $n$-clique cycle with attached star-clique graphs, there exists a connecting vertex $x\in\{x_{1},\ldots,x_{n}\}$ such that $|V(\operatorname{SC}(x))|\geq r+1$. Let $V(\operatorname{SC}(x))=\{z_{1},\ldots,z_{m},x\},~\text{with }m\geq r.$ We consider the following two cases:
Case 1: $\{z_{1},\ldots,z_{m}\}\cap V(\mathcal{H})=\emptyset$. We further divide this case: Suppose $x\in V(\mathcal{H})$. Then, since $x$ forms part of a star-clique of size at least $r+1$ and its neighbors are excluded from $\mathcal{H}$, we have $\{x\}\in E(\mathcal{H})$. By Lemma [26, Lemma 3], $\mathcal{H}\setminus x=(\mathcal{C}_{r}(G)/S)\setminus x=\mathcal{C}_{r}(G\setminus x)/S.$ Observe that $G\setminus x$ is a $\mathcal{T}_{1}$-graph. By Theorem 4.15, the complex $\mathcal{C}_{r}(G\setminus x)$ has a simplicial vertex, and hence so does $\mathcal{H}\setminus x$. Therefore, $\mathcal{H}$ itself has a simplicial vertex. Suppose $x\notin V(\mathcal{H})$. Choose a vertex $c\in V(\mathcal{H})\cap\{x_{1},\ldots,x_{n}\}$ such that $\operatorname{d}(c,z_{m})\leq\operatorname{d}(a,z_{m})$ for all $a\in V(\mathcal{H})\cap\{x_{1},\ldots,x_{n}\}.$ That is, $c$ is the connecting vertex closest to $z_{m}$ among those present in $\mathcal{H}$. Therefore $\{c\}\in E(\mathcal{H})$. Repeating the same argument as in above, we conclude that $\mathcal{H}$ contains a simplicial vertex.
Case 2: $\{z_{1},\dots,z_{m}\}\cap V(\mathcal{H})\neq\emptyset$. Let $y\in\{z_{1},\dots,z_{m}\}\cap V(\mathcal{H})$. If $y$ is not a singleton edge in $\mathcal{H}$, then Lemma 3.2 implies that $\mathcal{H}$ contains a simplicial vertex. If $y$ is a singleton edge in $\mathcal{H}$, then by an argument analogous to Case 1, $\mathcal{H}$ must also contain a simplicial vertex. In all cases, the $c$-minor $\mathcal{H}$ has a simplicial vertex. Thus, by Theorem 2.2(1), $\mathcal{I}_{r}(G)$ is shellable. ∎
The following example illustrates that the conclusions of Theorems 5.1, 5.4, and 5.5 may fail when $\operatorname{CCG}(H,S,t)$, $G^{\pi}_{t}$, or $G(t)$ is considered with $t<r$.
Example 5.6.
Let $H$ be the $4$-cycle on vertices $\{x_{1},x_{2},x_{3},x_{4}\}$, and let $G$ be the graph obtained by attaching a whisker to each vertex of $H$. Clearly, $G$ lies in the class of graphs considered in Theorems 5.1, 5.4, and 5.5. Now, set $r=2$. Let $K$ denote the set of newly added whisker vertices, i.e., $K=V(G)\setminus\{x_{1},x_{2},x_{3},x_{4}\}$. Using [25, Proposition 2.3] and [33, Lemma 2.2(1)], we obtain $\operatorname{link}_{\mathcal{I}_{r}(G)}(K)=\mathcal{I}(\mathcal{C}_{r}(G)/K)=\mathcal{I}(C_{4}).$ By [32, Theorem 10], the independence complex $\mathcal{I}(C_{4})$ is not shellable. Hence, $\mathcal{I}_{r}(G)$ is not shellable either.
Acknowledgments: We express our sincere gratitude to Priyavrat Deshpande, Anurag Singh, and Russ Woodroofe for their valuable insights and clarifications, which significantly contributed to the development of this work. The second author was partially supported by the National Board for Higher Mathematics (NBHM) and the Science and Engineering Research Board (SERB).
Data availability statement. Data sharing not applicable to this article as no datasets were generated or analysed during the current study.